A Bibliography of Publications about Typesetting

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

08 May 2024 Version 2.85

Title word cross-reference

5 × 7 [MBG77].

1.5A1 [Ano86f]. 10 [Sey66]. 10/100 [Com83]. 10/66 [Sey66]. 10/66-2R [Sey66]. 100-font [BN94]. 106 [FN85a, FN85b]. 1130 [IBM67]. 14th [Ano93a]. 15th [GR94]. 16507-2 [Deu98]. 1846 [Ano96a]. 1966 [Jas67]. 1970 [Lan71]. 1975 [Ano75]. 1984- [LA84]. 1988 [ZD91, vV88]. 1992 [Zla92]. 1995 [Mac96]. 1998-04 [Deu98]. 1st [ICC65a].

2 [ISO09, Sam63]. 202 [CKT80]. 23 [WW03]. 24-Wire [BM87]. 25th [WH86a]. 27 [ZD91]. 2R [Sey66].

30th [WH86a]. 3rd [Ano92].

4th [HAB98].

66-2R [Sey66].

7-10 [VC92]. 76 [Car80]. 7th

[HAB98, Zla92].

80000 [ISO09]. 80000-2 [ISO09]. 80’S [Dun80]. 8400 [Lev84].

’92 [VC92, Zla92]. ’95 [Lev95]. 9541

[SBSG89]. ’98 [HAB98].

= [Can93, ICC65a, MB96].

ABC [Lin82]. ABCs [Col89]. ABP [Lip78]. Abstract [Bro69]. academic [Wal82]. Accept [LMT^{+}21]. acceptance [Gul62]. Access [Day83, PRB87]. accessories [Lin30, Linxx]. According [Ano86b]. accuracy [AIDG95, WT28]. achieve [MR92b]. ACM [SIG82, Abr81]. acoustooptic [Ano96d]. acquisition [CBT88]. ACS [Dod97]. Adaptive [CW85, Ano96f]. Addressable [Ano86a]. Addressing [Ano86d]. Adjust [Ano85b]. Adobe [Ben95, WG19]. Advanced [Her81, Kim80, LK93, RS81]. Advances [Blu64, CTC66, Jas67]. Advantage [Zur85]. advertisement [Wil37]. advertisers [Ame30]. advertising [Anoxx]. affect [Ber00]. African [WH86a]. after [Ber23]. age [MH98, San99, ZD91]. Ago [Ano96a, Ano19, Tho21]. Aid [Dow98, Pin81, TBF^{+}77]. Aided [Ulu86, Ber67, Fri70, Gol69, Hol84, SPD72]. Aids [CBB54, CBB65, CBB87]. AIP [Ano90a]. Aldine [Fle88]. Aldus [Fle88]. Aleph [SW92]. Algebraic [Lev95]. ALGOL [von67]. Algorithm [BDN83, Smi84, HYH89, HYH91, MP94, dG99]. All-Points-Addressable [Ano86a]. allowed [Tho21]. almost [Ano19]. Alphabet [Zap07a]. Alphabetgeschichten [Zap07b]. Alphabetical [Ros15]. Alphabets [ST85, Kel11]. Also [Phe89]. alternative [Bee86]. alternatives [Le 88]. am [SV73]. amateur [Har83]. America [Ann75, Ann94, Ano97]. American [Ano97, Big80, FM94, Lin52, McG86, McG93, Par75, Sax92, Tho77, Tho96]. Americanism [Lin52]. among [Gul62]. Ampersands [Hou13a]. AMS [Spi86]. AMS-TEX [Spi86]. Analysis [MMN79, RS84, Eze87, Hau94, SO94]. annealing [Lag95]. anniversary [Fri83]. annotated [Sch95]. Annotations [Fre96]. Annual [Ano91, Ano93a, GR94, Mor93, WW03]. Ansaetze [ZK88]. any [KPB87]. Anything [Per88]. APL [APL87, Dic87, GSMN93, Hoh88, How86]. APL2TEX [GSMN93]. App [SH19]. Appendix [MC75a]. Apple [Knu91c, Wag85]. appliances [Wil31]. Application [Sch83, Lop94]. Applications [HL87, Zey74, Ano87b]. Applied [CKM86, CK86, Pin94]. appreciated [Kra96]. Approach [GW85, Gra86, Kor75, LZ86, BS96, Krt94, Won86a, Won90]. approaches [MRT89]. April [HAB98, VC92, vV88]. APSQ [Com84]. arabe [Ham75]. Arabic [Ano92, AA89, All94, AM86, Ano96e, EEN90, Lag92a, Lag92b, Mac90, Niz90, Ploxx, Str93, TA82, ZA87]. ArabTEX [Lag92b, Lag92a]. Arbitrary [MN77]. architecture [Ano96f]. Area [Her78a, Her78b, WH86b, DS83]. Arizona [USE87]. Arrangement [Dow66, Dow93, Dow95]. arrows [Ros92a]. Art [Abi88, Bor85, Ham68, Lin52, Ste65, Ste70, Ano19, Har83, Kel10, Smi17, SL67, Ano97]. article [VDM^{+}93]. Artist [Wri83]. artistic [HAB98]. Asian [WH86a]. ASIS [Lan71]. aspect [FM94, Som90]. Aspects [Ric78]. ASQA [Ano96f]. assembly [Ano90b]. assessing [UB95]. assessment [Ber00, Hus76]. assistants [Swa71, Swa79, Swa86, Swa98]. Assumptions [BM87]. Athens [Mac96]. Atkyns [Bli82]. attitudes [Skr59]. attribute [Lin95]. ATU [TA82]. audio-visual [Tex84]. August [WH86a, Lop94]. Ausgabe [Deu98]. Aussagen [Zap68, ZB72]. Author [Doo89, Ano52, Til52, Til55]. Authored [Kee80]. Authors [CBB54, CBB65, CBB87, Dod97, Kel85, Mas05, McS84, Swa71, Swa79, Swa86, Swa98, Zap80]. authors/typists [Zap80]. automate [CBT88]. Automated [Bra09, Gar80, Kee80, Lag95, MR92b, RM74, Com64, Com66, Com69, Des89, SS75]. Automatic [Ano86b, ICC65a, MP94, MPZ95, Ste65, SL67, Ste70, SOE77, TCLM82, Buc64, Cla64, JK74, MH70, Pla81, Ter68, TH73].

Automatically [Ano85b, BS96, Hob93]. Automation [For89, Fre96, Rad77, YKT90]. automatique [ICC65a]. Automatische [Roe85]. Automatisches [ICC65a]. Auxiliary [Her71]. available [Lan40, Col30, Lan49].

B [MC75a]. Babbage [Bux88]. Back [HYH89, HYH91, Ano19]. Back-propagataion [HYH89]. Back-propagation [HYH91]. Bangla [PC94]. Bangladeshi [PC94]. Bangor [Uni74]. Barbara [GR94]. Barnett [Sey66]. Based [Gel91, Pow86, BS96, GMZZ94, KAT90, Lev87, MPZ95, TC92]. Basic [BL84, CBM99, Gho84, Col62, Cra83a, Kle79, RM74, Sac86, Tay16, Her95]. BBN [Ano96e]. Be [Pod84, Ame30, ISO09, Lan49]. beautiful [Ber23]. Beauty [Kad11]. Bedeutung [Thi83]. beginner [Ano86i]. Beginning [SP87]. Begriffe [Deu94]. Being [Kra96, Ano19, Col30, Gou78]. Beispiel [SV73]. Bell [Lan40]. Benefits [Met75, Gib95]. Benjamin [Sta06]. Benton [Cos86]. beta [LH91]. beta-spline [LH91]. Better [Phi79, PK82, UBR90]. between [Ploxx, WT26, WT33]. Bézier [FB85, MP94]. Bible [Har94b]. Bibliographic [AKWC82, All81]. bibliography [Gil91, Gra66, Gra69, Nel88]. Binary [PS84]. bindery [Sta71b, Sta71a]. binding [Bar71]. biographical [ERBR91]. biographies [BMD87]. biography [Mer98, MS89]. Birmingham [Ano93a]. birth [Fri83]. bitmaps [Hob93]. Bits [Bra13a]. black [Smi17]. blind [Tho21]. block [SO94]. block-class-specific [SO94]. Board [WT28, WT26]. Boards [Boa84]. boiling [Smi17]. Book [Ano31, Bor85, Her78a, Her78b, Mas05, MW05, Ric78, San99, Sey66, Woh98, Aus88, Ame30, Ano52, Bur61, Bli82, Buc64, Dun83, Fos81, Kel10, Kun79, LA84, LA85, Til52, Tho77, Tho96]. Book-Like [Woh98]. Books [PRB87, Col62, Cro80, Fos81, Gou78, HK99, HK04, Mul94, Mul05]. borders [Ame30]. Both [LZ86, Ame30, Lan49, Tay16]. Box [Krt94]. Boyd [Cos86]. Braille [AJ88b, AJ88a, AJ90]. Branding [RPP96]. Breaking [Hol06, KP81, Ano07]. breaks [PK82]. Brewer [Uni74]. Bridging [dG99]. brief [Phe61]. Bringing [Mun88]. Brings [Blu64]. Brochure [Ano14]. Brockmann [Mül96]. brought [Ano19]. Browne [Ano80]. Brushless [JSG78]. Buchstaben [Poh15]. budget [Eis90]. Buechern [Man87]. Buffer [Ano86h]. Buffers [Buy85]. build [Kle79]. Bulletin [Bro69]. Burns [Bur61]. business [Ame83, Ame84, Ame88, Buh80, Buh81, Int06]. Buyer [Ano89a]. Buyers [Lab84]. Buying [Nie80]. Byblos [Ano96e]. BYSO [Lev87].

C [Hol87a]. CA [GR94]. caching [FK85]. calligraphic [Par93b]. Calligraphy [Her67, Cus11]. Cambridge [Sey66, All81, AKWC82, UCBB83]. Camera [Ano95a, Ano95b, Ano95c, Ano98, Zap80]. Camera-ready [Ano95a, Ano95b, Ano95c, Ano98, Zap80]. campus [Rob72]. Can [LMT^{+}21, MR92b]. Canada [Lev95]. cap [Ano80]. capitalization [Int06]. cards [Buc64, Cus11]. care [Wil31]. Cartography [Her69, Her79, Her95]. Cartridge [Ano86g]. Case [Per88, Phi79, Ste19]. Caslon [Bli82]. cast [Lan49]. Catalan [FF93]. catalog [Buc64]. catalogs [Ann75, Ann94, Buc64]. catalogue [Sch95]. Cathode [Kim80]. Cathode-Ray [Kim80]. Celestial [Her79]. census [Ano84, Dep79]. centenary [Bia93]. Center [AKWC82]. Centre [All81]. centroid [AIDG95]. century [Gou78, HK99, HK04, Kel10, Skr59, Sou05, Tho21, Wal84, McG86, McG93]. Certain [Des89]. Chain [TKE83]. change [Ano95d, Edv94, May69]. Changeable [BG73]. changed [Fre90, War12]. Changes [BM87, Mas84, Kir16, Knu86f, Knu89]. Changing [Ano86b, Cle82, Sha85, Eze87]. chap [Gou78]. Chapter [Lan71]. Character [Ano86e, BG73, CKM86, CK86, KM87, MRS83, MS84, NSA90, Rei66, SP87, Som90, Ste87, THL80, Tri69, UKKS84, UM85, Wol86, All94, Ano96f, AIDG95, ELS94, FTBM96, HD88, KD94, Krt94, LK95, Lin95, LCO94, Smi90, UBR90]. Characteristics [RF77]. Characters [Ano85a, Ano85c, Hou13a, Hou13b, KB86, KPB87, LZ86, MBG77, Ume82, Ano96b, Hus95, Pow90, TC92]. Charles [Bux88, GS21]. chart [Szo82]. Check [NIS04]. CHEM [BJK87]. Chemical [BJK87, Boa84, HO87, Haa86]. Chemlaw [Mar83]. Chicago [SIG82, Ano82a, Ano93c, Ano03]. child [FM94]. Chinese [Lun99, Ano83b, Ano96b, Ano96f, BS96, CH85, IBC87, LZ86, Lin95, MPZ95, RRLG83, TC92, Ume82, UM85, WH95]. Chinese/Japanese [IBC87]. chip [Ano96f]. Choosing [PK82, Szo82]. Christopher [Dur22]. Chronicle [Zap07a]. Chronik [Zap07b]. chronology [Wal88]. Church [Hus76]. CICERO [AC79]. Circuit [Boa84, Pad87]. circuits [Sim87]. circular [AAP94]. CIS [Com64]. Civil [Par75, Int06]. CJKV [Lun99]. Clark [ZD91]. Class [Ano93b, CSR10a, SO94]. Classic [BHH^{+}99, PR83, PR04]. Classical [ZD91]. Classification [ELS94, LCO94]. classifier [BN94, CDV95]. Classifiers [SV73]. Classroom [Clo80]. Claude [Ber23]. clichage [Sta71b, Sta71a]. Closer [BDH^{+}94, BHH^{+}99, BDH02]. Club [WW03]. cluster [UB95]. cockpit [UB95]. code [Lev87, Lin95]. Coded [TKE83]. codes [Aus88]. Coding [Kor75]. Cold [Gra75, Kro67, Jac74, Rob72]. Colines [Sch95]. collation [BMD87]. collection [CIS67]. Columbus [Col30]. Column [Ano85b]. Command [Ano86h]. commands [Com83]. comments [AA89]. Commercial [Dun80, Gul62, Sey82]. commerciale [Sta71b, Sta71a]. Committee [WT28, WT26]. Common [Cra87a, MC75a]. commonly [Fos76]. Communication [CDM02, Pod84, IEE93, MB96]. Communications [Her78a, Her78b, Mor87, Goo82]. Como [Luc85]. Comp [Gra75]. Companion [GRM97, MFW23, MFB^{+}23]. Company [Ano31, Ber23, Col30, Rom14]. Comparative [Tay16, SZ91]. Comparison [MRT89, SV73, Gul62]. Comparisons [MBG77, LCO94]. Compilation [Hau22]. Complete [Fel03, Har83, Lin82]. Completely [Ale92]. Complex [BKKW98, Woh98]. components [WH95]. composer [Gar73, Bro69]. composing [Lin30, Phi68, Ploxx]. Composition [Ale92, BE92, Com83, GS70, ICC65a, Lan71, Mac75, MNM79, Par75, Rub88, Sha85, Bar87, Lan40, Ber75, Can93, Hen73, Hol84, Hus85b, Int06, KAT90, Kro67, Kun79, Lan49, LK95, MM73, Deu94, Sta71b, Sta71a]. Compositors [CBB54, CBB65, CBB87, Har91, BMD87]. Comprehension [Cle82]. comprehensive [Fos76]. Compression [Ano86c, BHW83, IKM82, SO94]. Compression/Decompression [Ano86c]. Compugraphics [Lev84]. Compugraphics-8400 [Lev84]. Computation [Lev95]. Computational [Par93a]. Computer [AKWC82, AK73, Bar65, Ber67, CTC65, Cat83, CSR10a, Cle82, CW82, Dav69, DEM^{+}63, Gel91, Gol69, Gra66, Gra69, Hah93, Her72, Her81, Jas67, Kee80, Lan71, Le 88, LK77, Lin90, Mar72, McS84, MC75b, Neg88, Phi68, Pie71, Pow86, Rub88, Sac80, Sam82, Ser84, SPD72, Wag85, ZJ84, von67, All81, Ano98, BS88, Bee86, Com64, Com66, Com69, Car80, Dal80, FN87, Fri70, GB82, Goo82, Hah91, Hen73, Hol84, CTC66, Int82, Int86, KR85, MH70, Pub71, Rom75, Spi80, Spi82, Spi83, Com84, WH76, ZD91, Knu86d, Sey66, Ulu86, WH17]. Computer-aided [Gol69, SPD72, Fri70]. Computer-Based [Pow86]. computerisation [Hus95]. Computerized [Ano82b, AJ88b, AJ88a, AJ90, Bea77, CCT65, ICC65b, LB74, Met75, RRLG83, ST85, SDB88, CIS67, Nat87a, Nat87b, TH73]. Computerphile [Bra13a, Bra13b]. Computers [Ano91, Ghe87, Her67, Knu86e, Knu86f, Knu89, SOE77, Win90]. Computing [Lab17, Lun99, Ano92]. COMTEXT [Car80]. Concept [Wri81]. concepts [Com00, FSS82]. concise [Wal88]. Conference [APL87, Ano75, Ano91, CIS67, CTC65, Dun80, For89, HAB98, CCT65, ICC65b, CTC66, Jas67, LBJ95, Luc85, USE87, VC92, Zla92, vV88, SIG82, Ano92, Des86, SPD72]. Congress [Buc64, ICC65a, WH86a]. Connectionist [GSKM92]. connectivity [Krt94]. Considerations [Ben85, Boe76, Gol73, PRB87]. Constructing [Wal87]. Construction [LLW95, UM85]. contemporary [Sha77]. Content [Bra09, Pow62]. Contents [EGL^{+}97, VDM^{+}93]. context [KH94]. Continued [Ano97, Hel99]. Continues [Eis88]. Continuity [FB85]. contour [NSA90]. contribution [Cus11, Ham75, WH76]. contributions [Cos86]. Control [Rac70, RS81, SP87, FZ93, Pos78]. controller [Nas94]. Controlling [Ser84]. conventions [NIS04]. Conversion [Ano96b, BDN83]. Converting [ZJ84]. coordinates [WH76]. Coping [Smi89]. Copy [Kas72, Bar71, Lip78]. copyediting [Swa71, Swa79, Swa86, Swa98]. Copyflow [Kun79]. CORE [EGL^{+}97]. Corporation [Le 88]. Correct [Ghe87]. correcting [BN94, Ste19]. correction [Ano52]. Corrections [Sta06, Til52, Til55].

Corrosion [Boa84]. Cost

[BG73, Buy85, Col86, HH83, Pin81, Ser88]. Cost-Effective [Pin81]. Costs [Mac75, Ano80, Le 88, MM73, Tay16]. Council [CIS67]. count [FBJ90]. Counting [San95]. Course [CBM99, Cra83a, RM74]. CPU [KKM^{+}93, KKM^{+}94]. Craft [MR92b]. Crashkurs [Kha00]. Create [Buy85]. creation [Buc64]. credited [Ber23]. Critical [BDH^{+}94, BDH02, HM01]. crosswords [vdL92]. CRT [Gol69, Kim80, MRS83, Pad87, Rac70, Ric81, THL80]. Crystal [Ano86d]. Crystallographic [All81, AKWC82]. CTS [Kok87]. Cueing [FC77]. Culture [LR94, Ham75]. Curiosities [Hou13a]. Cursor [Ano85c]. curve [LH91, MP94]. Curves [FB85]. Custom [RS81]. cuts [MM73]. Cwpr [IBC87]. Cyber [Car80]. Cyber-76 [Car80]. Cyrillic [Str93]. Czechoslovakia [Zla92].

daily [Dep72, May69, Ten76]. Dallas [APL87]. Dànielle [Dur22]. Data [All81, AKWC82, Ben88, CSR10a, IKM82, RS84, Buc64, CBT88, FZ93, LA84, LA85, Rom75, SO94]. Database [HYC82, ZJ84, Myh90]. Databases [Kim90]. Datasouth [Bry84]. d’Athènes [Mac96]. Day [Dow98, Ame30]. days [Str93]. DC [Lan71, CCT65, ICC65b, JSG78]. December [Ano92, Lan71, Ano96a]. Decisions [Wag85, Ros92a]. Decompression [Ano86c]. defects [UB95]. deformations [Par93a]. Delivers [RS81]. Delivery [Tul85]. demands [Nas94]. démocratisation [Ham75]. demonstration [Col30]. Denotational [Pin94]. department [Jac74, Wil37]. Description [Ges86, Fos81]. descriptions [CDV95]. Design [Ben85, BDH^{+}94, BHH^{+}99, BDH02, BCJR87, CDM02, Dow98, Gol73, Hel99, Kad11, Kim80, Kin93, Meg98, MMN79, MNM79, Mül96, NKMO89, PRB87, Ric78, Rub88, SOE77, SDB88, Tri69, Wag85, ZL87, ZA87, Cos86, Cra83b, Cus11, Gho84, Gou78, Hal80, Hal81, Lab88, LLW95, LK95, Mas05, MR92a, MB96, Par93a, Par93b, Ros92a, Som90, Sou05, Tho77, Tho96, Uni78, Pri80]. designed [Kel19]. Designer [Hel98, Mül96, PK02, TR02, MW05]. designers [Kel10, MB96]. Designing [Cra83a, CBM99, WH95]. designs [Gou78, Zap68, ZB72]. Desk [Nat86]. Desk-top [Nat86]. Desktop [Ano96c, Cis87, Cra87b, FN87, Gel91, Mor87, RM19, WH17, YKT90, Bee86, Cra87a, Kle87, Kle90, Nel88, Ser88, GK00]. Desktop-publishing [GK00]. details [Har83]. Detection [MC75b]. Deutsche [BN85]. develop [Rom76]. developed [Ano96f]. Development [Ano83b, Dow98, SP87, Wag85, Hus73, Hus85b, She50]. Developments [Zap07a, Hol84, Wal88]. Device [Bro85, IBC87, Wat84, WW82]. Device-Independent [Bro85, IBC87]. devices [WW82]. DeVoss [Dur22]. DFORMAT [Ben88]. Diagnostics [EM84]. diagram [Ros92a]. Diagrams [BJK87, Ros92b, Ing91]. Dictionary [Fos76]. differ [BL84]. Difference [Ploxx, Pap91]. Different [Ale92, WT28, Zap68, Zap70, ZB72]. Digiset [Zey74]. Digital [Ano19, BBK13a, BBK13b, BD83, Col86, EGL^{+}97, Knu99, McC20, Rac70, Rub88, Wad84, Ano96f, Bry96, BW95, Col89, HAB98, Hop12, Lev84, MH98, Mor89, San99, Sey84b, Sey84a, Deu98]. Digitaler [Deu98]. Digits [Ast72]. dimensional [MB96]. Dimensions [AKWC82, All81]. DIN [Deu98]. DIPS [FN85a, FN85b]. DIPS-106 [FN85a, FN85b]. direct [KNS90]. Directions [Knu79]. Directors [WW03]. directory [ERBR91]. Discrete [WH86b]. Discrete-Element [WH86b]. Display [AMS83, Ano86h, Lan49, Pad87, RS84, Tuf83, UB95]. Displayed [Kor75]. Displays [MRS83, WH86b, Niz90]. disquisition [Kra96]. Distributed [BW84]. distribution [Buc64, ELS94]. dithering [AIDG95]. ditroff [AB89, BD87, Pow90, BD88, Guyxx]. diversions [WBS82]. Division [Sam82]. Do [Bry84, Cle82, Clo80]. Do-It-All [Bry84]. Document [BW82, BW84, FSS82, Geh86, Geh87, GL88, Lam86, Lam94, NCNS82, SO94, Sey82, SP87, Sri93, Tul85, VC92, VP94, vV88, ANP81, But94, CHM00, Bea88b, KL82, KH96, Kin93, McC88, WBS82, Won90]. Document-Preparation [BW82]. Documentary [Fle88]. Documentation [CW82, HYC82, Luc85, MPP81, Mar72, Des86]. Documents [Bar64, BKKW98, GRM97, Hah93, TCLM82, Woh98, AB89, ACG94, Bea77, Hah91, Uni85a]. does [Ber00]. Dokumentation [Ulu86]. Dokumenteneditor [Sch86a]. dominate [Ben95]. Don’t [CSR10b]. Doors [Dun83]. Dot [BM87, DiV84, MBG77, LBS85, LJ90, Niz90, UB95]. Dot-Matrix [DiV84, LBS85, UB95]. dot/millimeter [LJ90]. Down [Bor96]. downsample [ACG94]. DP [Cat83]. Dr. [Hus76]. Draft [BDN83, Kit84]. Draft-To-Text [BDN83]. Drive [JSG78, FBJ90]. Druckkunst [Zap68, ZB72]. Drucktechnik [Deu98]. drukwerk [KVM02]. Dual [TA82]. Dual-Font [TA82]. Durham [Ano92]. Dwiggins [Fri83]. Dynamo [Bry84].

Early [Car02, LR94, Hus85a, She50, Lin52]. earn [Pos78]. Ease [Abi88]. eases [Nas94]. Easy [BF80, LU92, Wil89]. economic [Dav88, Hau94, Mor89]. Economical [Sac80]. economically [Lin39]. écriture [PAL16]. Ed. [Zap80]. EDIAN [KAT90]. Edit [MH70, TH73]. Edit-insertion [MH70]. edit/insert [TH73]. edited [Dur22]. Editing [Cla64, Kas72, LB74, Mas84, SOE77, Ran86, Com79]. Editor [Wal87, ANP81]. Editorial [GdMMP58, Swa71, Swa79, Swa86, Swa98].

Editors [CBB54, CBB65, CBB87, Dod97, McS84]. Education [Hel98, FM94, LBJ95]. Edward [Pec18]. effect [BSS92, Pow62, Ran86, WT28]. Effective [Gar73, Pin81, Pod84]. effectiveness [FZ93]. Effects [AMS83]. Effektivnosti [RF77]. Efficiency [RF77, WT26, WT33]. Efficient [Ghe87, IKM82, KB86, TC92]. efficiently [Lin39]. Einfuehrung [Sch86b]. Electric [Lab17]. Electronic [Ber75, Ber82, Bor85, Cle82, Clo80, Dun80, GS70, HAB98, Hol84, Mar87, Pie71, SP87, VC92, Wal84, Zey74, vV88, Kan87, Nat87a, Nat87b]. electrotyping [Fos81]. Element [WH86b]. Elements [Bri02, Bri12, Int06, LK95]. eleven [Kel10]. embedded [CK88, KKM^{+}93, KKM^{+}94]. employed [Ste19]. Encoding [PS84, SP87]. encyclopaedia [GMZZ94]. Encyclopedia [Sha77]. End [RS65, GSMN93]. enduring [FM94]. engineer [PEE58]. Engineering [CIS67, Kad11]. Engineers [Par75]. English [RPP96, BL84, Bli82, Cry15, Int06, MC75a, Ock71, Ock75]. engraving [Fos81]. engravings [Har83]. Enhancement [Bra09]. enjoy [Buh80, Buh81]. Entwicklungen [Zap07b]. enunciated [Hus76]. Environment [Hay87, Edv94, Wal82]. Envisioning [Tuf90]. EP [HAB98]. EP92 [VC92]. Eq [Kin91, Kin92]. eqn [Bra84]. equalization [ACG94]. Equations [CH85, Fod78]. Equipment [Nie80, PEE58, Gag29, Kro67, Phi68]. Erasable [Sta06]. Ergonomic [MRS83]. Ergonomics [Kil83]. Eric [Gil91]. errata [Knu86f, Knu89]. Error [Gra86, Ran86, VDM^{+}93]. Errors [AMS83, MC75b, ELS94, MC75a, San95, WW93]. Erste [WF99]. Esq [Bux88]. essay [Smi17]. Essays [Fle88, Dun83, Wal93]. essential [PR83, PR04]. establish [Jac74]. established [Bea88b]. esteem [FM94]. estimation [Kop93]. etc [Fos76]. European [Des86, Luc85, Zla92]. EuroTEX [Zla92]. evaluate [Rom75]. evaluation [DEM^{+}63]. every [Ros15]. Everyone [Hah93, Hah91]. Evidence [Tuf97]. evolution [Hau94]. examination [Win90, Bra84]. Example [SV73, GSKM92]. examples [Sch95]. Executive [Gar67]. exhibit [Ber23]. Experience [Bro69, CKT80, ZJ84]. Experimental [FC77]. Experiments [Sey66, Bar65]. Expert [BCJR87]. Explanations [Tuf97]. explanatory [Har83]. explicit [Spi80, Spi82]. Expressions [Hol87b, Pin94]. Extra [Bra13a]. extraction [MPZ95, TC92].

F.R.S. [Bux88]. Face

[Dow98, Jur02, Sha85, Ber23, IEE93]. faces [Ame30, Col30, Flo73, Lud89, Lan49]. Fachsprachlichen [Roe85]. facsimile [Gou78]. Factors [SDB88]. facts [BMD87, Bir80]. Fail [Ast72]. Fail-Safe [Ast72]. Failure [HL87]. failures [UBR90]. Falsified [TCLM82]. Family [DiV84]. famous [BMD87]. Fancy [Hof83]. Farewell [Mas84]. Fast [Hau22, BMD87]. Fatigue [WT26, WT28]. features [Fos81, HD88]. February [ZD91]. Federal [VC92]. feedback [Mar91a, Mar91b]. feedforward [MRT89]. Fehlerlose [Roe85]. Fewer [Pri81]. fifteen [Ber23]. Figures [BD87]. File [CSR10a]. Files [LMT^{+}21, Sac80, BD88, CSR10b]. Find [SG93]. findphrases [AB89]. fine [HK99, HK04, WT26, WT28, WT33]. Finer [Dow66, Dow93, Dow95]. Fingertips [Lun02]. finishing [Fos76]. First [Luc85, Ber23, Bli82]. Five [Ste96]. flow [FZ93]. Fluorescent [RF77]. FOAM [GW85]. Font [AMS83, AM86, Ano85a, Ano85c, Ano86b, Ano86c, Ano86d, Ano86e, Ano86g, Ano86h, Ast72, BDN83, BG73, BHW83, Bro85, CKM86, CK86, EM84, Fow73, FN85a, FN85b, GL84, Hof83, HD88, JSG78, KB86, KPB87, KH94, KH96, Kit84, KM87, LO90, LZ86, LH91, MBG77, McC84, MRS83, MN77, MS84, PS84, RS81, Rei66, SV73, SP87, SBSG89, SSST83, TKE83, TCLM82, Tri69, TA82, UKKS84, Ume82, UM85, Wad84, WH86b, Wol86, ZL87, Ano96b, Ano96e, Ano96f, BN94, BSS92, EEN90, FK85, Gho84, Guyxx, HFW^{+}96, KMNO89, KKM^{+}93, KKM^{+}94, Kop93, LK95, Lin95, LCO94, MP94, MPZ95, MRT89, NKMO89, Par93a, Ris88, SB93, Tow15, TC92, UBR90, WB95, HFW^{+}96]. Font-and [HD88]. Fonts [Col86, Hoe98, Mar87, ST85, BSS92, HBBG95, Int07, Pow90, WH76]. Forensics [BBK13a, BBK13b]. Form [CDM02, Buc64]. Formal [HYC82, ZA87]. Format [Hos24, But94]. Formats [Ben88, CSR10a, Bea88b]. Formatter [Fuj94, HO87, KV89, ZA87, AA89, Haa86]. Formatting [BF80, GW85, Geh87, Noo83, CHM00, FSS82, Geh86, GL88, Kin93, Com79, dG99]. Formeln [Sch86a, Thi83]. Formelsatz [Deu94]. Formelzeichen [Deu94]. Forms [CH85, Jac79, WD18]. Formulae [Fuj94]. Formulas [HO87, Haa86, Thi83]. Forth [For89, Smi89, Smi90]. FORTRAN [Her69, Her70, JK74, Her71]. Forty [Cha97, LJ90, Ber23]. Founding [WG19]. foundries [Ann75, Ann94]. Fourth [Ano91, Win90]. Frame [Mur19]. FrameMaker [Ben95, Mur19]. francais [Mac96]. France [Des86, HAB98, vV88]. Franklin [Anoxx, Lin52, Mil55, Sta06]. free [KD94]. French [And05, BL84, PAL16]. front [GSMN93]. Full [PRB87]. Fuller [Cos86]. function [KH96]. functional [dG99]. Functions [HL87]. Future [Ber82, Mor87, Opp85, MR92a].

gap [dG99]. Garamond [Ber23]. Gartom [RF77]. Gee [Wil37]. Gegenwart [Zap68, ZB72]. gender [Edv94]. general [Par93a, She50]. generalization [GSKM92]. generalized [Lag95]. Generating [Hob93, AB89, BS96, HFW^{+}96]. Generation [Kit84, PS84, ST85, UKKS84, Ano88, HBBG95, Kir05, LH91, MP94, MPZ95, Ris88, Win90, Won86a]. Generator [SSST83, THL80]. Generic [Aus88]. genesis [She50]. Geometric [FB85]. Georgia [Win90]. German [Thi83, Bau41, Deu98, Deu94, Zap68, ZB72]. Geschke [GS21]. Gestaltung [MB96]. Gestaltungen [Zap68, ZB72]. Ghost [Ano85c]. Giant [Poy01a]. Gill [Gil91]. given [VDM^{+}93]. Gks [SS86]. Gks-Graphiken [SS86]. Global [Hol06, KH94]. Glossary [Ano83a, Com66, Com69, EAE^{+}94, EAE^{+}96, Goo82, Nat87a, Nat87b, Com64, SS75]. Going [Hoy13]. Goodbye [Kir05]. Gospel [Gos90]. Goudy [Gou78, Hoy13]. Goudy’s [Ber23]. Gourmet [Spi90, Spi80, Spi82, Spi83, Spi86]. Grafica [Gar80]. Grafik [Blu86]. grammar [Int06]. grap [GL88, BK86, BK90]. Graphic [BDH^{+}94, BHH^{+}99, BDH02, Hel98, Hel99, Her78a, Her78b, Meg98, Mül96, SDB88, MB96, WH76]. Graphical [Mol93]. Graphics [CW82, Dun80, GRM97, Hoe98, RS81, Wal87, Bea88a, CK88, Ike95, Ker81, Ker82a, Ker82b, Van81a, Van80, Van81b, WW82]. Graphik [SS86]. Graphiken [SS86]. Graphs [BK90, BK86]. gratified [Ber23]. gray [Woo87]. grayscale [HBBG95]. Great [Bra13b]. Greater [Uni74]. Greek [Mac96]. greyvalue [HD88]. Grid [MB96, TR02]. gridfont [GSKM92]. Growing [WG19]. growth [Bli82]. Guide [Dod97, Fel03, Hah93, Lam86, Lam94, Spi90, Ano86f, Ano86i, Ano87a, Ano87c, Ano89a, Bar71, Ber75, BW95, Dal80, Fos76, Hah91, Hal80, Hal81, Hol84, KS87, KC78, Lev87, Lin82, Lip78, Mas05, McS84, Mye83, Mye84a, SS75, Spi80, Spi82, Spi83, Spi86, Szo82, Tex84, Uni78, Pri80, Wil89]. guidelines [Zap80]. Guild [Hod03]. Gutenberg [Fis00, Kir05, Ploxx]. GVIP [Ike95].

Hackers [Kad11]. half [Gou78]. half-century [Gou78]. Hallmark [Cus11]. Hamburg [WH86a]. hand [Lan49, Sax92, Ste19, Tay16, WT26, WT28, WT33]. Handbook [ABO06, Kra01, MW17, ERBR91, Hod03, Kle87, Kle90, Kro67, PR83, PR04, Ric81, vL92, Bau41]. Handbuch [BF10, Poh15, Bau41]. Handling [Ano86g, KM87, Mar87]. Handwritten [LZ86]. Hangul [Kim90]. Hansen [San99]. Happens [Wic23]. Hard [RF77]. Hardware [Boe76]. Hart [Har91]. Hashing [Hol87b]. having [Kra96]. HDTV [MRK93]. Head [Ano93b, FBJ90, LJ90]. Headlines [LB74]. hear [Tho21]. Hebrew [Har94b, Str93]. Height [Ano85b]. held [HAB98, Mac96, SPD72]. hello [Kir05]. Help [Nel88, Dow83]. Hermann [Cus11, Kel11, Kel19]. Hershey [WH76]. Heuristics [Sam82]. hidden [CWB93, HYH89, HYH91]. hierarchical [ANP81]. High [AIDG95, BHW83, CDFV12, DiV84, Int82, Int86, KMNO89, ST85, Ser88, UKKS84, ZL87, Ano80, Ano96b, Ike95, KKM^{+}93, KKM^{+}94, MP94, MR92a]. high-performance [Ike95]. High-Quality [DiV84, ST85, UKKS84, CDFV12, MP94]. High-Speed [ZL87, KKM^{+}93, KKM^{+}94]. High-tech [Int82, Int86]. Highest [MR92b]. Highlights [Ano97]. Highly [IKM82]. Hilfe [WF99]. hint [KKM^{+}93, KKM^{+}94]. Histoire [PAL16]. historical [Mor89]. History [Bra13a, Dow97, Dow98, MLS82, Meg98, Per88, Rom14, RM19, WH17, Hus85b, Kir16, Kle78, Kur17, PHM80, Smi17, PAL16]. Hitac [Hay87]. hoax [Hus76]. holy [Har94b, Har92]. home [BS88, Har83]. hood [Ano90c]. Hot [Mas84, MH98, San99]. hour [Wil89]. house [Ano52, Bir80, Bra84, Eis90, Hal80, Hal81, Mun88, Til52, Til55, Vác92, Gos90]. HPWord [Smi84]. Human [KB86, SDB88]. humanities [Ano87b]. hundred [Ste96]. hundredth [Fri83]. Hyphenate [Ock71, Ock75]. Hyphenating [RS65]. Hyphenation [MMN79, Smi84, Wan79]. Hyphens [Pri81]. hz [LK93]. hz-Program [LK93].

IBM [Gar67, Gar73, Pap89]. Iceland [Edv94]. ideas [Dow83, Kra96, War12]. Identification [Dow98, TCLM82, KH94, KH96]. IEC [SBSG89]. IEEE [Ano75]. IFIP [DS83]. II [MFB^{+}23]. III [Cra87a, RRLG83]. Illumination [WT28, WT26, WT33, WT26]. Illustrated [Dow98, Kle87, Kle90]. Illustrating [GRM97]. illustration [Hod03, Tex84]. illustrations [Spi80, Spi82]. Image [Ano88, IKM82, Poy01a, HD88, SO94]. Images [Tuf97, CWB93, Kop93, Lev84]. imaging [HAB98, WW82]. Impact [DiV84, ZL87, Mye84b, UB95]. Impaired [GL84]. Implement [BF80]. Implementation [Hol87a, AA89, Kin93]. Impressions [Ano97]. Improve [UBR90]. Improved [Buy85, Ano86i]. Improving [Cle82]. In-house [Bir80, Bra84, Eis90, Hal80, Hal81, Mun88]. Inc. [WG19]. Inclusion [CW82]. incorporating [Ano52, Phi68, Til52]. Incremental [CHM00, Hau22]. Independent [Bro85, IBC87, Ker82c, WW82]. indexes [AB89]. Indexing [Bra09, Col62, Cro80, Mul94, Mul05, Ste65, Ste70]. Indo [PC94]. Indo-Bangladeshi [PC94]. Industrial [For89, WT28, WT26]. Industrie [Sta71b, Sta71a]. industries [YKT90]. industry [Bir76, Bir83, Hau94, Kil83, Sta71b, Sta71a, Dep79]. Indx [AB89]. influential [Kel10]. Infoline [Opp85]. Information [Ano80, Ano83b, Ano87b, Bro85, CH85, ICC65a, II88, Lun99, Mar72, Opp85, Tuf83, Tuf90, Kan87, Mye83, Mye84a, Sli81, Sli84, SBSG89, Ste19, Myh90]. Initial [Ano86h]. Innovations [Par75]. Innovative [UB95]. inquiries [Sim87]. insert [TH73]. insertion [MH70]. Institut [Mac96]. Institute [VC92]. Institution [Hus85a]. instruction [Ten76]. Instructions [Ano95a, Ano95b, Ano95c, Ano98]. Instructor [Lin52, Har83]. Instrument [LR94]. Integrated [Gol73, GMZZ94]. Integration [App86, Blu86, Sch86a]. Intensive [Ben20]. Interactive [BW84, GM83, TBF^{+}77, Lau90, Par93b]. Interaktive [App86]. Interchange [SP87, Bea88b, SBSG89]. interesting [Fos81]. Interface [ASW89, GL84, LA84, Lau90, Phi68, LA85]. Interleaf [Blu86]. International [Ano92, HAB98, ICC65b, CTC66, ICC65a, Jas67, Lev95, VC92, WH86a, Mac96, Par90, PR83, PR04, SPD72, Ano14, vV88]. internationalization [Mac90]. Interpolating [FB85]. Interrobangs [Hou13a]. Intertype [Lud89]. Introduction [Dow97, Dow98, KR85, PHM80, Rad77, Rub88, Ano88, Com00, Grä96, Hen73, Pub71]. invariant [AAP94, HD88]. invention [Fre90, Rom86a, Rom86b]. inventions [Hus76]. inventor [Mer98, MS89]. Investigation [SDB88]. IPCC [IEE93]. iron [Sax92]. isn’t [Woo87]. ISO [SBSG89]. ISO/IEC [SBSG89]. isomath [Mil12]. ISSAC [Lev95]. issues [FSS82, Gho84]. italic [Int07]. Italics [Fow73]. Italy [Luc85]. items [Sli81, Sli84]. IV [Her69].

Jailbreak [Bra13b]. Japanese [Lun99, IBC87, ST85]. John [Lan40]. joint [WT26, WT28]. jointly [HAB98]. Josef [Mül96]. journal [MM73]. journalism [Ten76]. Journals [Ano75, LK77]. Joy [Spi90, Spi80, Spi82, Spi83, Spi86]. July [Ano93a, GR94, CTC66, Jas67, Lev95]. June [Abr81, Des86, For89, Mac96, USE87]. Just [Lan88]. Justification [Pri81]. Justified [Reh78, Sam82]. Justifying [Sam63, Ste19].

Kanji [FN85a, FN85b, SSST83, TKE83]. Kanji-Printers [FN85a]. Karel [KVM02]. Kathleen [San99]. keepsake [Bia93]. Kent [DS83]. Key [BM87]. Keyboard [Fri70, Lun02]. Khmer [Har94a]. Killer [SH19]. Klassifikatoren [SV73]. knowledge [TC92]. knowledge-based [TC92]. Knuth [Ano07]. Korean [Lun99, Kim90]. Kurs [GK00].

labels [SB93]. labours [Bux88]. Lamp [RF77]. Lamps [RF77]. Language [BK90, BW82, Fox87, Ges86, Pad87, SW92, BK86, CK88, Gau84, Ing91, Ker81, Ker82a, Ker82b, Kin93, PC94, Van81a, Van80, Van81b, WBS82]. languages [Zap68, Zap70, ZB72]. Lanston [Hop12]. Lara [Sch86a]. Large [Buy85, GL84, KM87, WH86b, Smi89]. Large-Area [WH86b]. Laser [Ano96d, KNS90, Nas94, FN87, LBS85, Mar91a, Mar91b, MRK93]. Laser-printer [Nas94]. Lasers [Dun80, Sch83]. late [Bux88]. Later [LMT^{+}21]. latest [Ben95]. LaTeX [MFW23, MFB^{+}23]. Latin [TA82]. Latin/Arabic [TA82]. Lausanne [VC92]. Law [Bor96, Hus76]. Laying [Bor96]. Layout [Kha00, SDB88, Ben95, Knu91c, Kro67]. Lead [RF77, Smi17]. leaded [Tay16]. Lear [Pec18]. Learning [CKM86, CK86, GH97]. Least [Kop93]. Least-squares [Kop93]. lecons [And05]. LePrint [LBS85]. lessons [And05]. Letter [WD18, Int06, Ros15]. Letterfontäne [Poh15]. Letterforms [Dur22]. lettering [Cus11, Lin82]. letterpresses [Ano19]. Letters [GF00, Tho21, Lin82, Mac96]. Level [GW85]. levels [Can93]. Libraries [Cis87]. Library [Buc64, EGL^{+}97, PRB87, ZD91]. Life [Fle88, Hou13b, Poy01a, Ano19, Bux88]. ligatures [Lag92b]. Lighting [RF77, WT28]. Like [Woh98]. Likelihood [BSS92]. Line [Hol06, RS65, Sam82, ZL87, KR85, Lan49, Mar91a, Mar91b, PK82, Ano07]. line-breaking [Ano07]. Linear [SV73, dG99]. linear-time [dG99]. Linearer [SV73]. Linee [Cos86]. lines [KP81]. lingual [Ano92, Gho84]. link [MH98, San99]. Linotron [CKT80]. Linotype [Lud89, Rom86a, Rom86b, Tay16, Col30, Mer98, MS89, Rom14]. Linsco [Linxx]. Linux [Lip99]. Liquid [Ano86d]. Lisa [Wag85]. List [NIS04]. listing [Lan49]. LITA [Cra87b]. Literacy [Hel99]. Literary [LR94, Kir16]. Literate [Knu84a, Sew89]. Load [Ano86h]. Loading [FN85a, FN85b]. local [DS83]. logon [McC88]. London [CTC65, SPD72]. Look [Gra75]. Looking [BDH^{+}94, BHH^{+}99, BDH02]. Lout [Kin91, Kin92, Kin93]. Low [BG73, Buy85, Col86, HH83, Thi91, Ser88]. Low-Cost [BG73, Buy85, Col86, HH83]. Low-tech [Thi91]. Lower [Mac75, Phi79]. LSI [KMNO89, NKMO89]. Ltd [Wil37]. Lyuminestsentnykh [RF77].

M [GS21, Zap80]. MACC [Ano87a]. machen [RPP96]. Machine [AM86, Rei66, Rom86a, Rom86b, Lan40, Buc64, Lan49, Ned89, Phi68, She50]. machinery [Sli81, Sli84]. machines [Fre90, Hus85a, Lin30, Lin39, Linxx, She50, Skr59, Wil31]. Machtig [RPP96]. Macintosh [KS87, CS86, How86, Knu91c, Mol93, SH19]. macro [Spi86, Spi90, WM88]. MACSYMA [Fod78]. Made [Wag85, Lan40, Ber23, Fos81, Lan49]. MAGIC [MPP81]. Magnetic [Kok87, Mye84a, Bro69]. Main [Dow98]. maintain [Lin39, Pos78]. Major [Blu64]. Make [LB74, Par90]. Makeup [KV89]. Making [Cry15, EGL^{+}97, Hos24, Mar83, Ste19]. making-up [Ste19]. Malo [HAB98]. man [Hoy13, Phi68]. man-machine [Phi68]. management [Rom75]. Manifesto [Sie86]. Manipulation [Abr81, VC92, Lab88, vV88]. Manual [Ano82a, Ano93c, Ano03, BK90, BF10, Dod97, Fel03, Gos90, Her78a, Her78b, Lab84, Lam86, Lam94, AC79, Ano90a, Bee86, Col62, Com83, Cra83b, Ker82a, MW05, MB96, Ros92b, Uni85b, Zap68, ZB72]. Manuale [BF10, Zap54, Zap70, Zap68, ZB72]. Manuel [BF10]. manufacture [Fos81]. manufactures [Ano84, Dep79]. manufacturing [Sou05]. Manuscripts [Ano95a, Ano95b, Ano95c, NIS98, NIS04, Ano98, Car80]. Manutius [Fle88]. Maple [Ano31, Lop94, Pin94]. Maps [GM83]. March [HAB98, CCT65]. March/April [HAB98]. Margination [Fis00, GF00]. mark [AJ88a, AJ90]. mark-up [AJ88a, AJ90]. Marken [RPP96]. marketing [Bar71]. markets [Rom76]. Markov [CWB93]. Marks [Hou13b, Deu94]. Marriage [Bra09]. Martens [KVM02]. Martin [San99]. Mass [Sey66]. Materials [SDB88, Ame30]. Math [Grä96, Grä00a, Grä00b]. Mathematica [Soi95]. Mathematical [CH85, Kas72, Kra96, Kra01, Mil12, MK89, Soi95, CS86, Hod88, Deu94, Smi17, Thi83, ISO09]. Mathematics [CBB54, CBB65, CBB87, Day83, KC75, Kor75, Pod84, Swa71, Swa79, Swa86, Swa98, Wil54, Wic65, FTBM96, FT96, KC78, KC80, Kin91, Kin92, Lop94]. mathematische [Deu94]. Mathematischer [Sch86a, Thi83]. maths [Tow15]. MathWriter [CS86]. matrices [Lan49]. Matrix [BM87, DiV84, MBG77, Hus85b, LBS85, Niz90, UB95]. matter [KVM02, Tay16]. matter/drukwerk [KVM02]. Matters [Dur22, WD18]. Maximizing [Ser88]. May [APL87, Luc85, Mor93]. MCS [Com83]. meaning [Thi83]. Means [Cle82, She50]. Mechanical [Bow47, Hus73, She50]. mechanism [WW93]. mecum [Lip99]. Media [Ano86e, Mye84a]. Meeting [WH17, Ano93a, GR94, Mor93]. Memoir [Bux88]. Memory [Buy85, Nas94]. Mergenthaler [CKT80, Guyxx, Mer98, MS89, Rom86a, Rom86b]. Merging [Gar80]. Merit [FM94]. Merrymount [Bia93]. Messages [PK02]. METAFONT [Knu87]. METAFONTbook [Knu86g, Knu86c]. Metal [Mas84, McG86, McG93]. Method [Ano85b, Ano85a, Ano86d, FN85a, FN85b, Ghe87, IKM82, PS84, RS65, Wad84, FBJ90, NSA90, Ploxx, TC92]. methods [Hus73, SO94, Sou05]. metric [Kop93]. metropolitan [Dep72]. Michael [Sey66]. Micro [LK93, Ano86i, BS88, CBT88]. micro-computer [BS88]. Micro-Typography [LK93]. Microcomputer [GW85, Kei83, Kle82]. Microcomputers [ZA87]. Microprogrammable [THL80]. MicroTEX [Ano86f]. millimeter [LJ90]. minutes [McC88]. Mode [Ano86h]. model [WW82, Zap80]. Models [Hus85a, KB86, CWB93]. Modern [Her78a, Her78b, LR94, Hol84, Knu86d]. Modular [Com83]. modulating [Ano96d]. modulation [FBJ90]. molecular [Ing91, All81, AKWC82]. MOLLY [Ing91]. Monotype [Ano97, Tay16, Lan40, Col30, Hop12, Lan49]. Monthly [Bro69]. Montreal [Lev95]. morphological [KD94]. Morris [Cos86, Tho77, Tho96]. most [Fos76]. Motor [JSG78]. Mr. [Ber23]. ms [GL88]. MT [Bro69]. MT/SC [Bro69]. Müller [Mül96]. Müller-Brockmann [Mül96]. Multi [AM86, Ano85a, Ano86e, Ano92, Ano96e, CKM86, CK86, Day83, EEN90, KM87, LZ86, Pad87, Ris88, SSST83, TKE83, Ume82, UM85, Ano96f, Gho84, Lin95, LCO94, TC92]. Multi-Access [Day83]. Multi-Font [Ano85a, Ano86e, CKM86, CK86, KM87, LZ86, SSST83, TKE83, Ume82, UM85, Ano96e, Ano96f, EEN90, Lin95, LCO94, TC92]. Multi-Language [Pad87]. Multi-lingual [Ano92, Gho84]. Multi-phase [EEN90]. Multi-system [Ris88]. multifont [Krt94]. multilingual [SZ91]. Multiple [Ano85c, Hol06, Met75, Flo73]. Museum [Ano97, WH17]. Music [Fox87, Lin90, CBT88, Con88]. Musicians [TBF^{+}77]. Must [Pod84]. Mut [GK00]. mv [GL88]. my [Lan88, Ploxx].

Name [GM83]. Narrative [Tuf97]. NASA [WM88]. National [Ano97]. natural [ISO09]. Nature [Pec18]. NCA [NCA81]. NCPA [NCP89]. necessary [Lam93]. needs [Cus11]. Neither [Big80]. Network [FN85a, FN85b, Ano96f, CDV95, LO90]. networked [KAT90]. networks [Ano96f, AIDG95, DS83]. Neural [CDV95, Ano96f, AIDG95, LO90]. neuron [Ano96f]. newly [Ano96f]. News [GS21]. Newsletter [Bea88a, Cra87b]. Newspaper [Dun80, IKM82, Kok87, Mas84, Eze87, Lag95, Pow62, Rob72, Win90]. Newspapers [MH98, Dep72, May69, Ran86, San99]. Nice [vV88]. Nicole [Dur22]. night [Ame30]. nineteenth [Skr59]. Niveaux [Can93]. no [VDM^{+}93]. non [Tay16]. non-technical [Tay16]. Nonfiction [PRB87]. Noninteractive [BW84]. nor [Big80]. Nordic [Edv94]. Normen [Deu94]. North [WH86a]. notation [Con88, Deu94]. Nov [SIG82]. novel [FBJ90]. NPP [BF80]. NR [Hol87a]. Nroff [Hol87a, Hol87b]. number [HYH89, HYH91]. Numerals [Hol87b]. numerous [Spi80, Spi82]. NY [Lop94].

Ob [RF77]. Obey [Poy01a]. object [LLW95]. object-oriented [LLW95]. Objective [Hol06]. occidental [WH76]. O’Collegian [Ten76]. O’Connor [Zap80]. OCR [ACG94, But94, CDV95, PC94, San95, SV73]. OCR-A [SV73, SV73]. October [ICC65b, IEE93, Knu89]. Office [Cat83, Ano80, YKT90]. Oklahoma [Ten76]. old [Ano19]. Omni [KB86, Wol86]. Omni-Font [KB86, Wol86]. on-line [KR85]. one [Lan49, Wil89]. one-line [Lan49]. Online [Opp85]. Only [CSR10a, EM84]. OpenType [Hos24]. operate [Buh80, Buh81]. operating [Jac74]. operation [Fri70]. operations [Ste19]. Operator [AMS83, MK89]. opportunities [Uni74]. Optical [AAP94, FTBM96, Mac75, Mar91a, Mar91b, Tri69, AIDG95, ELS94]. Optimal [FK85, Pla81]. options [Mun88]. optophone [Tho21]. Oregon [Abr81]. organizing [MRT89, WW93]. Oriental [LK95]. orientation [MRT89]. oriented [LLW95]. origin [Hop12, Hus85b, Phe61, She50]. original [Bli82]. origination [Fos76]. Origins [Lab17, War18]. ornaments [Ame30]. Osveshchenii [RF77]. Other [Blu64, Hou13a, Hou13b, Met75, Fos81, Gul62, Guyxx, Sim87, Ste19, TH73]. Ottmar [Mer98, MS89]. Our [CKT80, Cus11]. Outline [KKM^{+}93, KKM^{+}94, Ano96b, KMNO89, MP94, MPZ95, NKMO89]. outlines [Hob93]. Output [Bro85, KV89, Wat84, WT28]. overcomes [Mar91a, Mar91b]. Own [Clo80, Nie80]. Oxford [CBB54, CBB65, CBB87, Har91].

PA [IEE93]. Package [Cas91, Kin91, Kin92, Spi86, Spi90]. Page [Ale92, Ges86, KV89, Sha85, Ben95, Int06]. PageMaker [Ben95]. pages [Zap70]. Pagination [BKKW96, BKKW98, Pow86, Woh98, CDFV12, Lag95, Pla81]. paging [Kur17]. Pair [Bli82]. Panel [AMS83]. Paper [Ano80, Kur17, Fos76]. papers [CIS67, Car80, ZD91, Bra84]. papier [dBD02]. paragraph [Bir86, dG99]. paragraphs [KP81]. Paris [ICC65a]. Parlor [Dur22]. Parsing [Smi90, FTBM96]. Part [Com00, Deu98, ISO09, MFW23, MFB^{+}23]. parts [Linxx]. Password [BS88]. past [Mil55, Zap68, Zap70, ZB72]. paste [Jac79]. paste-up [Jac79, Tex84]. pasteup [Sac86]. Pattern [FN85a, FN85b, KB86, AAP94]. Patterns [Ano86c, PS84]. PC [Lev87, Pap89, Ros92b]. PCs [Zur85]. PEN [ANP81]. Penmanship [Int06]. people [IEE93, Tho21]. Perceived [SB93]. Perception [KB86, Dun83]. Perceptually [HBBG95]. Perfect [Fel03, Tow15]. Performance [AMS83, LCO94, Des89, Ike95]. Pergamon [Opp85]. peripherals [Phi68]. persnickety [Cry15]. Personal [Pow86, Ser84, Ano86f, Int82, Int86]. personnel [Jac74]. Perspective [MK89, Dav88]. Perspectives [Kin02]. Petites [And05]. phase [EEN90]. Philadelphia [IEE93, Hoy13]. Philological [Won86b]. Phoenix [USE87]. Photo [PT’65, Kle76]. photo-composition [Hen73]. Photocomposing [Zey74]. Photocomposition [AK73, Kor75, Mac75, Hol84, Pub71]. Photographic [Boa84, War85, War69]. Photoplotting [Ned89]. photos [Bea88a]. photoscript [EEN90]. photosetter [BS88]. Phototypesetter [CKT80, Kim80]. Phototypesetting [BJK87, Cra83b, Ham68, Rac70, Roh74, THL80, Tul85, Ber67, Gul62]. Phototypography [Hal80, Hal81]. PIC [Ker81, Ker82a, Ker82b, Ros92a]. Picture [Buy85]. Pioneer [Mül96, WH17]. pixel [UBR90]. pixels [Mac96]. Placement [CW82, GM83]. Places [PK02]. Plans [Opp85]. Plasma [AMS83]. Plass [Ano07]. Platemaking [Sta71b, Sta71a]. plates [Sli81, Sli84]. Plotting [Gar80]. point [Cry15, Fos76, Tay16]. Points [Ano86a, Dow66, Dow93, Dow95]. Poligrafii [RF77]. polyglot [SZ91]. Polygonal [MRK93]. Portable [WB95]. Portland [Abr81]. portraits [BMD87]. Possible [MC75a]. Postprocessing [KV89]. POSTSCRIPT [BD87, BD88, BW95, Ges86, GRM97, Per88, Pow90, War18]. Potomac [Lan71]. Pour [Ham75]. power [WBS82]. pp. [Sey66]. Practical [BW95, Lab84, Hol84]. practice [LBJ95]. Prague [Zla92]. Pre [Sie86]. Pre-TEX [Sie86]. premiums [Fos76]. prepaging [FK85]. Preparation [All81, AKWC82, Bar64, BW82, BW84, Bro69, Her70, Lam86, Lam94, Bar71, GB82, KL82, NCNS82, Sey82, WBS82, Zap80]. Preparatory [Her78a, Her78b]. prepared [Ano80]. Preparing [Ano93d, Uni85a]. prepress [KNS90]. Preprocessor [BF80, BD87, BD88, Dew86]. Present [Dow98, Mil55, Opp85, Zap68, Zap70, ZB72]. Presentation [II88, Ano86i]. presented [CIS67, ZD91]. Press [Ano31, CBB54, CBB65, CBB87, Har91, Sey66, Cav71a, Cav71b, Cav83, Sch95, Bia93, Dur22]. presses [Ano19, Gul62, Sax92]. pretty [Ros92a]. pri [RF77]. Pricing [Bir76, Bir83]. Primer [Kra96, FN87, Ste19]. principles [Col62, MTZ94]. Print [LR94, LJ90, Lev87]. Printed [AM86, Boa84, FC77, KPB87, LZ86, RS65, Sta06, Ume82, UM85, Ano96e, HK99, HK04, KVM02, Lin95, Tho21]. Printer [BG73, BM87, Cat83, DiV84, Kok87, Sou05, Ano52, Big80, FN87, Har83, Hus73, Hus85b, LBS85, Nas94, Til52, Til55]. Printers [BHW83, FN85a, FN85b, HH83, Sch83, ZL87, Ame30]. Printing [PP’30, PW’54, PEE58, Ano86b, Ano86a, Ano89b, Bli82, Blu64, Bra13a, CBB54, CBB65, CBB87, Dow97, Dow98, Dun80, Eis88, Fow73, Gib95, Gra86, MN77, Rad77, RF77, Ano95d, Ano19, Bar71, FBJ90, Fos81, Fos76, Lin82, Sli81, Sli84, Ste96, Uni78, Pri80, Wil31, War12, Zap68, Zap70, ZB72, Deu98]. printout [MH70]. Prints [Per88]. private [Cav71a, Cav71b, Cav83]. Problem [Sam82, Bir86, MRT89, Ter68]. Problems [Ste87, Des89, DEM^{+}63]. Procedural [BW82]. Procedure [MMN79]. procedures [Kun79, MAV82]. Proceedings [Abr81, Ano91, DS83, Dun80, ICC65b, Jas67, Lan71, Lev95, Luc85, Mac96, USE87, VC92, WH86a, Zla92, SIG82, APL87, CTC65, Des86, HAB98, CTC66, IEE93, SPD72, vV88]. Process [Ano91]. Processes [Eis88, Gul62, IEE93]. Processing [Abi88, Ano83b, CH85, DO87, Gol73, IBC87, Lun99, RRLG83, Sri93, BR87, Chr87, Des89, KKM^{+}93, KKM^{+}94, Kir16, LA84, LA85, Lip78, Mye84b, Rah87, Uni85a, Won90]. processing/typesetting [LA84, LA85]. Processor [Lab84, Ike95, KKM^{+}93, KKM^{+}94]. Producing [Pin81, Wil89]. Product [Pap89]. Production [Dun80, Gra86, HYC82, PP’30, Bra84, GSKM92, Mas05, Ten76]. productivity [LBJ95]. products [IEE93]. Produktion [Man87]. Professional [Hod88, KS87]. Professionals [Gel91]. professionell [Gor01]. profits [Pos78]. Program [Ano86h, Bar64, BJK87, Ben88, Gel91, Knu86a, Knu86b, Ock71, Ock75, Pie71, Sam63, Zey74, Dal80, Knu87, Knu88, Knu91a, Knu93a, Sew89, Thi91, TH73, dG99, LK93]. Programmed [Ano86b]. Programming [Her69, Her71, Bir86, Knu84a, Sew89]. programs [Guyxx, MH70]. Progress [FT96, Mar72, FM94]. Project [EGL^{+}97, Fis00]. projectors [MRK93]. promotions [Fos76]. proof [Ano52, Til52, Til55]. proofreading [Int06, Swa71, Swa79, Swa86, Swa98]. propagataion [HYH89]. propagation [HYH91]. proposal [Ber00]. Proposals [Roh74]. Prospects [Ber82, Opp85, Bar65, Sey66]. Prosthesis [GL84]. Providing [Ano85a]. Provisional [GM83]. Psfig [BD87, BD88]. pt [Sli81, Sli84]. Public [PRB87]. publication [Bar71, MM73]. Publications [Ben85, Gar80, Pin81, Ano93d]. publish [Bar71]. published [Kra96]. Publisher [Ber82, Gil89, Jan86]. publishers [Ame30, Gul62, Mas05, McS84, Mun88]. Publishing [Ano96c, Ber82, Bor85, Cis87, Dun80, Gos90, Gel91, Mar87, Met75, Mor87, Myh90, Rad77, RM19, Ulu86, VC92, WH17, vV88, Ano82b, Bee86, CDFV12, CBT88, Cra87a, FN87, HAB98, Kle87, Kle90, Nat87a, Nat87b, Nat86, Nel88, Ser88, Vác92, War12, YKT90, GK00]. Pulling [Dun80]. Pulse [FBJ90]. Punctuation [Hou13b, Cry15, Int06]. Punkt [Bau99]. Pursuit [MR92b]. Push [Blu64]. put [Ano80]. Pwn [CSR10a].

Quadratic [SV73]. quadratischer [SV73]. qualité [Can93]. Quality [Buy85, Can93, DiV84, MR92b, ST85, UKKS84, Ano96b, CDFV12, LBJ95, MP94, MR92a, SL67]. Quantitative [Tuf83]. Quantities [Tuf97, ISO09]. quantizer [Ano96f]. Quark [Ben95]. quarter [Wal84]. queuing [Com84]. Quick [Ghe87]. Quietwriter [Ano86g]. Quirky [Lun02]. quotations [Zap70]. Qur’an [Har92]. Qwerty [Lun02].

Rabot [RF77]. Raster [HAB98, WW82, MB96]. Raster-Systeme [MB96]. rasterizing [KMNO89, NKMO89]. rate [Ran86]. Ratgeber [WF99]. ratio [Som90]. Ray [Kim80]. Raznogo [RF77]. Re [Bra09]. Re-typesetting [Bra09]. Read [EM84, Kra96]. Read-Only [EM84]. readability [SB93, UBR90]. Reader [UM85]. Readers [CBB54, CBB65, CBB87, Har91, Kir05].

Reading [Cle82, BSS92, Pow62]. ready [Ano95a, Ano95b, Ano95c, Ano98, Zap80]. realities [CIS67]. really [Lam93]. recipes [LU92]. Recognition [AM86, Ano86e, CKM86, CK86, CW85, KB86, KPB87, LZ86, Rei66, Ste87, Tri69, Ume82, VP94, Wol86, All94, Ano96e, Ano96f, AAP94, AIDG95, EEN90, ELS94, FTBM96, HD88, KH96, KD94, Krt94, LO90, Lin95, LCO94, PR83, PR04, WW93]. recognizing [FT96]. recommendations [AJ90]. Reconsidered [BKKW96]. Record [Ano75, Buc64, MH98, San99]. recorded [Kra96]. REDUCE [ASW89]. Reduction [Gra86]. Reference [Hah93, Lam86, Lam94, ZJ84, Com83, Hah91, Lip78, Mac90]. reflected [Wil37]. refuses [Ano19]. Registers [Bra09]. rejecting [WW93]. related [Com64, Com66, Com69, Deu98, Wal93]. relation [WT26, WT33]. Relational [Lin95]. relative [BMD87]. reliable [Buc80]. religious [FM94]. reliure [Sta71b, Sta71a]. Remarkable [Pec18]. Rendering [Mor89, KKM^{+}93, KKM^{+}94]. repertory [WH76]. Report [Bor85, CTC65, Ste65, Ste70, Ben20, Int06, SS75, WT26, WT28]. Reports [Her70, NCP89, NCA81, WM88]. representation [Hus95, Lin95]. Request [Kan87]. required [Ano90b]. Requirements [Ste87, Wat84]. rescheduling [Ano89b]. Research [Ano83b, Blu64, Lin90, WT26, WT28, Car80, CIS67]. Resistive [Ano86a, LJ90]. Resolution [BHW83, MRK93, Ser88]. Resource [Bee86]. resources [Guyxx, Nel88]. Rest [Hos24]. Restoration [BBK13a, BBK13b]. Retaining [But94]. Reverse [Kad11]. Review [Dur22, San99, Sey66, Dav69, SL67]. Reviewing [NIS04]. Reviews [Pap89]. revision [Uni85b]. Revisited [Ano07]. Revisiting [BBK13a, BBK13b]. Reviving [Jur02]. Revolution [Blu64, Eis88, Ulu86, Ano82b, Ber75].

Rewards [FM94]. Rex [Anoxx]. Rex-Franklin [Anoxx]. Rhetoricity [Dur22, WD18]. Ribbon [Ano86a, LJ90]. Richard [Bli82]. RIDT [HAB98]. right [Szo82]. Ring [DS83, UCBB83]. RISC [KKM^{+}93, KKM^{+}94]. Rochester [For89]. Role [Ber82]. Roman [Ano92, Hol87b, Int07]. Rookledge [ERBR91, PR83, PR04]. Room [Cat83, Lin30]. Rotary [Ano95d]. rotation [AAP94]. Row [Ano85b]. RPO [IBM67]. Rules [CBB54, CBB65, CBB87, Har91, Jur02, Wic65, Ame30, BL84, Lan64, NIS04, Wan79]. RUNOFF [Bar64].

s [RF77]. Safe [Ast72, CSR10a]. saga [dBD02]. sale [Fos76]. sampling [AAP94]. sand [Sim87]. Santa [GR94]. Sättningsregler [Lan64]. Satz [Deu98, Thi83]. Saved [SH19]. Scalable [Ike95]. Scale [SOE77]. Scaling [Wad84, NSA90]. scan [Mar91a, Mar91b]. scan-line [Mar91a, Mar91b]. Scanned [Bra09, CWB93]. scanners [MRK93]. Scanning [Mac75]. SCAT [Kee80]. scholars [vL92]. Schrift [Zap68, ZB72, WF99]. Schriftart [SV73]. Schriftgrößen [Deu98]. Schriftsatz [Man87]. Schriftsetzer [Bau41]. Science [LR94, RS84, Mil12]. Sciences [Kra01, ISO09]. Scientific [Ano75, Her72, Luc85, NIS98, Des86, Hod03, Int07]. Scores [Fox87]. Scott [Dur22]. screen [Ano95d]. Screendesign [Kha00]. Script [Ano92, Kim90]. Searchable [ZJ84]. Seattle [Mor93]. second [Des86]. Secret [Hou13b]. secular [FM94]. Security [Sch83]. Segment [Ast72]. Segmentation [KD94, Ano96f, WW93]. Segmentation-free [KD94]. select [Gra66, Gra69]. selected [Nel88, Ran86, Wal93]. Selection [Ano86h, PR83, PR04, Ten76]. Selectric [Bro69]. Self [BN94, MRT89, WW93]. Self-correcting [BN94]. self-organizing [MRT89, WW93]. sell [Dow83, Rom76]. sells [RPP96]. Semantics [Pin94]. semi [Lau90]. semi-interactive [Lau90]. Seminar [Mor93, ZD91]. sense [Cra87a]. Sept [DS83]. September [Zla92, Knu86f]. series [She50, Dep79]. server [UCBB83]. service [Int06, Int82, Int86, KR85, Bur61]. Services [Opp85, SIG82]. Set [Ano86b, KM87, Ben95, Tay16, WM88]. SETAB [TH73]. Setting [Clo80, Fel03, Jus72, She50, Ste19, Wic65]. setzen [ICC65a]. Seven [Ast72]. Seven-Segment [Ast72]. Seybold [SS75]. Shady [Hou13a, Hou13b]. Shannon [San99]. shape [Par93a]. Share [Wat84]. Share-Output [Wat84]. Sheep [SG93]. Shift [Wic23]. Sholes [Rom86a, Rom86b]. shown [Ame30]. SI [Ano14, Ano14, NIS04]. SIC [Dep79]. Sichere [BN85]. signs [ISO09]. SIGOA [Abr81]. SIGPLAN [Abr81]. SIGUCCS [SIG82]. Silbentrennung [BN85, Roe85]. Simon [Sch95]. Simple [Dew86, War12]. Simplification [Gra86]. simulated [Lag95]. Simulation [MS84]. Single [Ber21]. Singular [Ano97]. Sinnentsprechende [BN85]. Sinotype [RRLG83]. Site [Ber00]. Size [KPB87, HD88]. sizes [BSS92, Deu98, Lan49, SB93]. Skeleton [BS96]. Skeleton-based [BS96]. sketch [Phe61]. Skill [Edv94]. skills [FN87]. Small [And05, Sac80, SOE77]. Small-Scale [SOE77]. SMI [KAT90]. SMI/EDIAN [KAT90]. Smithsonian [Hus85a]. Society [Par75, Cha97]. sociological [Eze87]. Software [Ano96c, Boe76, CW85, GL84, LBJ95, MNM79, Pad87, Ano86f, Ano98, Hod88, Ser88]. solid [Tay16]. Some [Ano90b, Skr59, AJ90, BS96]. sourcebook [Bry96]. SP [TA82]. Spacing [Dow66, Dow93, Dow95, Ste19]. spatial [MRK93]. SPEC [CK88, Whi87]. Special [Wil31, Mac90]. specific [SO94]. specification [Car80, WB95]. specify [Rom75]. Specimen [Lan40, Ano52, Bur61, Bli82, Til52]. specimens [Ame30, Gou78, Lan49]. Spectral [RF77]. spectroscopic [TH73]. speech [Ano96e]. Speed [Cle82, ZL87, KMNO89, KKM^{+}93, KKM^{+}94, Pow62]. Speedier [Blu64]. speeds [MM73]. Spektra [RF77]. spelling [MC75a]. Spent [CKT80]. Sperry [Ano87a]. spite [UBR90]. spline [LH91]. sponsored [CIS67]. spotting [CWB93]. Sprache [BN85]. Sprachen [Zap68, ZB72]. Spreadsheet [Ano85b]. Spreadsheets [Ano85a]. squares [Kop93]. St [HAB98]. Stages [Dow98]. Standard [Fow73, Hos24, SP87, Tri69, Ano52, SBSG89, Til52, Til55, Ber23]. Standards [Hos24, MR92b, Deu94, AJ88a, AJ90]. Stanford [McC20]. Star [Lin30]. start [Buh80, Buh81]. State [Bor85, Ste65, Ste70, Ten76, SL67, Ham68]. State-of-the-Art [Ste65, Ste70, SL67]. statements [Zap68, ZB72]. Stealing [SG93]. Stop [SG93]. Store [EM84]. Stories [Zap07a]. Story [Lor19, Lun02, Cry15, Gou78, Hoy13, Rom86a, Rom86b, Ros15]. Strangers [LMT^{+}21, CSR10b]. Strasbourg [Des86]. Strategies [Hoe98, Bir76, Bir83, FZ93]. Strike [Rob72]. Strike-on [Rob72]. strings [Smi90]. stroke [Ano96b, LK95, MPZ95, TC92]. Strokes [TKE83]. Structural [Fuj94, CDV95]. Structure [BJK87, HO87, Haa86, Ing91, MR92a]. Structured [Ano96f, BW82, Noo83, Won90, LK95]. Structures [AKWC82, All81]. Student [Her78a]. Studies [WH86a, Fle88]. Study [FZ93, FC77, Lin90, Mil55, PRB87, Kel10, Le 88, Phi68, SZ91]. Style [Ano31, Ano82a, Ano93c, Ano03, Bri02, Bri12, Dod97, MTZ94, MW17, WH86b, Aus88, Ano52, Ano90a, Car80, Mil12, NIS04, Til52, Til55]. Styles [AMS83, Pow62, SB93]. subject [Ano95d]. subjects [Wal93]. Substitution [McC84]. successful [Buh80, Buh81]. Summary [Dow98, Sli81, Sli84]. Summer [BBK13a, BBK13b, CKT80, USE87, Lop94]. Supplement [Com69, Sey84a]. supplies [Lin30]. Supply [Le 88]. supporting [Kim90]. survey [FSS82, Phi68]. Sussex [CTC66, Jas67]. Sweden [Edv94]. Swedish [Lan64]. swift [BMD87]. Swiss [VC92, Mül96]. Switzerland [VC92]. Symbol [Ano86b]. Symbolic [Lev95]. symbology [UB95]. Symbols [Hou13b, NIS98, Int07, ISO09, WH76, Deu94]. Symposium [Abr81, Lev95, Lop94, Mac96]. Syntactic [Wol86]. Synthese [ZK88]. Synthesis [TKE83]. System [Ano83b, Ano14, Boe76, Com83, GW85, Geh87, GS70, Her70, Her71, Her72, IBM67, IBC87, KC75, Lam86, Lam94, MN77, MS84, Rac70, Rub88, SOE77, SDB88, THL80, Tul85, UKKS84, Wol86, Zey74, AB89, Ano86f, Ano96b, Ano96d, Ano96e, AAP94, Buc80, Chr87, Geh86, GL88, GMZZ94, HFW^{+}96, JK74, Kan87, KAT90, KC80, Kle79, KNS90, Lag92a, Lam93, Lau90, LLW95, MM90, Mye83, Mye84a, Par93b, Ris88, Rom75, Smi89, Ten76, Tut83, Com79, Kee80]. Systematic [Ric78]. Systeme [MB96]. Systems [BCJR87, Col86, Day83, Gar80, Gol73, Mar87, Pow86, Ste87, WG19, Zey74, Cla64, Des89, EP87, FSS82, Hol84, MB96, NCNS82, Pla81, Sey82, WT28, Ano80].

T [Sey66]. table [VDM^{+}93]. Tables [Dew86, Ghe87, MM90, TH73, WH76]. tablets [Mac96]. tabulating [Int06]. Take [Zur85, CSR10b]. talks [She50]. tally [Mor73]. Tape [Bro69, Gar73]. tariff [Sli81, Sli84]. TBFLP [Dew86]. tbl [Bra84]. Teach [Phe89]. Teacher [Her78b, FM94]. tech [Int82, Int86, Thi91]. Technical [AK73, Ben85, Boe76, ICC65a, Jan86, LK77, MPP81, May69, Mor87, Pod84, SH94, SDB88, Wri81, Zap07a, ABO06, Bea77, Edv94, IEE93, MC75a, Spi80, Spi82, Spi83, Tay16]. Techniken [Deu98]. technique [Ham75, Kro67, Par93a]. Techniques [CKM86, CK86, Lin90, Ano87b, Cla64, Deu98, Jac79, LCO94, Pla81, SL67, WH76]. Technischen [Ulu86]. technischer [Zap07b]. technological [Hau94, Wal84]. technologies [Gib95]. Technology [BM87, LR94, Mas84, RS84, VC92, Ano87b, Ano89b, Ano96d, Cos86, Deu98, DS83, Eze87, ISO09, Lab88, Mil12, Mor89, PHM80, Ran86, Mur19]. Teil [Deu98]. Telecommunicating [Mye83]. Teleprinter [TA82]. tells [Ros15]. temporal [MRK93]. ten [McC88, LMT^{+}21]. Tenth [SIG82]. Terminal [GL84]. Terminology [Wal81, Wal83]. Terms [EAE^{+}94, EAE^{+}96, Ano83a, Com64, Com66, Com69, Fos76, Goo82, Nat87a, Nat87b, Deu94]. Terrestrial [Her79]. testing [FZ93]. TEX [Des86]. TEX-embedded [CK88]. TEX-REDUCE [ASW89]. TEX-Texten [SS86]. TEX/LaTEX [Haa86, Fuj94, HO87]. Texas [APL87]. TEXbook [Knu84b, Knu91b, Knu93b]. TEXnical [Myh90]. TEXnology [Pap89]. Text [Abi88, Abr81, AK73, BF80, BR87, BDN83, Blu86, CW85, CSR10a, DO87, Dun80, FC77, Fuj94, GW85, GB82, Gol73, HO87, Hos24, KV89, LB74, MNM79, Noo83, PRB87, RRLG83, Sac80, Sam82, SDB88, Com79, ZA87, AA89, Ano86i, Chr87, EEN90, Haa86, Hod88, Int07, Niz90, San95, Spi80, Spi82, Spi83, Tay16]. Text-Editing [LB74]. Text-Graphic [SDB88]. Text-Only [CSR10a]. Textbook [ZJ84]. Texte [SS86]. Texten [SS86, Roe85, ICC65a]. textes [ICC65a]. Texts [AM86, HM01, Kee80, Kin02, FF93]. TEXTURES [Ano87c, Knu91c, Ale92]. Textverarbeitung [ZK88]. Textverarbeitungsumgebung [App86].

Thai [Str93]. their [Ann75, Ann94]. theological [Won86a]. theory [LBJ95]. There [Jus72, Tow15]. thermal [FBJ90, LJ90]. Things [Cle82, PK02]. third [Won86a]. those [Wil31]. three [MB96, Pow62]. tight [Eis90]. Tillotsons [Ano52, Til55]. time [dG99]. Timetable [SOE77]. tipografia [BF10]. tipografico [BF10]. title [Int06, VDM^{+}93]. title-page [Int06]. TJ [Sam63]. TJ-2 [Sam63]. Today [Ham68, Lab88, Mye84b]. Together [Dun80, BMD87]. tokens [FM94]. Tolbert [Hop12]. tolerances [MRK93]. Tones [Tho21]. tool [UB95]. tools [Ano87b, Wil31]. top [Nat86]. Topography [Doy93]. Total [Wri81]. Touch [HFW^{+}96, Kle78]. Tour [Mol93]. Tpsnew [Pow90]. Track [Kir16]. trade [Phe61, Sta71b, Sta71a, Sli81, Sli84]. tradeoffs [Som90]. Train [SOE77, Jac74]. Training [Gel91, Ben20, LBJ95]. Transformational [Bir86]. transition [APL87]. translation [AAP94]. Transporting [Hay87]. traps [WBS82]. Trend [Wil37]. Tribute [Fri83]. Troff [EP87, GL88, Dic87, Ker82c, IBC87, Bra84]. Troy [Lop94]. True [Her95]. TSUS [Sli81, Sli84]. Tube [Kim80]. tuned [HBBG95, Hob93]. TURES [KS87]. Turning [Tho21]. Tutorial [BK90, Flo73, Hah93, Hah91, Lev87]. TV [MRK93]. twentieth [Kel10, Sou05, McG86, McG93]. Two [GW85]. Two-Level [GW85]. Type [Ann75, Ann94, BH02, Col30, CBM99, Deu98, Dow66, Dow93, Dow95, Dow98, Fel03, HM01, Lab84, Lab88, Lie77, Mil55, Rub88, Sam63, SG74, SG93, Swa71, Swa79, Swa86, Swa98, Til52, TCLM82, Tri69, Wil54, Wag85, Whi87, Wic65, WD18, Ame30, Ano52, Bur61, Bli82, Col89, Cra83a, Cus11, Flo73, Gho84, Gou78, Jus72, Kro67, Lan49, Lan88, MH98, Mor89, Par93b, RPP96, San99, She50, Sou05, Ste19, Szo82, WH76, Dur22, WW03].

Type-setting [Wic65]. typedesigners [ERBR91]. typeface [Cos86, PR83, PR04, WH95]. Typefaces [Knu86d, NIS98, BS96, Lie77, Lin82, McG86, McG93]. typefinder [PR83, PR04]. Typen [RPP96]. Types [Dow97, Dow98, Lan40, Ber21, Ber23, Kel11, Mor73, Zap70]. typescripts [Zap80]. Typeset [Lab84, FTBM96, FT96, Kra96, McC88, Ros92a, San95, Pow90, Uni85b]. Typesetter [BBK13a, BBK13b, Ker82c, Lab17]. Typesetters [Bau41, Lab84]. Typesetting [Dep72, Ame83, Ame84, Ame88, AKWC82, Ano83c, Ano86i, Ano95a, Ano95b, Ano95c, Bur61, Bar87, BE92, Ben88, BK90, Ber23, Blu64, Boe76, Bra13a, Col30, CTC65, Cas91, CH85, Cle82, Clo80, Col86, Con88, CW82, Deu98, Dav88, Day83, Dic87, Doo89, Doy93, EAE^{+}94, EAE^{+}96, Eze87, FF93, Fod78, Fox87, Fuj94, Gos90, Gag29, Gar67, Gar80, Geh87, Gel91, GSMN93, Ghe87, Gil89, Haa86, HO87, Hah93, Har92, Har94b, Har94a, Hau22, Her78a, Her78b, Hoh88, How86, IBM67, CCT65, ICC65b, Jas67, Kel85, KC75, KC78, Kle82, Knu79, Knu86e, Lag92b, LK93, LK77, Lev84, LB74, MAV82, Mar72, MM90, MR92b, Mol93, MK89, Mye84a, Nie80, Phe89, Pie71, Pin94, Pow90, Reh78, Ros92b, RRLG83, Sac86, Sac80, SW92, Ser84, Sey66, Sie86, Sma79, SH94, Soi95]. Typesetting [Spi90, Sri93, Ste19, Str93, Tex84, Thi83, Tut83, T’770, VDM^{+}93, Wan79, Wal82, Won86b, Woo87, Zur85, vdL92, von67, Aus88, AC79, All81, Ano52, PT’65, Ano80, Ano82b, Ano83a, Ano86f, Ano88, Ano89a, Ano96d, Ano98, Anoxx, AJ88b, AJ88a, AJ90, BS88, BMD87, Bar65, Bar71, BR87, BL84, Bea77, Bea88a, Bee86, Lan40, Ben95, Ben20, BK86, Ber75, Ber67, Big80, Bir76, Bir80, Bir83, Bow47, Bra09, Buc80, Buc64, Buh80, Buh81, BW95, Can93, Com64, Com66, CIS67, Com69, Car80, CK88, Cla64, Com83, Com00, CS86, Cos86, Cra87a, Cra87b, Dal80, Dav69, Dow83, DEM^{+}63, Edv94, Eis90, EP87, FN87, Fos81, Fos76, Fre90, Fri70, Gar73, Gau84, Geh86, GL88, Gho84, GB82, Gol69, Goo82, Gra66, Gra69, GMZZ94]. typesetting [Gul62, MLS82, Hah91, Hal80, Hal81, Har83, Hau94, Hen73, Hol84, Hop12, Hus73, Hus85a, Int06, CTC66, ICC65a, Ing91, Int82, Int86, Jac74, Jac79, JK74, Jan86, Jus72, Kan87, KS87, KR85, KC80, Ker81, Ker82a, Ker82b, Kil83, Kin91, Kin92, Kle76, Kle78, Kle79, Kle87, Kle90, Knu86f, Knu89, Knu91c, Kro67, Kun79, LBS85, Lan49, Lin30, Lin39, Linxx, LA84, LA85, Lab88, Lag92a, Lam93, Lau90, Le 88, LLW95, Lev87, Lin82, Lip78, Mar91a, Mar91b, May69, McS84, MH70, MR92a, Mun88, Mye83, Mye84b, Myh90, Nat82, Nat87a, NCA81, Nat87b, Nat86, Ned89, Neg88, PHM80, Pub71, Pap91, Phe61, Phi68, Pla81, Ploxx, Pos78, Pow62, Ran86, Ric81, Rob72, RM74, Rom75, Rom76, Rom86a, Rom86b, Sta71b, Sta71a, SZ91, SS75, Sey82, Sey84b, Sey84a]. typesetting [Sha77, She50, Skr59, SPD72, Sli81, Sli84, Spi80, Spi82, Spi83, Spi86, SL67, Til52, Til55, Tay16, Ten76, Ter68, Thi91, TH73, Com84, Uni78, Pri80, Com79, Uni74, Dep79, Ano93d, Uni85a, UCBB83, Vác92, Van81a, Van80, Van81b, Wil31, War85, War69, Wal84, Wal88, Wal93, WT26, WT28, WT33, Wil89, Wil37, Win90, WM88, WH76, Won86a, Kee80, Lan64, HH83]. typesetting/computer/communications [Goo82]. Typesetting/illustration/paste [Tex84]. Typesetting/illustration/pasteup/audio-visual/guide [Tex84]. typesetting/photo [Hen73]. typesetting/photo-composition [Hen73]. typesetting/photocomposition [Pub71]. Typewriter [MM73, Rom86a, Rom86b]. typewriters [Gar67]. Typical [Zey74]. typists [Zap80]. Typo [Kha00, MC75a]. Typografie [BF10, Poh15, WF99]. typographers [Ame30, MB96].

Typographic [BCJR87, Bri02, Bri12, CDM02, FC77, Her70, Her71, LR94, MW17, Ame30, PAL16, Phi68, SL67, Zap70]. typographic-quality [SL67]. Typographica [Poy01b]. Typographical [Hou13a, Hou13b, MC75b, Wal81, Wal83, Hus76, Ran86, Zap68, ZB72]. typographicum [Zap54, Zap68, Zap70, ZB72]. Typographie [Aic84, Aic89, Gor01, GK00, Zap68, ZB72, And05, Ham75]. typographique [BF10, PAL16]. typographische [Bau99, Zap68, ZB72]. Typography [BH02, BD83, Bor96, Car02, CBM99, Fel03, HM01, Her69, Her72, Her81, Her95, Her78a, Her78b, Jur02, Kei83, Kin02, Knu99, Kra01, Lab17, Lab88, LK93, McC20, MR92b, Rub88, SZ91, Sin98, Typ02, VC92, WW03, Zap68, ZB72, And05, Anoxx, Ber00, BF10, Bry96, Com00, Cra83a, Dun83, Fre90, Gou78, HAB98, MW05, Nat87a, Nat87b, PHM80, Smi17, Wil37, ZD91, vV88, WW03]. Typomania [Wal93]. Typophiles [Gou78]. TYPSET [Dal80].

U.K [Dav69]. UCSB [Com79]. UK [DS83]. ultimative [Poh15]. Umgang [WF99]. Unbound [Hoe98]. Understanding [Hel99, Kle76]. Unicode [Hos24]. Unit [NIS04]. Units [Ano14, HYH89, HYH91, ISO09]. Univ [DS83]. University [For89, CTC66, Jas67, Gul62, CTC65, CBB54, CBB65, CBB87, Har91, Ten76]. UNIX [BR87, Chr87, DO87, EP87, Geh86, Geh87, GL88, IBC87, KL82, LK77, MM90, Sri93, Tut83, Com79]. Unjustified [Kin02, Reh78]. up/audio [Tex84]. upheaval [Wal84]. Upon [AMS83]. USA [ICC65b, USE87]. Use [CSR10a, IBC87, Met75, Sie86, Zur85, Flo73, Gul62, Int07, Par93a, She50, WW82]. used [Fos76, ISO09, Lan49, Lin82, Ros92a]. User [SIG82, Ano87c, BK90, Dal80, Ham68, Lam86, Lam94, Ano87a, KS87, KC78, Ker82a, Lau90, Mye83, Mye84a, Ros92b, Uni85b]. Users [PRB87, Ano86i, Int82, Int86]. Uses [II88]. Using [ACG94, BCJR87, CW85, CBT88, Fow73, Gau84, Hah93, KB86, SV73, SH94, WW93, Wri83, Ano96d, Ano96e, Ano96f, Ano98, AIDG95, Bra84, Car80, CWB93, Con88, CDV95, Dic87, Hah91, How86, KH94, Lab88, Lag95, LLW95, MP94, Pow90, SO94, Com79, WH95]. USP [ZJ84].

v [RF77, Lop94]. Vacation

[BBK13a, BBK13b, CKT80]. vade [Lip99]. Valley [Lan71]. Vapors [Boa84]. Variable [Ano85c, Rei66]. Variable-Width [Ano85c]. varied [SB93]. varies [HYH89, HYH91]. variety [Col30, HFW^{+}96]. Various [Guyxx, RF77, She50]. Vax [Ano87a]. Vein [Ale92]. Ventura [Lor19, Wil89]. Vergangenheit [Zap68, ZB72]. Vergleich [SV73]. Verheissung [Won86a]. verschiedenen [Zap68, ZB72]. version [Ano86f, Spi80]. versions [Ben95]. Versprechen [Won86a]. versus [MRK93]. vertical [Int06]. verwandte [Deu98]. via [vdL92]. Video [RS81]. Vietnamese [Lun99]. View [Car02, Doo89, Gil89, Opp85, AJ88a, Jan86, Mor89]. viewed [YKT90]. Viewing [BW84]. Viewpoint [Ham68]. Vinyas [Par93b]. Virtual [GL84, Gou78]. Visible [Le 88]. Visual [Tuf83, Tuf97, KH94, MB96]. visual/guide [Tex84]. Visually [Ben85, GL84]. visuelle [MB96]. vital [CIS67]. VMS [Ano87a]. Vol [GL88]. Volume [Dun80]. Volumes [AKWC82, All81]. vom [Kha00]. Vorbereitung [Man87]. VOS3 [Hay87]. vowels [Lag92b]. VTEX [Cas91].

W. [Fri83]. WA [Mor93]. warning [SB93]. warnings [BSS92]. Washington [ICC65b, Lan71, CCT65]. Waterloo [Buc80]. Watson [Wil37]. Way [Cle82, Ploxx]. Weaving [Sew89]. WEB [Sew89, Sin98]. weigh [Bir80]. Well [Bea88b]. Well-established [Bea88b]. West [Ber23]. WG [DS83]. WhatTheFont [Anoxx]. Wheel [BG73]. where [BL84]. wherein [Ame30]. which [HYH89, HYH91, Hus76, Lan49]. White [Ano80]. Who [Ano83c, Nat82]. Wider [Sie86]. widow [Big80]. Width [Ano85b, Ano85c]. widths [Guyxx]. wife [Big80]. William [Hus76, Bli82, Tho77, Tho96]. wiped [Ano19]. Wire [BM87]. Within [Ano85a]. wobble [Mar91a, Mar91b]. women [Big80]. wood [Fos81]. Word [CWB93, IBC87, Lab84, Mye84b, Des89, KH96, Kir16, LA84, LA85, Tho21, Uni85a]. Word-Processing [IBC87]. Worde [Cha97]. Words [Ock71, Ock75, SG74, MC75a, Rah87]. Wordworking [Cra87a]. Work [Fle88, Gra86, WT26, WT28, WT33]. working [Ste19]. Works [RF77, SG93]. workshop [DS83, Lop94, Lan71]. Workstation [KAT90, Wri83, Gho84]. Workstation-based [KAT90]. Workstations [KM87]. World [Hos24, Poy01a, PW’54, Ano19, Kel19, Sey84b, Sey84a, Str93]. Wp [Pin81]. writers [Kir05]. Writing [Ano91, Ben85, Kra96, Lin52, Sta06, Wri81, ABO06, Int06, PAL16, Rom86a, Rom86b, Zap68, ZB72]. Writings [BDH^{+}94, BHH^{+}99, BDH02, HM01]. wrought [Hoy13]. Wyatt [Dur22]. Wynkyn [Cha97]. WYSIWYG [LLW95].

XIII [Gou78]. XIV [Gou78]. XPress [Ben95]. XXXII [WH86a]. XY [Ros92b]. XY-pc [Ros92b]. Xy-pic [Ros92a].

Years [Ano96a, Cha97, LMT^{+}21, Ano19, Ste96].

Zapf [Cus11, Kel11, Kel19]. Zeichen [Deu94]. Zeilenfall [Kha00]. zum [Kha00]. zur [GK00, Poh15]. zweier [ZK88].

References

Abi-Akar:1989:ATF

[AA89] Ramez Abi-Akar. Arabic text formatter: comments and implementation. Computer Languages, 14(1):53–60, 1989. CO-DEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Asselin:1994:OCS

[AAP94] Daniel Asselin, Henri-H. Arsenault, and Donald Prevost. Optical circular sampling system for translation and rotation invariant pattern recognition. Optics Communications, 110(5-6): 507–513, September 1994. CO-DEN OPCOB8. ISSN 0030-4018.

Abe:1989:IFS

[AB89] Kris K. Abe and Daniel M. Berry. Indx and findphrases, a system for generating indexes for ditroff documents. Software —Practice and Experience, 19 (1):1–34, January 1989. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abikoff:1988:TEA

[Abi88] William Abikoff. TEX: The ease and art of text processing. Abacus, 5(4):10–29, summer 1988. CODEN ABACEJ. ISSN 0724-

Alred:2006:HTW

[ABO06] Gerald J. Alred, Charles T.

Brusaw, and Walter E. Oliu. Handbook of technical writing. Bedford/St. Martins, Boston, MA, USA, eighth edition, 2006. ISBN 0-312-43613-0 (paperback), 0-312-35267-0 (hardcover). xxiv + 636 pp. LCCN T11 .B78 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0708/2005921350-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0708/
2005921350-d.html.

Abrahams:1981:PAS

[Abr81] P. Abrahams, editor. Proceedings of the ACM SIGPLAN SIGOA Symposium on Text Manipulation, Portland, Oregon, June 8–10, 1981, volume 16(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1981. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6.

Adams:1979:CTM

[AC79] R. C. Adams and G. A. Cleave.

CICERO typesetting manual. ARL-STRUC-NOTE; 453 Aeronautical Research Laboratories, Fishermens Bend, Victoria. Structures note; 453. Aeronautical Research Laboratories, Melbourne, Victoria, Australia, 1979. ISBN 0-642-90048-5. 128 pp. LCCN ????

Agazzi:1994:UOE

[ACG94] O. E. Agazzi, K. W. Church, and W. A. Gale. Using OCR and equalization to downsample documents. Proceedings — International Conference on Pattern Recognition, 2:305–309, 1994. CODEN PICREG. ISSN 1051-4651. IEEE catalog number 94CH3440-5.

Aicher:1984:T

[Aic84] Otl Aicher. Typographie. Ernst, Berlin, Germany, 1984. ISBN 3-433-02090-6. 256 pp. LCCN Z246 .A45 1988. In German and English. With a contribution by Josef Rommen.

Aicher:1989:T

[Aic89] Otl Aicher. Typographie. Ernst, Berlin, Germany, 2. Auflage edition, 1989. ISBN 3-433-02090-6. 256 pp. LCCN ???? In German and English. With a contribution by Josef Rommen.

Avi-Itzhak:1995:HAO

[AIDG95] Hadar I. Avi-Itzhak, Thanh A. Diep, and Harry Garland. High accuracy optical character recognition using neural networks with centroid dithering. IEEE Transactions on Pattern Analysis and Machine Intelligence, 17(2):218–224, February 1995. CODEN ITPIDJ. ISSN 0162-8828.

Arrabito:1988:CBTb

[AJ88a] R. Arrabito and H. Jurgensen. Computerized Braille typesetting: another view of markup standards. Electronic Publishing—Origination, Dissemination, and Design, 1(2):117– 131, September 1988. CODEN EPODEU. ISSN 0894-3982.

Arrabito:1988:CBTa

[AJ88b] R. Arrabito and Helmut Jurgensen. Computerized Braille typesetting. Report, Dept. of Computer Science, University of Western Ontario, London, Ontario, Canada, 1988. ISBN 0-7714-0910-9. 11 pp.

Arrabito:1990:CBT

[AJ90] R. Arrabito and H. (Helmut) Jurgensen. Computerized Braille typesetting: some recommendations on mark-up and Braille standards. Thesis (M.Sc.), Dept. of Computer Science, University of Western Ontario, London, ON, Canada, 1990. vii + 88 pp.

Alt:1973:CPT

[AK73] Franz L. Alt and Judith Yuni Kirk. Computer photocomposition of technical text. Communications of the ACM, 16 (6):386–391, June 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Describes the typesetting system used by the American Institute of Physics.

Allen:1982:CCD

[AKWC82] Frank H. Allen, Olga Kennard, David G. Watson, and Kathleen M. Crennell. Cambridge Crystallographic Data Center — 6. preparation and computer typesetting of “Molecular structures and dimensions” bibliographic volumes. J. Chem. Inf.

Comput. Sci, 22(3):129–139 (or 129–138??), August 1982. CO-DEN JCISD8. ISSN 0095-2338.

Alexander:1992:TTP

[Ale92] George Alexander. TEX and TEXtures: Page composition in a completely different vein. The Seybold Report on Desktop Publishing, 7(1):3–7, September 4 1992. CODEN SRDFED. ISSN 0889-9762.

Allen:1981:PCT

[All81] Frank H. Allen. The Preparation and computer typesetting of the “Molecular structures and dimensions” bibliographic volumes for the Cambridge Crystallographic Data Centre. Science and Engineering Research Council, Rutherford Appleton Lab., Chilton, Eng., 1981. 21 + [17] pp.

Allam:1994:ACR

[All94] May Allam. Arabic character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 2181:351–359, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Amin:1986:MRM

[AM86] Adnan Amin and Gerald Masini. Machine recognition of multi font printed Arabic texts. Proceedings — International Conference on Pattern Recognition, pages 392–395, 1986. CODEN PICREG. ISBN 0-8186-0742-4.

IEEE Service Cent. Piscataway, NJ, USA.

ATC:1930:BSB

[Ame30] American Typesetting Corporation. The book of specimens: the book of the American Typesetting Corporation, typographers to printers, publishers and advertisers, wherein are shown type faces, rules, ornaments, borders and typographic materials to be had of American typographers, both by day and by night. The Corporation, Chicago, IL, USA, 1930. 270 pp.

AEA:1983:TB

[Ame83] American Entrepreneurs Association, Los Angeles, CA, USA. Typesetting business, 1983. 119 pp.

AEA:1984:TB

[Ame84] American Entrepreneurs’ Association. Typesetting business. AEA business manual; no. X1245. Entrepreneur Magazine, Los Angeles, CA, USA, 1984. 119 pp.

AEA:1988:TB

[Ame88] American Entrepreneur Association, Los Angeles, CA, USA. Typesetting business, 1988. 191 pp.

Abramson:1983:EDE

[AMS83] Sandra R. Abramson, L. Hardy Mason, and Harry L. Snyder. Effects of display errors and font styles upon operator performance with a plasma panel. Proceedings of the Human Factors Society, 1:28–32, 1983. CODEN PHFSDQ.

Andre:2005:PLT

[And05] Jacques André. Petites leçons de typographie. (French) [Small lessons of typography]. IRISA, Rennes, France, May 16, 2005.

50 pp.  URL http://www.
irisa.fr/faqtypo/lessons.

pdf. Revision of earlier versions of 1990 and 2003.

Annenberg:1975:TFA

[Ann75]  Maurice Annenberg.  Type

foundries of America and their catalogs. Maran Print. Services, Baltimore, MD, USA, 1975. ISBN ???? 245 pp. LCCN Z250.A2 A55.

Annenberg:1994:TFA

[Ann94]  Maurice Annenberg.  Type

foundries of America and their catalogs. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, second edition, 1994. ISBN 1-884718-06-X. xxi + 286 pp. LCCN Z250.A1 A56 1994. US$49.95. With additions and an introduction by Stephen O. Saxe, and an index by Elizabeth K. Lieberman.

Anonymous:1931:MPC

[Ano31] Anonymous. The Maple Press Company Style Book. Maple Press Company, York, PA, USA, 1931. ???? pp. LCCN ???? Reprinted 1942.

Anonymous:1952:TTS

[Ano52] Anonymous. Tillotsons type specimen book, incorporating the house style of typesetting and standard printers’ and authors’ proof correction. Tillotsons, Bolton England, 1952. 133 pp.

Anonymous:1975:ICS

[Ano75] Anonymous. IEEE Conference on Scientific Journals, Record, 1975. IEEE Transactions on Professional Communication, PC-18(3):??, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Anonymous:1980:HPC

[Ano80] Anonymous. How to put a cap on the high costs of office typesetting / a White Paper prepared by Browne Information Systems. Browne Information Systems. Marketing Department, New York, NY, USA, 1980. 18 pp.

Anonymous:1982:CMS

[Ano82a] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 13th edition, 1982. ISBN 0-226-10390-0. ix + 737 pp. LCCN Z253 .U69 1982. UK£30.

Anonymous:1982:CTR

[Ano82b] Anonymous. Computerized typesetting: a revolution in publishing, 1982. 1 sound cassette (ca. 26 min.).

Anonymous:1983:GTT

[Ano83a] Anonymous. Glossary of typesetting terms. Compugraphic Corp., Wilmington, MA, USA, 1983. 25 pp.

Anonymous:1983:RDC

[Ano83b] Anonymous. Research and Development of a Chinese Information Processing System. Instrument Society of America, Research Triangle Park, NC, USA, 1983. ISBN 0-87664-739-5. 271– 278 pp. LCCN TA165.M52 1983.

Anonymous:1983:WWT

[Ano83c] Anonymous. Who’s Who in Typesetting, 1983. National Composition Association, Arlington, VA., 1983. 54 pp.

Anonymous:1984:CM

[Ano84] Anonymous. 1982 census of manufactures. Report MC 82-I-27 C-4 (P), U.S. Dept. of Commerce, Bureau of the Census: For sale by Supt. of Docs., U.S. G.P.O., Washington, DC, USA, 1984. 6 pp.

Anonymous:1985:MPM

[Ano85a] Anonymous. Method for providing multi-font characters within spreadsheets. IBM Technical Disclosure Bulletin, 28(2):823– ??, July 1985. CODEN IBM-TAA. ISSN 0018-8689.

Anonymous:1985:MAA

[Ano85b] Anonymous. Method to automatically adjust row height and column width of a spreadsheet.

IBM Technical Disclosure Bulletin, 28(2):819–??, July 1985. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1985:VGC

[Ano85c]  Anonymous.  Variable-width

ghost cursor for multiple font characters. IBM Technical Disclosure Bulletin, 27(11):6510–??, April 1985. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:APR

[Ano86a] Anonymous. All-points-addressable printing with a resistive ribbon. IBM Technical Disclosure Bulletin, 29(2):609–610, July 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:ACP

[Ano86b] Anonymous. Automatic changing of programmed symbol set according to printing font. IBM Technical Disclosure Bulletin, 29 (4):1682–??, September 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:CFP

[Ano86c]  Anonymous.  Compression/

decompression of font patterns. IBM Technical Disclosure Bulletin, 28(8):3563–3564, January 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:LCF

[Ano86d] Anonymous. Liquid crystal font addressing method. IBM Technical Disclosure Bulletin, 29(2): 807–808, July 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:MMC

[Ano86e] Anonymous. Media and multifont character recognition. IBM Technical Disclosure Bulletin, 29 (4):1562–??, September 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:MPT

[Ano86f] Anonymous. MicroTEX personal typesetting software, version 1.5A1: system guide. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-17207-0. iv + 56 pp. LCCN ????

Anonymous:1986:QFC

[Ano86g] Anonymous. Quietwriter font cartridge handling. IBM Technical Disclosure Bulletin, 28(11): 5062–??, April 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:SDM

[Ano86h] Anonymous. Selection of the display mode, display font, and display buffer by initial program load command. IBM Technical Disclosure Bulletin, 29(4):1524– 1525, September 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:TMU

[Ano86i] Anonymous. Typesetting for micro users: a beginner’s guide to improved text presentation. Quorum Technical Services, Cheltenham, 1986. ISBN 0-948726-00-8 (spiral). 248 pp. LCCN ????

Anonymous:1987:TAM

[Ano87a] Anonymous. TEX at MACC: user guide for the VMS Vax and Sperry 1100. Typesetting series. University of Wisconsin-Madison, MACC Academic Computing Center, Madison, WI, USA, 1987. 56 pp.

Anonymous:1987:ITH

[Ano87b] Anonymous. Information technology in the humanities: tools, techniques and applications. Computers and their applications. Halsted Press, New York, USA, 1987. 184 pp.

Anonymous:1987:UGT

[Ano87c] Anonymous. User’s guide, TEX-TURES. Addison-Wesley, Reading, MA, USA, 1987. 123 pp.

Anonymous:1988:IGI

[Ano88] Anonymous. Image generation an introduction to typesetting, 1988. 1 videocassette (20 min.).

Anonymous:1989:BGT

[Ano89a] Buyers’ guide to typesetting, page various, 1989. Benn Publications, Tonbridge, KY, USA.

Anonymous:1989:PTR

[Ano89b] Anonymous. Printing technology and rescheduling. Folio: the Magazine for Magazine Management, 18(10):187–??, October 1, 1989. ISSN 0046-4333.

Anonymous:1990:ASM

[Ano90a] Anonymous. AIP style manual. American Institute of Physics, Woodbury, NY, USA, fourth edition, 1990. ISBN 0-88318-642-X (paperback). 64 pp.

LCCN QC5.45 .A45; QC28 .A5 1990; T11 .A45 1990.

Anonymous:1990:SAR

[Ano90b] Anonymous. Some assembly required. BYTE Magazine, 15 (8):297–300, August 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:UH

[Ano90c] Anonymous. Under the hood. BYTE Magazine, 15(8):289–??, August 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PFA

[Ano91] Anonymous, editor. Proceedings of the Fourth Annual Conference on Computers and the Writing Process. Univ. Sussex, Brighton, UK, March 1991.

Anonymous:1992:MCA

[Ano92]  Anonymous, editor.  Multi-

lingual computing: Arabic and Roman Script: 3rd International conference — December 1992, Durham. University of Durham, Durham, UK, 1992.

Anonymous:1993:AMJ

[Ano93a] Anonymous, editor. 14th Annual meeting: — July 1993, Birmingham, volume 14(3) of TUGboat. TEX Users Group, San Francisco, CA, USA, 1993. ISSN 0896-3207.

Anonymous:1993:AHC

[Ano93b] Anonymous. At the head of the class. In-plant Reproductions, 43 (8):14–??, August 1, 1993. ISSN 0198-9065.

Anonymous:1993:CMS

[Ano93c]  Anonymous.  The Chicago

Manual of Style. University of Chicago Press, Chicago, IL, USA, 14th edition, 1993. ISBN 0-226-10389-7. ix + 921 pp. LCCN Z253 .U69 1993. US$40.00.

USGAO:1993:PPT

[Ano93d] Anonymous. Preparing publications for typesetting. Report GAO/OIMC-12.14.1, Office of Information Management and Communications, United States. General Accounting Office, Washington, DC, USA, 1993. ???? pp.

Anonymous:1995:ITCa

[Ano95a] Anonymous. Instructions for typesetting camera-ready manuscripts. International Journal of High Speed Computing, 7(2):??, ???? 1995. CODEN IHSCEZ. ISSN 0129-0533.

Anonymous:1995:ITCb

[Ano95b] Anonymous. Instructions for typesetting camera-ready manuscripts. International Journal of High Speed Computing, 7(3):??, ???? 1995. CODEN IHSCEZ. ISSN 0129-0533.

Anonymous:1995:ITCc

[Ano95c] Anonymous. Instructions for typesetting camera-ready manuscripts. International Journal of High Speed Computing, 7(4):??, ???? 1995. CODEN IHSCEZ. ISSN 0129-0533.

Anonymous:1995:RSP

[Ano95d] Anonymous. Rotary screen printing subject to change. International Textile Bulletin, Dyeing/Printing/Finishing (English Edition), 41:50–52, 3rd Quarter 1995. CODEN ITBDD2. ISSN 1012-8417.

Anonymous:1996:YAD

[Ano96a] Anonymous. 50, 100, and 150 years ago: December 1846. Scientific American, 275(6):14, December 1996. Describes the new (in 1846) Clay and Rosenborg steam type setting machine; it required one man and four boys to work. An illustration of the machine is provided.

Anonymous:1996:CSH

[Ano96b] Anonymous. Conversion system of high quality Chinese characters from stroke font to outline font. Ruan Jian Xue Bao/Journal of Software, 7(5): 257–263, May 1996. CODEN RUXUEW. ISSN 1000-9825.

Anonymous:1996:DPS

[Ano96c] Anonymous. Desktop publishing software. Home office computing: building better business, 14 (1):74–??, January 1, 1996. ISSN 0899-7373.

Anonymous:1996:LTS

[Ano96d] Anonymous. Laser typesetting system using acoustooptic modulating technology. Yadian Yu Shengguang/Piezoelectrics and Acoustooptics, 18(4):235–239, August 1996. CODEN YASHE7.

Anonymous:1996:MRP

[Ano96e] Anonymous. Multi-font recognition of printed Arabic using the BBN Byblos speech recognition system. ICASSP, IEEE International Conference on Acoustics, Speech and Signal Processing — Proceedings, 4:2136–2139, 1996. CODEN IPRODJ. ISSN 0736-7791. IEEE catalog number 96CB35903.

Anonymous:1996:SNN

[Ano96f] Anonymous. Structured neural networks for multi-font Chinese character recognition using a newly developed digital neural network chip with adaptive segmentation of quantizer neuron architecture (ASQA). Neural Networks for Signal Processing — Proceedings of the IEEE Workshop, pages 330–336, 1996. IEEE catalog number 96TH8205.

Anonymous:1997:HCS

[Ano97] Anonymous. Highlights continued: “Singular Impressions: The Monotype in America,” National Museum of American Art. Smithsonian, 28(2):30–??, May 1997. CODEN SMSNA5. ISSN 0037-7333.

Anonymous:1998:ITC

[Ano98] Anonymous. Instructions for typesetting camera-ready manuscripts using computer software. Parallel Processing Letters, 8 (2):269–273, June 1998. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Anonymous:19xx:RTA

[Anoxx] Anonymous. Rex-Franklin typesetting: advertising typography. Rex-Franklin typesetting, ????, 19xx. various pp.

Anonymous:2003:CMS

[Ano03] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 15th edition, 2003. ISBN 0-226-10403-6. xvii + 956 pp. LCCN Z253 .U69 2003. US$55.00.

Anonymous:2007:KPL

[Ano07] Anonymous. Knuth & Plass line-breaking revisited. Web site, July 23, 2007. URL

                                        https://defoe.sourceforge.
                                        net/folio/knuth-plass.html.

Anonymous:2014:SBI

[Ano14]  Anonymous.  SI Brochure:

The International System of Units (SI). Bureau International des Poids et Mesures (BIPM), F-92312 Sèvres Cedex, France, eighth edition, 2014. ISBN 92-822-2213-6. 186 pp. LCCN QC90.8 .S99 8th ed.

                                        (2006). URL https://www.
                                        bipm.org/en/publications/
                                        si-brochure/;   https://www.
                                        bipm.org/utils/common/pdf/
                                        si_brochure_8.pdf. Issued in

2006, and updated in 2014.

Anonymous:2019:HWO

[Ano19] Anonymous. How the world’s old printing presses are being brought back to life: Digital printing almost wiped out the world’s letterpresses 40 years ago, but the art refuses to die. The Economist, ??(??):22–25, December 19, 2019. CODEN EONOEH. ISSN 0013-0613 (print), 1476-8860 (electronic).

URL https://www.economist.
com/christmas-specials/2019/
12/18/how-the-worlds-old-
printing-presses-are-being-
brought-back-to-life.

Anonymous:20xx:W

[Anoxx]  Anonymous.  WhatTheFont.

Web site, 20xx. URL http://

www.myfonts.com/WhatTheFont/

.

Allen:1981:PHD

[ANP81] Todd Allen, Robert Nix, and Alan Perlis. PEN: a hierarchical document editor. ACM SIGPLAN Notices, 16(6):74–81, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

APLC:1987:CPA

[APL87] Conference proceedings: APL 87 Dallas: APL in transition, May 10–14, 1987, Dallas, Texas. ACM Press, New York, NY, USA, 1987.

Appelt:1986:ITI

[App86] W. Appelt. Die Integration von TEX in eine Interaktive Textverarbeitungsumgebung. Informationstechnik it: Computer, Systeme, Anwendungen, 28(6):342– 347, 1986. CODEN ICSAES.

Astle:1972:FFS

[Ast72] Brian Astle. Fail-safe font of seven-segment digits. Electronics, 45(18):93–94, August 1972. CODEN ELECAD. ISSN 0883-4989.

Antweiler:1989:TRI

[ASW89] Werner Antweiler, Andreas Strotmann, and Volker Winkelmann. A TEX-REDUCE interface. SIGSAM Bulletin, 23 (2):26–33, April 1989. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

AGPS:1988:GTC

[Aus88] Australian Government Publishing Service, Canberra, ACT, Australia. Generic typesetting codes: style book, 1988. various pp. Three volumes.

Barnes:1964:RPP

[Bar64] Larry Barnes. RUNOFF: A program for the preparation of documents. Memo CC-244 / MAC-M-193, Project MAC, Massachusetts Institute of Technology, Cambridge, MA, USA, December 6, 1964. URL http:

                                        //web.mit.edu/Saltzer/www/
                                        publications/CC-244.html.

Barnett:1965:CTE

[Bar65]  Michael P. Barnett.  Com-

puter typesetting; experiments and prospects. MIT Press, Cambridge, MA, USA, 1965. xvi + 245 pp. LCCN Z253 .B29.

Barnett:1971:HPG

[Bar71] Mitzi Barnett. How to publish: a guide for copy preparation, typesetting, printing, binding, and marketing your publication. Miran Publishers, Ft. Worth, Tex. (246–250 Ridglea Bank Bldg., 3327 Winthrop St., Ft. Worth 76116), 1971. i + 15 pp. Revised July 1973.

Barlow:1987:TC

[Bar87] Geoff Barlow. Typesetting and composition. Publisher’s guide series. Blueprint, London, UK, 1987. ISBN 0-948905-04-2. xiv + 270 pp. LCCN Z253.3.B35 1987.

Bauer:1941:HSG

[Bau41] Friedrich Bauer. Handbuch für Schriftsetzer (German) [Handbook for Typesetters]. Klimschs graphische Bücherei. Verlag von Klimsch & Co., Frankfurt/Main, Germany, tenth edition, 1941. viii + 312 pp.

Bauer:1999:TP

[Bau99] Friedrich L. Bauer. Der typographische Punkt. Informatik Spektrum, 22(1):41–42, February 1999.

Bagley:2013:RSVa

[BBK13a] Steven R. Bagley, David F.

Brailsford, and Brian W. Kernighan. Revisiting a summer vacation: Digital restoration and typesetter forensics.

Report, School of Computer Science, University of Nottingham, and Department of Computer Science, Princeton University, Nottingham NG8 1BB, UK and Princeton, NJ 08540, USA, July

20, 2013. URL http://www.
cs.princeton.edu/~bwk/202/.

See also the original and reconstructed papers [CKT80].

Bagley:2013:RSVb

[BBK13b] Steven R. Bagley, David F. Brailsford, and Brian W. Kernighan. Revisiting a summer vacation: Digital restoration and typesetter forensics. In Proceedings of the 2013 ACM Symposium on Document Engineering, pages 3–12. ACM Press, New York, NY, USA, 2013. URL

https://dl.acm.org/doi/10.
1145/2494266.2494275.

Biles:1987:UES

[BCJR87] Al Biles, Frank Cost, Guy Johnson, and Ken Reek. Using expert systems in typographic design. IEEE Transactions on Professional Communication, PC-30(2):102–111, June 1987. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Bigelow:1983:DT

[BD83] Charles Bigelow and Donald Day. Digital typography. Scientific American, 249(2):??, August 1983. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Batchelder:1987:PDP

[BD87] Ned Batchelder and Trevor Darrell. Psfig — A DITROFF preprocessor for POSTSCRIPT figures. In USENIX Association [USE87], pages 31–42.

Batchelder:1988:PDP

[BD88] N. Batchelder and Trevor Darrell. Psfig — a Ditroff preprocessor for PostScript files. Internal report, Computer and Information Science Department, University of Pennsylvania, Philadelphia, PA, USA, 1988.

Bierut:1994:LCC

[BDH^{+}94] Michael Bierut, William Drenttel, Steven Heller, D. K. Holland, and Bill Drenttel, editors. Looking Closer: Critical Writings on Graphic Design. Allworth Press, New York, NY, USA, 1994. ISBN 1-880559-15-3 (vol. 1), 1-880559-56-0 (vol. 2). 256 (est.) pp. LCCN NC997.L63 1994. US$18.95. Two volumes.

Bierut:2002:LCC

[BDH02] Michael Bierut, William Drenttel, and Steven Heller, editors. Looking Closer 4: Critical Writings on Graphic Design. Allworth Press, New York, NY, USA, 2002. ISBN 1-58115-235-3. xi + 291 pp. LCCN NC997.L6333 2002. US$21.95.

Berkebile:1983:DFC

[BDN83] H. E. Berkebile, J. M. Dunn, and L. R. Nielsen. Draft-to-text font conversion algorithm.

IBM Technical Disclosure Bulletin, 25(10):5173–5174, March 1983. CODEN IBMTAA. ISSN 0018-8689.

Barlow:1992:TC

[BE92] Geoff Barlow and Simon Eccles. Typesetting and Composition. Blueprint, London, UK, second edition, 1992. ISBN 0-948905-72-7. xiv + 274 pp. LCCN ????

Beach:1977:CTT

[Bea77] Richard J. Beach. Computerized typesetting of technical documents. Technical report, Dept. of Computer Science, University of Waterloo, Waterloo, ON, Canada, 1977. 10 pp.

Beach:1988:NTG

[Bea88a] Mark Beach. Newsletter typesetting graphics and photos. Coast to Coast, Portland, OR, USA, 1988. 36 pp.

DeLaBeaujardiere:1988:WDI

[Bea88b] Jean-M De La Beaujardiere. Well-established document interchange formats. In J. C. van Vliet, editor, Document manipulation and typography: Proceedings of the International Conference on Electronic Publishing, Cambridge series on electronic publishing, pages 83–94 (of vi + 288). Cambridge University Press, New York, NY, USA, 1988. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Beechick:1986:RMT

[Bee86] Al Beechick. Resource manual for typesetting with your computer: alternative to desktop publishing. Arrow Connection, Pollock Pines, CA (P.O. Box 899, Pollock Pines 95726), second edition, 1986. 46 pp.

Benham:1920:ITT

[Ben20] Charles Edward Benham. Intensive training in typesetting: a report. Maqua Co., Schenectady, NY, USA, 1920. 62 pp.

Benson:1985:WVD

[Ben85] Philippa J. Benson. Writing visually: Design considerations in technical publications. Technical Communication, 32(4):35– 39, 4th Quarter 1985. CODEN TLCMBT. ISSN 0049-3155.

Bentley:1988:DPT

[Ben88] J. L. Bentley. DFORMAT — a program for typesetting data formats. Computing Science Technical Report 142, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1988. 14

pp. URL https://github.com/
arnoldrobbins/dformat. 142.

Benedek:1995:AIS

[Ben95] A. Benedek. Adobe is set to dominate typesetting and page layout: the latest versions of PageMaker, quark XPress and FrameMaker. The British Printer: a journal for the printing trades, 108(12):6O, ???? 1995. ISSN 0007-1684.

Berry:1921:ST

[Ber21] Edward D. Berry. Single types. Technical report, ????, East Aurora, NY, USA, 1921. 16 pp.

Berry:1923:STC

[Ber23] Edward D. Berry. Standard Typesetting Company is gratified to exhibit first in the West Mr. Goudy’s beautiful types made after a face credited to Claude Garamond in fifteen forty. ????, Chicago, IL, USA, 1923. 8 pp.

Bernstein:1967:CAT

[Ber67]  Hans H. Bernstein.  Com-

puter aided typesetting and phototypesetting. Joint Nuclear Research Center, Ispra Establishment, Scientific Information Processing Center — CETIS, Ispra, Italy, 1967. 122 pp. LCCN Z253.3 .B4 Oversize. EUR 3661e. At head of title: European Atomic Energy Community–EURATOM.

Berg:1975:ECG

[Ber75] N. Edward Berg. Electronic composition: a guide to the revolution in typesetting. Graphic Arts Technical Foundation, Pittsburgh, PA, USA, 1975. ISBN 0-88362-011-1. 374 pp. LCCN Z253.3 .B44 Special Coll.

Berul:1982:PFE

[Ber82] Lawrence H. Berul. Prospects for the future of electronic publishing: the role of the publisher. International Online Information Meeting, pages 207– 217, 1982. CODEN IOIMDT. ISBN 0-904933-39-3.

Berleant:2000:SDT

[Ber00]  Daniel Berleant.  On site:

does typography affect proposal assessment? Communications of the ACM, 43(8): 24–25, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/cacm/
2000-43-8/p24-berleant/.

Barach:1980:NEI

[BF80] David R. Barach and David M.

Fram. NPP: An easy to implement preprocessor for text formatting. Software—Practice and Experience, 10(5):335–346, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bodoni:2010:MTM

[BF10] Giambattista Bodoni and Stephan Füssel. Manuale tipografico: Manual of typography: Handbuch der Typografie: Manuel typographique: Manual de tipografia. Taschen, Köln, Germany, 2010. ISBN 3-8365-0553-3 (hardcover). lxxii + 267 (vol. 1), 279 (vol. 2) pp. LCCN Z232.B66 A23 2010.

Branson:1973:CWL

[BG73] C. A. Branson and J. B. Gillett.

Character wheel for low-cost changeable font printer. IBM Technical Disclosure Bulletin, 16 (5):??, October 1973. CODEN IBMTAA. ISSN 0018-8689.

Baines:2002:TT

[BH02] Phil Baines and Andrew Haslam. Type & Typography. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 2002. ISBN 0-8230-5524-8. 192 pp. LCCN ???? US$29.95.

Bierut:1999:LCC

[BHH^{+}99] Michael Bierut, Jessica Helfand, Steven Heller, Rick Poynor, and Michael Beirut. Looking Closer 3: Classic Writings on Graphic Design (Looking Closer). Allworth Press, New York, NY, USA, 1999. ISBN 1-58115-022-9. xix + 283 pp. LCCN NC997.L633 1999. US$18.95.

Bresenham:1983:FCH

[BHW83] J. E. Bresenham, J. S. Heath, and C. N. Wallis. Font compression in high resolution printers. IBM Technical Disclosure Bulletin, 25(9):4582–4584, February 1983. CODEN IBMTAA. ISSN 0018-8689.

Bianchi:1993:MPC

[Bia93] Daniel Berkeley Bianchi. The Merrymount Press: a centenary keepsake. Stinehour Press, Bridgewater, CT, USA, 1993. ISBN ???? 24 + 55 pp. LCCN Z232.M57 B5 1993. US$200.00.

Biggs:1980:NPW

[Big80] Mary Biggs. Neither printer’s wife nor widow: American women in typesetting, 1830– 1950. Library Quarterly, 50(4): 431–452, 1980.

Birkett:1976:PST

[Bir76]  William A. Birkett.  Pric-

ing strategies for the typesetting industry. National Composition Association, Arlington, VA, USA, 1976. 42 pp.

Birkett:1980:ITW

[Bir80] William H. Birkett. In-house typesetting: weigh the facts. The Association, Arlington, VA, USA, 1980. 22 pp.

Birkett:1983:PST

[Bir83]  William A. Birkett.  Pric-

ing strategies for the typesetting industry. National Composition Association, Arlington, VA, USA, 1983. 42 pp.

Bird:1986:TPP

[Bir86] R. S. Bird. Transformational programming and the paragraph problem. Science of Computer Programming, 6(2):159– 189, March 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bentley:1987:CPP

[BJK87] Jon L. Bentley, Lynn W. Jelinski, and Brian W. Kernighan. CHEM — a program for phototypesetting chemical structure diagrams. Computers and Chemistry, 11(4):281–297, 1987. CODEN COCHDK. ISSN 0097-8485.

Bentley:1986:GLT

[BK86] Jon L. Bentley and Brian W.

Kernighan. GRAP: a language for typesetting graphs. Communications of the ACM, 29(8): 782–792, August 1986. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1990:GLT

[BK90] J. L. Bentley and B. W. Kernighan. GRAP — a language for typesetting graphs. tutorial and user manual. Unix Research System Papers, II:109– 146, 1990.

Bruggemann-Klein:1996:PR

[BKKW96] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. Pagination reconsidered. Technical Report 205, Fernuniversität Hagen, Germany, 1996.

Bruggemann-Klein:1998:PCD

[BKKW98] Anne Brüggemann-Klein, Rolf Klein, and Stefan Wohlfeil. On the pagination of complex documents. Technical Report 234, Fernuniversität Hagen, Germany, July 27, 1998.

Baudot:1984:BRT

[BL84] Alain Baudot and Therese Lior. Basic rules for typesetting in French: where they differ from rules for English. Technical report, Groupe de recherche en études francophones (G.R.E.F.), Toronto, ON, Canada, 1984. ISBN 0-919604-96-X. v + 10 pp.

Bliss:1982:PPR

[Bli82] Carey Bliss. A Pair on Printing: Richard Atkyns’: The original and growth of printing; William Caslon: the first English type specimen book. Bird & Bull Press, North Hills, PA, USA, 1982. ISBN ???? 137 pp. LCCN Z250 B593 1982. US$60.00.

Blundell:1964:PRR

[Blu64] W. E. Blundell. Printing revolution: Research push brings speedier typesetting, other major advances. The Wall Street Journal, ??(??):??, December 10, 1964.

Blumenfeld:1986:IIT

[Blu86] Matthias Blumenfeld. Interleaf:

die Integration von Text und Grafik. Informationstechnik it: Computer, Systeme, Anwendungen, 6:377–385, 1986. CODEN ICSAES.

Butterfield:1987:DMT

[BM87] Steve Butterfield and Shinichi Mori. 24-wire dot matrix technology changes key printer assumptions. Computer Technology Review, 7(9):101, 104, 108–109, July 1987. CODEN CTERES. ISSN 0278-9647.

Barnes:1887:CFR

[BMD87] William C. Barnes, Joseph W.

McCann, and Alexander Duguid. A collation of facts relative to fast typesetting: together with portraits and biographies of the more famous swift compositors. Concord Co-operative Printing Co., New York, NY, USA, 1887. 80 pp.

Barth:1985:SSS

[BN85] Wilhelm Barth and Heinrich Nirschl. Sichere sinnentsprechende silbentrennung fuer die deutsche sprache. Angewandte Informatik, Applied Informatics, 27(4):152–159, April 1985. CODEN AWIFA7.

Baird:1994:SC

[BN94] Henry S. Baird and George Nagy. Self-correcting 100-font classifier. Proceedings of SPIE — The International Society for Optical Engineering, 2181:106– 115, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Boan:1984:CPC

[Boa84] B. J. Boan. Corrosion of printed circuit boards by photographic chemical vapors. In Corrosion/ 84, International Corrosion Forum Devoted Exclusively to the Protection and Performance of Materials. NACE, Houston, Tex, USA, 1984.

Boehm:1976:SHC

[Boe76]  Peter J. Boehm.  Software

and hardware considerations for a technical typesetting system. IEEE Transactions on Professional Communication, PC-19 (1):15–19, March 1976. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Borko:1985:EBP

[Bor85] Harold Borko. Electronic book publishing: a state of the art report. Microcomputers for Information Management, 2(1):1–10, March 1985. CODEN MIIMEW. ISSN 0742-2342.

Borzo:1996:TLL

[Bor96] Jeanette Borzo. Typography:

Laying down the law. Publish!: the how-to magazine of desktop publishing, 11(6):50–??, June 1, 1996. CODEN PBLIEJ. ISSN 0897-6007.

Bower:1947:MT

[Bow47] W. M. Bower. Mechanical typesetting. Printing theory and practice; 3. Sir I. Pitman, ????, 1947. 161–224 pp.

Barron:1987:TPT

[BR87] David Barron and Mike Rees.

Text processing and typesetting with UNIX. International computer science series. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-14219-8. xiv + 447 pp. LCCN Z286.D47 B37 1987. US$24.95.

Brailsford:1984:HPE

[Bra84] D. F. Brailsford. In-house production of Examination Papers using troff, eqn and tbl. In Proceedings PROTEXT I Workshop, Dublin, pages 21–28. Boole Press, ????, October 1984. URL

http://eprints.nottingham.
ac.uk/363/1/protext.pdf.

Brailsford:2009:ART

[Bra09] David F. Brailsford. Automated re-typesetting, indexing and content enhancement for scanned marriage registers. In Proceedings of the ACM Symposium on Document Engineering (DocEng09), 15–18 September 2009, pages 29–38. ACM Press, New York, NY, USA, 2009.

Brailsford:2013:EBP

[Bra13a] David Brailsford. Extra bits — printing and typesetting history — computerphile. Web video lecture (14m35s)., December 15, 2013. URL https:

//www.youtube.com/watch?v=
HdModNEK_1U.

Brailsford:2013:GJC

[Bra13b] David Brailsford. The great 202 jailbreak — computerphile. Web video lecture (19m54s)., December 12, 2013. URL http://www.

cs.princeton.edu/~bwk/202;
 http://www.eprg.org/papers/
202paper.pdf;   https://www.
youtube.com/watch?v=CVxeuwlvf8w.

Bringhurst:2002:ETS

[Bri02] Robert Bringhurst. The Elements of Typographic Style. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, 2002. ISBN 0-88179-132-6. 350 pp. LCCN Z246.B74 1996. US$29.95.

Bringhurst:2012:ETS

[Bri12] Robert Bringhurst. The Elements of Typographic Style 4.0. Hartley and Marks, Point Roberts, WA, USA, fourth edition, 2012. ISBN 0-88179-211-X (hardcover), 0-88179-212-8 (paperback). 398 pp. LCCN Z246 .B74 2012.

Brown:1969:EMT

[Bro69] C. L. Brown. Experience with Magnetic Tape Selectric Composer (MT/SC) in preparation of monthly abstract bulletin. In Proceedings of the American Society for Information Science, 32nd Annual Meeting, San Francisco, CA, USA, Oct 1–4 1969, volume 6, pages 69–73. American Society for Information Science, Medford, NJ, 1969.

Brown:1985:FID

[Bro85] Heather Brown. Font information and device-independent output. NATO ASI Series, Series F: Computer and Systems Sciences, 17:397–407, 1985. CO-DEN NASFEG. ISBN 3-540-13920-6. ISSN 0258-1248.

Bryan:1984:DDD

[Bry84] Shawn Bryan. Datasouth’s doit-all dynamo. Microcomputing, 8(9):56–58, September 1984. CODEN MIRCDC.

Bryan:1996:DTS

[Bry96] Marvin Bryan. The digital typography sourcebook. John Wiley and Sons, Inc., New York, NY, USA, 1996. ISBN 0-471-14811-3. 416 pp. LCCN Z250.7.B79 1996. US$39.95. Includes CD-ROM.

Bailey:1988:PTH

[BS88] Paul Bailey and Peter Southeran.

Password: typesetting from home micro-computer to photosetter. Sotheran, London, UK, 1988. ISBN 0-905032-19-5 (spiral). 56 pp. LCCN ????

Bo:1996:SAA

[BS96] Lu Bo and Cai Shijie. Skeletonbased approach of automatically generating some Chinese typefaces. Journal of computer science and technology, 11(1):30– 38, January 1996. CODEN JCTEEM. ISSN 1000-9000.

Braun:1992:LRW

[BSS92] Curt C. Braun, N. Clayton Silver, and Barry R. Stock. Likelihood of reading warnings: the effect of fonts and font sizes. Proceedings of the Human Factors Society, 2:926–930, 1992. CO-DEN PHFSDQ. ISSN 0163-5182.

Buckland:1964:CMR

[Buc64] Lawrence F. Buckland. The creation of a machine record of Library of Congress catalog data for distribution and automatic typesetting of cards and book form catalogs. Inforonics, Maynard, MA, USA, 1964. 15 pp.

Buccino:1980:RTS

[Buc80] Joseph H. Buccino. A reliable typesetting system for Waterloo. Report CS-80-20, Dept. of Computer Science, University of Waterloo, Waterloo, ON, Canada, 1980. 111 pp.

Buhl:1980:HSO

[Buh80] Tom Buhl. How to start, operate and enjoy a successful typesetting business. Homefront Graphics, Santa Barbara, Calif. (P.O.

Box 4114, Santa Barbara, CA 93103), 1980. 131 pp.

Buhl:1981:HSO

[Buh81] Tom Buhl. How to start, operate and enjoy a successful typesetting business. Homefront Graphics, Santa Barbara, CA, USA, second edition, 1981. ISBN 0-939374-00-5. 133 pp. LCCN Z244.6.U5 B83 1981.

BTS:1961:BTS

[Bur61] Burns Typesetting Service, Inglewood, CA, USA. Burns Typesetting Service [type specimen book], 1961. various pp.

Butler:1994:RDF

[But94] Timothy Butler. Retaining document format in OCR. Proceedings of SPIE — The International Society for Optical Engineering, 2181:78–86, 1994. CO-DEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Buxton:1988:MLL

[Bux88] H. W. (Harry Wilmot) Buxton.

Memoir of the life and labours of the late Charles Babbage Esq. F.R.S. Reprint series for the history of computing, vol. 13. MIT Press, Cambridge, MA, USA, 1988. ISBN 0-262-02269-9. xix + 401 pp. LCCN QA29.B2B89 1988. US$50. From Computing Reviews: “Charles Babbage was a nineteenth-century polymath who initially studied mathematics at Cambridge and subsequently pursued interests in fields as diverse as philosophy, political economy, and mechanical computing. While attempting to realize a machine for calculating and printing error-free mathematical tables, he visited factories to learn metalworking techniques, invented a system of mechanical notation to represent the states of the machine, and devised a method for correct typesetting. As the editor of this series notes, the scope of Babbage’s work makes him a difficult subject to understand in our modern age of specialized scholarship.”.

Buynak:1985:LLM

[Buy85] William Buynak. Large lowcost memory buffers create improved picture quality. Computer Technology Review, 5(2): 169, 172, 174–175, Spring 1985. CODEN CTERES. ISSN 0278-9647.

Bonham:1982:SPD

[BW82] Mike Bonham and Ian H. Witten. Structured procedural document-preparation language. Proceedings — Canadian Information Processing Society, pages 6–12, 1982. CODEN PCISE7.

Bonham:1984:TDD

[BW84] Mike Bonham and Ian H. Witten. Towards distributed document preparation with interactive and noninteractive viewing. Proceedings — Canadian Information Processing Society, pages 365–372, 1984. CODEN PCISE7.

Byram-Wigfield:1995:PPG

[BW95] David Byram-Wigfield. Practical PostScript: a guide to digital typesetting. Cappella Archive, Great Malvern, UK, 1995. ISBN 0-9525308-0-5 (paperback). 64 pp. LCCN ????

CCGPS:1993:QLT

[Can93] Canada Communication Group, Printing Services, Ottawa, ON, Canada. Quality levels for typesetting = Niveaux de la qualité de la composition, re-edited, December 1993 edition, 1993. iv + 19 + 19 + iv pp.

Cartwright:1980:SST

[Car80] B. Cartwright. A style specification for typesetting manuscripts of research papers using COM-TEXT on the Cyber-76 computer. Division of Soils divisional report; no. 52. CSIRO, Melbourne, Victoria, Australia, 1980. ISBN 0-643-02742-4. 16 pp. LCCN ????

Carter:2002:VET

[Car02] Harry Carter. A View of Early Typography: Up To About 1600, volume 1968 of The Lyell lectures. Hyphen Press, London, UK, 2002. ISBN 0-907259-21-9. 137 pp. LCCN ???? US$35.00. Reprinted with an introduction by James Mosley. Originally published in 1969 by Clarendon Press, Oxford, UK.

Casti:1991:VTP

[Cas91] John L. Casti. VTEX typesetting package. Notices of the American Mathematical Society, 38(2):105–109, February 1991. ISSN 0002-9920 (print), 1088-9477 (electronic).

Catt:1983:CPD

[Cat83] B. R. L. Catt. Computer printer — out of the DP room and into the office. IERE Conference Proceedings, 57:133–139, 1983. CODEN IERCB6. ISBN 0-903748-52-5. ISSN 0538-0006.

Cave:1971:PPa

[Cav71a] Roderick Cave. The private press. Faber and Faber, London, UK, 1971. ISBN 0-571-09443-0. 376 pp. LCCN Z231.5.P7 C37.

Cave:1971:PPb

[Cav71b] Roderick Cave. The private press. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1971. ISBN 0-8352-1695-0. 376 pp. LCCN Z231.5.P7 C3.

Cave:1983:PP

[Cav83] Roderick Cave. The private press. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, second edition, 1983. ISBN 0-8352-1695-0. xvi + 389 pp. LCCN Z231.5.P7 C37 1983. US$64.95.

Chaundy:1954:PMA

[CBB54] T. W. (Theodore William) Chaundy, P. R. Barrett, and Charles Batey, editors. The Printing of Mathematics: Aids for Authors and Editors and Rules for Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1954. vi + 105 pp. LCCN Z250.6.M3 C5 1954.

URL https://www.gwern.net/
docs/design/1954-chaundy-
theprintingofmathematics.
pdf.

Chaundy:1965:PMA

[CBB65] Theodore W. Chaundy, P. R.

Barrett, and Charles Batey, editors. The Printing of Mathematics: Aids for Authors and Editors and Rules for Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1965. vi + 109 pp. LCCN ????

Chaundy:1987:PMA

[CBB87] T. W. (Theodore William)

Chaundy, P. R. Barrett, and Charles Batey, editors. The Printing of Mathematics: Aids for Authors and Editors and Rules for Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second revised edition, 1987. vi + 109 pp. LCCN Z250.6.M3 C5 1965.

Craig:1999:DTB

[CBM99] James Craig, William Bevington, and Susan E. Meyer, editors. Designing with Type: a Basic Course in Typography. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, fourth edition, 1999. ISBN 0-8230-1347-2. 176 pp. LCCN Z246.C69 1999. US$24.95.

Clarke:1988:UMA

[CBT88] A. T. Clarke, B. M. Brown, and M. P. Thorne. Using a micro to automate data acquisition in music publishing. Microprocessing and Microprogramming, 24(1-5):549–553, August 1988. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

ICCT:1965:CCT

[CCT65] The Conference on Computerized Typesetting, Washington, DC., March 2–3, 1965. Composition Information Services, Los Angeles, CA, USA, 1965.

Ciancarini:2012:HQP

[CDFV12] Paolo Ciancarini, Angelo Di Iorio, Luca Furini, and Fabio Vitali. High-quality pagination for publishing. Software—Practice and Experience, 42(6):733–751, June 2012. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carter:2002:TDF

[CDM02] Rob Carter, Ben Day, and Philip Meggs. Typographic Design: Form and Communication. John Wiley and Sons, Inc., New York, NY, USA, third edition, 2002. ISBN 0-471-38341-4. 309 pp. LCCN Z246.C217

2002.  US$50.00.  URL

http://www.loc.gov/catdir/
description/wiley0310/2002027392.
html.

Cordella:1995:NNC

[CDV95] L. P. Cordella, C. De Stefano, and M. Vento. Neural network classifier for OCR using structural descriptions. Machine Vision and Applications, 8(5):336– 342, 1995. CODEN MVAPEO. ISSN 0932-8092.

Chung:1985:TFM

[CH85] C. Y. Chung and H. R. Hwa.

Typesetting forms and mathematical equations in Chinese information processing. In IEEE 1985 Proceedings of the International Conference on Cybernetics and Society, Tucson, AZ, USA, pages 46–50. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. IEEE Service Cent. Piscataway, NJ, USA.

Chave:1997:FYW

[Cha97] Leonard Chave. Forty Years On: The Wynkyn de Worde Society 1957–1997. The Wynkyn de Worde Society, Hatfield, UK, 1997. ISBN 1-871224-14-4. ???? pp. LCCN ???? US$60.

Chen:2000:IDF

[CHM00] Pehong Chen, Michael A. Harrison, and Ikuo Minakata. Incremental document formatting. In ACM Conference on Document Processing Systems: proceedings: December 5–9, 1988, Santa Fe, New Mexico, pages 93–100. ACM Press, New York, NY, USA, 2000. ISBN 0-89791-291-8. LCCN QA76.9.T48 A25 1988. ACM order number 429882.

Christian:1987:UTP

[Chr87] Kaare Christian. The UNIX text processing system. John Wiley and Sons, Inc., New York, NY, USA, 1987. ISBN 0-471-85581-2. xii + 50 pp. LCCN QA76.76.O63 C47 1987. US$24.95.

CIS:1967:RCT

[CIS67] The realities of computerized typesetting; a collection of the 15 vital papers presented at the Conference … sponsored by the Research and Engineering Council. Composition Information Services, Los Angeles, CA, USA, 1967.

Cisler:1987:DPL

[Cis87] Steve Cisler. Desktop publishing in libraries. Online (Weston, Connecticut), 11(5):64–72, September 1987. CODEN ON-LIDN. ISSN 0146-5422.

Cannat:1986:LTAb

[CK86] J.-J. Cannat and Y. Kodratoff. Learning techniques applied to multi-font character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 635:469–479, 1986. CODEN PSISDG. ISBN 0-89252-670-X. ISSN 0277-786X (print), 1996-756X (electronic).

Chu:1988:STG

[CK88] Jiang-Hsing Chu and Gary D.

Knott. SPEC: a TEX-embedded graphics typesetting language. Computer science technical report series; CS-TR-2118 CAR; TR-395 Computer science technical report series (College Park, Md.); TR- 2118. CAR (Series); TR-395. University of Maryland, College Park, MD, USA, October 1988. 32 pp.

Cannat:1986:LTAa

[CKM86] J.-J. Cannat, Y. Kodratoff, and S. Moscatelli. Learning techniques applied to multi-font character recognition. Proceedings — International Conference on Pattern Recognition, pages 123–125, 1986. CODEN PI-CREG. ISBN 0-8186-0742-4. IEEE Service Cent. Piscataway, NJ, USA.

Condon:1980:EML

[CKT80] Joe Condon, Brian Kernighan, and Ken Thompson. Experience with the Mergenthaler Linotron 202 phototypesetter, or, how we spent our summer vacation. Technical Memorandum TM-80-1270-1, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 6, 1980. 11

+ 2 pp. URL http://www.
cs.princeton.edu/~bwk/202/

. Publication of this internal memo of 1980 was suppressed for 33 years, but both the original, and the reconstructed memo are now available on the Web. See [BBK13b] for the story of its reconstruction and Web publication.

Clancy:1964:ETA

[Cla64] Gerald Francis Clancy. Editing techniques in automatic typesetting systems. Thesis (B.S.), Massachusetts Institute of Technology, Department of Electrical Engineering, Cambridge, MA, USA, 1964. 31 pp.

Clement:1982:CCW

[Cle82]  Frank Clement.  The com-

puter: Changing the way we do things. electronic typesetting: The means to improving reading speed and comprehension? Performance & Instruction Journal, 21(4):39–41, May 1982. ISSN 8750-0191.

Clover:1980:WDY

[Clo80] Margaret Clover. Why not do your own “Setting” with electronic typesetting in the classroom? C.S.P.A.A. Bulletin, 38 (1):7–9, Summer 1980.

CTC:1930:TBD

[Col30] Columbus Typesetting Company, Columbus, O. Type … being a demonstration of the variety of Linotype and Monotype faces that are available at Columbus Typesetting Company, 1930. 8 + 158 pp.

Collison:1962:IBM

[Col62] Robert Lewis Collison. Indexing books: a manual of basic principles. De Graff, New York, NY, USA, 1962. 96 pp. LCCN Z695.9 .C64.

Collins:1986:DFL

[Col86] J. S. Collins. Digital fonts for low-cost typesetting systems. Proceedings — IEEE Computer Society International Conference, pages 82–87, 1986. CODEN PCICDQ. ISBN 0-8186-0692-4. IEEE Service Cent. Piscataway, NJ, USA.

Collins:1989:ADT

[Col89] John Collins. ABCs of digital type. BYTE Magazine, 14(12): 403–408, November 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

CompugraphicCorporation:1900:ITT

[Com00] Compugraphic Corporation. An introduction to typography and typesetting: Part I: Typography concepts. Compugraphic Corporation, Wilmington, MA, USA, 1900. 66 pp.

CIS:1964:CGA

[Com64] Composition Information Services, Los Angeles, CA, USA. CIS glossary of automated typesetting and related computer terms, 1964. 47 pp.

CIS:1966:GAT

[Com66] Composition Information Services, Los Angeles, CA, USA. Glossary of automated typesetting and related computer terms, second edition, 1966. 105 pp. First ed. published in 1964 under title: CIS glossary of automated typesetting and related computer terms.

CIS:1969:GAT

[Com69] Composition Information Services, Los Angeles, CA, USA. Glossary of automated typesetting and related computer terms: Supplement, second edition, 1969. 56 pp.

UC:1979:TEF

[Com79] Computer Center, University of California, Santa Barbara. Text editing, formatting and typesetting using the UCSB UNIX system, 1979. 244 pp.

Compugraphic:1983:MTC

[Com83]  Compugraphic  Corporation,

Wilmington, MA, USA. Modular Composition System (MCS) 10/ 100 typesetting commands: reference manual, 1983. 71 pp.

UACS:1984:AQC

[Com84] Computing Services, University of Alberta, Edmonton, AB, Canada. APSQ queuing computer typesetting, 1984. 3 pp.

Conrad:1988:TMN

[Con88] Phillip T. Conrad. Typesetting of music notation using TEX. Thesis (M.S.), West Virginia University, Morgantown, WV, USA, 1988. vii + 207 pp.

Cost:1986:CLB

[Cos86] Patricia Cost. The contributions of Linee Boyd Benton and Morris Fuller Benton to the technology of typesetting and typeface design. Thesis (M.S.), Rochester Institute of Technology, Rochester, NY, USA, 1986. ix + 257 pp.

Craig:1983:DTB

[Cra83a] James Craig. Designing with type: a basic course in typography. IEEE Transactions on Professional Communication, PC-26(4):213, December 1983. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Craig:1983:PDM

[Cra83b] James Craig. Phototypesetting:

a design manual. IEEE Transactions on Professional Communication, PC-26(4):213–214, December 1983. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Crawford:1987:CSW

[Cra87a] Walt Crawford. Common sense wordworking III: desktop publishing and desktop typesetting. Library Hi Tech, 5(1):43–52, Spring 1987. CODEN LIHTD2. ISSN 0737-8831.

Crawford:1987:DTL

[Cra87b] Walt Crawford. Desktop typesetting and the LITA newsletter. Information Technology and Libraries, 6(2):89–96, June 1987. CODEN ITLBDC. ISSN 0730-9295.

Cross:1980:IB

[Cro80] Ruth Canedy Cross. Indexing books. A Word Guild guide. Word Guild, Cambridge, MA, USA, 1980. ISBN 0-9602524-1-X. xiii + 145 pp. LCCN Z695.9 .C76.

Crystal:2015:MPP

[Cry15] David Crystal. Making a point: the persnickety story of English punctuation. St. Martin’s Press, New York, NY, USA, 2015. ISBN 1-250-06041-9 (hardcover), 1-4668-6564-4 (e-book). 400 (est.) pp. LCCN PE1450 .C89 2015.

Cooke:1986:MMT

[CS86] J. Robert Cooke and E. Ted Sobel. MathWriter: mathematical typesetting with the Macintosh. Cooke Publications, Ithaca, 1986. vii + 106 pp.

Checkoway:2010:TOD

[CSR10a] Stephen Checkoway, Hovav Shacham, and Eric Rescorla. Are text-only data formats safe? Or, use this LaTEX class file to pwn your computer. In ????, editor, 3rd USENIX Workshop on Large-Scale Exploits and Emergent Threats, LEET ’10, page ?? USENIX, Berkeley, CA, USA, 2010.

Checkoway:2010:DTL

[CSR10b] Stephen Checkoway, Hovav Shacham, and Eric Rescorla. Don’t take LaTEX files from strangers. ;login: the USENIX Association newsletter, 35(1): 17–22, August 2010. URL

https://www.usenix.org/system/
files/login/articles/73506-
checkoway.pdf.

CTC:1965:CTC

[CTC65] Computer Typesetting Conference, London University, 1964. Report of proceedings, Institute of Printing, Ltd. Publications. Institute of Printing, London, UK, 1965. LCCN Z253 .C69 1964c.

ICTCUS:1966:CTC

[CTC66] Advances in computer typesetting: proceedings of the 1966 International Computer Typesetting Conference. Sussex University, July 15–16, 1966. Institute of Printing, London, UK, 1966. LCCN Z 253.3 I61 1966.

Cusick:2011:WOL

[Cus11] Rick Cusick. What our lettering needs: the contribution of Hermann Zapf to calligraphy and type design at Hallmark Cards. RIT Cary Graphic Arts Press, Rochester, NY, USA, 2011. ISBN 1-933360-55-0 (paperback). xv + 117 pp. LCCN Z250.A2 Z374 2011.

Corbett:1982:IPD

[CW82] Chris Corbett and Ian H. Witten. On the inclusion and placement of documentation graphics in computer typesetting. The Computer Journal, 25(2):272– 277, May 1982. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Casey:1985:TRU

[CW85] R. G. Casey and K. Y. Wong.

Text recognition using adaptive software. In GLOBECOM ’85: IEEE Global Telecommunications Conference — Conference Record, New Orleans, LA, USA, pages 353–357. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. IEEE Service Cent. Piscataway, NJ, USA.

Chen:1993:WSS

[CWB93] Francine R. Chen, Lynn D. Wilcox, and Dan S. Bloomberg. Word spotting in scanned images using hidden Markov models. Proceedings — ICASSP, IEEE International Conference on Acoustics, Speech and Signal Processing, 5:V–1–V–4, 1993. CODEN IPRODJ. ISBN 0-7803-0946-4. ISSN 0736-7791. IEEE catalog number 93CH3252-4.

Dallwitz:1980:UGT

[Dal80] M. J. Dallwitz. User’s guide to TYPSET, a computer typesetting program. CSIRO Division of Entomology report; no. 18. Commonwealth Scientific and Industrial Research Organization Australia, Canberra, ACT, Australia, 1980. ISBN 0-643-02499-9. 23 pp. LCCN ????

Davis:1969:CTU

[Dav69] H. J. Davis. Computer typesetting in the U.K.: a review. OSTI/DPC report; no. 4001. Office for Scientific and Technical Information, Documentation Processing Centre, Manchester, UK, 1969. ISBN 0-85522-000-7. 22 pp. LCCN ????

Davis:1988:TEP

[Dav88] Ronnie H. Davis. Typesetting:

an economic perspective. National Composition Association, Arlington, VA, USA, 1988. 11 pp.

Day:1983:TMM

[Day83]  Robert A. Day.  Typeset-

ting mathematics on multiaccess systems. Software— Practice and Experience, 13(2): 131–138, February 1983. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deBiasi:2002:SDP

[dBD02] Pierre-Marc de Biasi and Karine Douplitzky. La saga du papier. Collection “Textures”. Arte Editions, Issy-les-Moulineaux (Hauts-de-Seine), 2002. ISBN 2-87660-350-0 (paperback), 2-87660-228-8 (hardcover). 255 pp. LCCN TS1094 .B53 1999.

URL http://catalogue.bnf.
fr/ark:/12148/cb38857166s.

Duncan:1963:CTE

[DEM^{+}63] C. J. Duncan, J. Eve, L. Molyneux, E. S. Page, and Margaret G. Robson. Computer typesetting: an evaluation of the problems. Printing Technology, 7(??):133– 151, ???? 1963.

ADL:1972:TMD

[Dep72] Dept. of Labour and National Service, Melbourne, Victoria, Australia. Typesetting in metropolitan daily newspapers, 1972. 42 pp.

USBC:1979:CMI

[Dep79] Dept. of Commerce, Bureau of the Census, Washington, DC, USA. 1977 census of manufactures: industry series, typesetting, SIC 2791, 1979. 7 pp.

Desarmenien:1986:TSD

[Des86] J. Desarmenien, editor. TEX for scientific documentation: second European conference, Strasbourg, France, June 19–21, 1986: proceedings, volume 236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. CO-DEN LNCSD9. ISBN 0-387-16807-9 (USA: paperback), 3-540-16807-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.4.T47 T491 1986.

Desyatnik:1989:CPP

[Des89] E. S. Desyatnik. Certain problems with the performance of automated word processing systems. Soviet Energy Technology (English translation of Energomashinostroenie), 9:73–74, 1989. CODEN SETEDW. ISSN 0734-1024.

Normung:1994:FFM

[Deu94] Deutsches Institut für Normung. Formelzeichen, Formelsatz, mathematische Zeichen und Begriffe: Normen (German) [Symbols, composition of mathematical notation, mathematical marks and terms: Standards]. Number 202 in DIN Taschenbuch. Beuth, Berlin, Germany, second edition, April 1994. ISBN 3-410-12954-5. xii + 478 pp. LCCN ????

DIN:1998:DAD

[Deu98] Deutsches Industrie Norm. DIN 16507-2, Ausgabe: 1998-04, Drucktechnik — Schriftgrößen — Teil 2: Digitaler Satz und verwandte Techniken (German) [Printing technology — Type sizes — Part 2: Digital Typesetting and related techniques]. Deutsches Institut für Normung, e.V., Burgrafenstaße 6, 10787 Berlin, Germany, April 1998. ISBN ???? 9 pp. LCCN

????  URL http://www.cl.
cam.ac.uk/~mgk25/volatile/
din-16507-2.pdf.  German

draft standard on metric typography.

Dewey:1986:TSP

[Dew86] Michael E. Dewey. TBFLP:

a simple preprocessor for tables. Software—Practice and Experience, 16(8):731–738, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deMoor:1999:BAG

[dG99] O. de Moor and J. Gibbons.

Bridging the algorithm gap: a linear-time functional program for paragraph formatting. Science of Computer Programming, 35(1):3–??, ???? 1999. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dickey:1987:TAU

[Dic87]  Leroy J. Dickey.  Typeset-

ting APL using TROFF. In John Haliburton, editor, APL in transition, pages 494–497. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-226-

8. LCCN ???? US$37.00.

DiVittorio:1984:HDI

[DiV84]  Mark J. DiVittorio.  High-

quality, dot-matrix impact printer family. Hewlett–Packard Journal: technical information from the laboratories of Hewlett– Packard Company, 35(11):30– 32, November 1984. CODEN HPJOAX. ISSN 0018-1153.

Dougherty:1987:UTP

[DO87] Dale Dougherty and Tim O’Reilly, editors. UNIX Text Processing. Hayden Books UNIX library system. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1987. ISBN 0-672-46291-5 (paperback). xiv + 665 pp. LCCN Z286.D47 D68 1987. US$26.95. URL ftp:

//ftp.ffii.org/pub/groff/
contrib/documentation/utp/
;   http://books.google.com/
books?id=H148AQAAIAAJ;   http:
//catalog.hathitrust.org/
api/volumes/oclc/15722288.
html;   http://proquest.safaribooksonline.
com/?fpi=9780810462915;
http://www.oreilly.com/openbook/
utp/UnixTextProcessing.pdf
 http://proquest.tech.safaribooksonline.
de/9780810462915.

Dodd:1997:ASG

[Dod97] Janet S. Dodd, editor. The ACS Style Guide: a Manual for Authors and Editors. American Chemical Society, Washington, DC, USA, second edition, 1997. ISBN 0-8412-3461-2 (hardcover), 0-8412-3462-0 (paperback). xi + 460 pp. LCCN QD8.5.A25 1997. US$27.50.

Doob:1989:TTA

[Doo89] Michael Doob. TEX and typesetting—an author’s view. Notices of the American Mathematical Society, 36(9):1203–1204, November 1989. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dowding:1966:FPS

[Dow66] Geoffrey Dowding. Finer Points in the Spacing and Arrangement of Type. Wace, London, UK, third edition, 1966. xiv + 79 pp. LCCN Z253 .D69 1966.

Dowdle:1983:IHY

[Dow83] Wade B. Dowdle. 101 ideas to help you sell more typesetting. National Composition Association, Arlington, VA, USA, 1983. 144 pp.

Dowding:1993:FPS

[Dow93] Geoffrey Dowding. Finer Points in the Spacing and Arrangement of Type. Hartley and Marks, Point Roberts, WA, USA, 1993. ISBN 0-88179-086-9. xvii + 90 pp. LCCN Z253 .D69 1993.

Dowding:1995:FPS

[Dow95] Geoffrey Dowding. Finer Points in the Spacing and Arrangement of Type. Hartley and Marks, Point Roberts, WA, USA, 1995. ISBN 0-88179-119-9 (paperback). xvii + 90 pp. LCCN Z253 .D69 1993.

Dowding:1997:IHP

[Dow97] Geoffrey Dowding. An Introduction to the History of Printing Types. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, 1997. ISBN 1-884718-43-4 (hardcover), 1-884718-44-2 (paperback). 304 pp. LCCN ???? US$29.95 (paperback), US$39.95 (hardcover).

Dowding:1998:IHP

[Dow98] Geoffrey Dowding. An Introduction to the History of Printing Types: an Illustrated Summary of the Main Stages in the Development of Type Design from 1440 Up to the Present Day: an Aid to Type Face Identification. British Library, London, UK, 1998. ISBN 0-7123-4563-9 (U.K. paperback), 0-7123-4577-9 (U.K. cased), 1-884718-43-4 (U.S. cased), 1-884718-44-2 (U.S. paperback). xxiv + 277 pp. LCCN Z250.A2 D6 1998.

Doyon:1993:TT

[Doy93] Roy Doyon. Topography and typesetting. Cartographica, 30 (??):108, Summer 1993. ISSN 0317-7173.

Dallas:1983:RTL

[DS83] I. N. Dallas and E. B. Spratt, editors. Ring technology: local area networks: Proceedings of the IFIP WG 6.4 Univ. of Kent workshop (Kent, UK, Sept. 28–30, 1983). Elsevier North-Holland, Inc., New York, NY, USA, 1983. ISBN 0-444-86852-6. LCCN TK5105.5 .I3427 1983.

Dunn:1980:PLG

[Dun80] S. Thomas Dunn, editor. Proceedings — Lasers in Graphics, Electronic Publishing in the 80’S Conference, Volume 1: Pulling It Together — Text and Graphics for Newspaper Production; Volume 2: Pulling It Together — Text and Graphics for Commercial Printing, 1980. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980.

Duncan:1983:DPE

[Dun83] Harry Duncan. Doors of perception: essays in book typography. W. Thomas Taylor, Austin, TX, USA, 1983. ISBN ???? 99 pp. LCCN Z116 .D6 1983. Reprinted in 1987.

Durazzi:2022:RTM

[Dur22] Allison Durazzi. Review of “Type Matters: The Rhetoricity of Letterforms edited by Christopher Scott Wyatt and Dànielle Nicole DeVoss,” (2018). Parlor Press. Communication Design Quarterly Review, 10(1): 54–56, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3507454.3507460.

Eckersley:1994:GTT

[EAE^{+}94] Richard Eckersley, Richard Angstadt, Charles M. Ellertson, Richard Hendel, Naomi B. Pascal, and Anita Walker Scott. Glossary of Typesetting Terms. Chicago Guides to Writing, Editing, and Publishing. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-18371-8. xi + 169 pp. LCCN Z253.G57 1994. US$20.00.

Eckersley:1996:GTT

[EAE^{+}96] Richard Eckersley, Richard Angstadt, Charles M. Ellertson, Richard Hendel, Naomi B. Pascal, Anita Walker Scott, and Linda L. Oestreich. Glossary of typesetting terms. Technical Communication, 43(2):179– ??, ???? 1996. CODEN TL-CMBT. ISSN 0049-3155.

Edvardsson:1994:SGT

[Edv94] I. R. Edvardsson. Skill, gender and technical change in a Nordic environment: typesetting in Iceland and Sweden. New Technology, Work, and Employment, 9 (1):30–42, 1994. CODEN NWE-MEJ. ISSN 0268-1072.

ElGowely:1990:MRM

[EEN90] Khaled El Gowely, Ossama El Dessouki, and Ahmed Nazif. Multi-phase recognition of multi-font photoscript Arabic text. Proceedings — International Conference on Pattern Recognition, 1:700–702, 1990. CODEN PICREG. ISBN 0-8186-2062-5. IEEE catalog number 90CH2898-5.

Entlich:1997:MDL

[EGL^{+}97] Richard Entlich, Lorrin Garson, Michael Lesk, Lorraine Normore, Jan Olsen, and Stuart Weibel. Making a digital library: The contents of the CORE project. ACM Transactions on Information Systems, 15(2):103–123, April 1997. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80/tois/abstracts/
entlich.html.

Eisner:1988:PPR

[Eis88] William S. Eisner. Printing processes — the revolution continues. Annual Meeting — Technical Section, Canadian Pulp and Paper Association, Preprints, A: a275–??, 1988. CODEN PM-CADQ. ISBN 0-919893-49-X. ISSN 0316-6732.

Eisenberg:1990:ITT

[Eis90] Daniel Eisenberg. In-house typesetting on a tight budget. Scholarly Publishing, 21(4):205–??, July 1990. ISSN 0036-634X.

Esakov:1994:CDO

[ELS94] Jeffrey Esakov, Daniel P. Lopresti, and Jonathan S. Sandberg. Classification and distribution of optical character recognition errors. Proceedings of SPIE — The International Society for Optical Engineering, 2181:204–216, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Elstner:1984:DFR

[EM84] J. Elstner and D. Maignan. Diagnostics for font read-only store. IBM Technical Disclosure Bulletin, 27(5):3049–??, October 1984. CODEN IBMTAA. ISSN 0018-8689.

Emerson:1987:TTU

[EP87] Sandra L. Emerson and Karen Paulsell. Troff typesetting for UNIX systems. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, 1987. ISBN 0-13-930959-4. xx + 359 pp. LCCN Z253.4.U53E45 1987. US$29.95.

Eason:1991:RHT

[ERBR91] Ron Eason, Sarah Rookledge, Phil Baines, and Gordon Rookledge, editors. Rookledge’s handbook of typedesigners: a biographical directory. Sarema, Carshalton Beeches, Surrey, UK, 1991. ISBN 1-870758-09-9. vi + 208 pp. LCCN ????

Ezell:1987:TSA

[Eze87] Christopher Clarence Ezell. Typesetting: sociological analysis of changing newspaper technology. Thesis (Ph.D. in sociology), Vanderbilt University, Nashville, TN, USA, 1987. x + 222 pp.

Fournier:1985:GCI

[FB85] Alain Fournier and Brian A. Barsky. Geometric continuity with interpolating Bézier curves. Proceedings — Graphics Interface, pages 337–341, 1985. CO-DEN PGINEK. ISSN 0713-5424.

Fiscella:1990:PCM

[FBJ90] M. D. Fiscella, S. A. Brownstein, and D. A. Johnson. Pulse count modulation. A novel head drive method for thermal printing. Proceedings of SPIE — The International Society for Optical Engineering, 1252:156–167, 1990. CODEN PSISDG. ISBN 0-8194-0299-0. ISSN 0277-786X (print), 1996-756X (electronic).

Foster:1977:EST

[FC77] Jeremy Foster and Peter Coles.

Experimental study of typographic cueing in printed text. Ergonomics, 20(1):57–66, January 1977. CODEN ERGOAX. ISSN 0014-0139.

Felici:2003:CMT

[Fel03] James Felici. The Complete Manual of Typography: a Guide to Setting Perfect Type. Adobe Press, Mountain View, CA, USA, 2003. ISBN 0-321-12730-7. xx + 360 pp. LCCN Z250.F44 2003. US$45.00. Foreword by Frank Romano.

Feruglio:1993:TCT

[FF93] G. V. Feruglio and R. Fuster.

Typesetting Catalan texts with TEX. In Anonymous [Ano93a], pages 252–259. ISSN 0896-3207.

Fishburne:2000:MPG

[Fis00] William Fishburne. Margination and Project Gutenberg.

Dr. Dobb’s Journal of Software Tools, 25(5):52, 54–56, May 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/2000/2000_
05/gutenbg.txt;   http://
www.ddj.com/ftp/2000/2000_
05/gutenbg.zip. See letter

[GF00].

Fuchs:1985:OPF

[FK85] D. R. Fuchs and D. E. Knuth. Optimal prepaging and font caching. ACM Transactions on Programming Languages and Systems, 7(1):62–79, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fletcher:1988:NAS

[Fle88] Harry George Fletcher, III. New Aldine Studies: Documentary Essays on the Life and Work of Aldus Manutius. B. M. Rosenthal, Inc., San Francisco, CA, USA, 1988. ISBN 0-9600094-1-8. xiii + 206 pp. LCCN Z232.M3 F57 1988. US$65.00.

Flores:1973:TUM

[Flo73] Ivan Flores. Tutorial use of multiple type faces. IEEE Transactions on Professional Communication, PC-16(1):2–6, March 1973. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Fenn:1994:RMT

[FM94] Patricia Fenn and Alfred P. Malpa. Rewards of Merit; tokens of a child’s progress and a teacher’s esteem as an enduring aspect of American religious and secular education. Howell Press, Inc., Charlottesville, VA, USA, 1994. ISBN 0-943231-68-X. 224 pp. LCCN ???? US$39.95. Includes nearly 500 color examples of engraving, calligraphy, printing, illustration, and folk art.

Fujieda:1985:FPLa

[FN85a] Hideo Fujieda and Hiroki Nagai. Font pattern loading method for Kanji-printers in the DIPS-106 Network OS. Denki Tsushin Kenkyusho Kenkyu Jitsuyoka Hokoku, 34(2):317–325, 1985. CODEN DTKKAA. ISSN 0415-3200.

Fujieda:1985:FPLb

[FN85b] Hideo Fujieda and Hiroki Nagai.

Font pattern loading method for Kanji printers in the DIPS-106 Network OS. Review of the Electrical Communication Laboratories (Tokyo), 33(4):725–731, July 1985. CODEN RELTAN. ISSN 0418-6338.

Felici:1987:DPS

[FN87] James Felici and Ted Nace.

Desktop publishing skills: a primer for typesetting with computer and laser printer. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-11537-9. vii + 180 pp. LCCN Z286.D47 F44 1987. US$19.95.

Foderaro:1978:TME

[Fod78] John Keith Foderaro. Typesetting MACSYMA equations. Thesis (Master of Science, Plan II), Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1978. ???? pp.

Forsley:1989:RFC

[For89] Lawrence P. Forsley, editor. 1989 Rochester Forth Conference. Industrial Automation: June 20–24, 1989, University of Rochester. Institute for Applied FORTH Research, Rochester, NY, USA, 1989. ISBN 0-914593-09-9. LCCN TJ212.2.A15 1989.

Foster:1881:HBM

[Fos81] Charles Foster. How books are made: a description of typesetting, wood engraving, electrotyping, printing, and other interesting features of book manufacture. ????, Philadelphia, PA, USA, 1881. 24 pp.

Foster:1976:DPT

[Fos76] Ian Foster. Dictionary of printing terms: a comprehensive guide to the terms most commonly used in printing, paper, typesetting, origination, finishing, point of sale, premiums and promotions etc. Print Buyer Magazine, London, UK, revised edition, 1976. ISBN 0-905768-00-0. 112 pp. LCCN Z118.D48 1976.

Fowler:1973:PIU

[Fow73] R. L. Fowler. Printing italics using a standard font. IBM Technical Disclosure Bulletin, 15 (9):??, February 1973. CODEN IBMTAA. ISSN 0018-8689.

Foxley:1987:MLT

[Fox87]  Eric Foxley.  Music — a

language for typesetting music scores. Software—Practice and Experience, 17(8):485–502, August 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freeman:1990:TCI

[Fre90] Vivian Leigh Freeman. Has typography changed with the invention of typesetting machines? Thesis (B.S.), East Texas State University, Commerce, TX, USA, 1990. iv + 27 pp.

Frederiksen:1996:AA

[Fre96] Paul A. Frederiksen. Automation annotations. Colorado Lawyer, 25(5):65–??, May 1, 1996. ISSN 0363-7867.

Friedland:1970:KOC

[Fri70] Andrew Mark Friedland. Keyboard operation for computeraided typesetting. Thesis (M.S. in Engineering), Univ. of California, Berkeley, Berkeley, CA, USA, September 1970. various pp.

Friends:1983:TWD

[Fri83] Friends of Hermann Puterschein.

A Tribute to W. A. Dwiggins on the hundredth anniversary of his birth. Inkwell Press, New York, NY, USA, 1983. ISBN ???? 153 pp. LCCN Z 116 A3 T74 1983.

Furuta:1982:DFS

[FSS82] R. Furuta, J. Scofield, and A. Shaw. Document formatting systems: survey, concepts, and issues. In Nievergelt et al. [NCNS82], pages 133–220. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Fateman:1996:PRT

[FT96] Richard J. Fateman and Taku Tokuyasu. Progress in recognizing typeset mathematics. Proceedings of SPIE — The International Society for Optical Engineering, 2660:37–50, 1996. CODEN PSISDG. ISBN 0-8194-2034-4. ISSN 0277-786X (print), 1996-756X (electronic).

Fateman:1996:OCR

[FTBM96] Richard J. Fateman, Taku Tokuyasu, Benjamin P. Berman, and Nicholas Mitchell. Optical character recognition and parsing of typeset mathematics. Journal of Visual Communication and Image Representation, 7(1):2–15, March 1996. CODEN JVCRE7. ISSN 1047-3203 (print), 1095-9076 (elec-

tronic).  URL http://www.
cs.berkeley.edu/~fateman/
papers/jvcir.ps.Z.

Fujita:1994:TSF

[Fuj94] Shinsaku Fujita. Typesetting structural formulae with the text formatter TEX/LaTEX. Computers in Chemistry, 18(2):109–116, 1994.

Foreman:1993:SEC

[FZ93] Lynn M. Foreman and Stuart H.

Zweben. Study of the effectiveness control and data flow testing strategies. The Journal of Systems and Software, 21 (3):215–228, June 1993. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gag:1929:TE

[Gag29] Wanda Gag. Typesetting equipment, 1929. 1 art print.

Gardner:1967:TIE

[Gar67] Christina Gardner. Typesetting with IBM Executive typewriters. Gardner Printing, Santa Rosa, CA, USA, 1967. 180 pp.

Gardner:1973:ETI

[Gar73] Christina Gardner. Effective typesetting with the IBM tape composer. North American Pub. Co., Philadelphia, PA, USA, 1973. ISBN 0-912920-20-3. 156 pp. LCCN Z253.3.G35.

Garner:1980:GED

[Gar80] Franklin J. Garner. Grafica — merging typesetting and plotting in automated publications systems. Research Publication — General Motors Research Laboratories, ??(??):83–92, ???? 1980.

Gauthier:1984:UTT

[Gau84] Richard L. Gauthier. Using the TEX typesetting language. Reston Publishing Co., Inc., Reston, VA, USA, 1984. ISBN 0-8359-8160-6 (paperback). viii + 248 pp. LCCN Z253.4.T47G38 1984.

GilmourBryson:1982:TPC

[GB82] A. Gilmour-Bryson. Text preparation for computer typesetting. ALLC Bull, 10(2):40–47, 1982. CODEN ALLCB5. ISSN 0951-1474.

Goudsmit:1958:Ec

[GdMMP58] S. A. Goudsmit, E. der Mateosian, M. McKeown, and S. Pasternack. Editorial. Physical Review Letters, 1(6):197– 198, September 15, 1958. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Gehani:1986:DFT

[Geh86] Narain Gehani. Document formatting and typesetting on the UNIX system. Silicon Press, Summit, NJ, USA, 1986. ISBN 0-9615336-0-9. xv + 364 pp. LCCN Z52.5.U54 G43 1986. US$33.95.

Gehani:1987:DFT

[Geh87] Narain Gehani. Document Formatting and Typesetting on the UNIX System. Silicon Press, Summit, NJ, USA, second edition, 1987. ISBN 0-13-938325-5, 0-9615336-2-5. xv + 377 pp. LCCN Z52.5.U54G43 1987. US$40.90.

Gelfand:1991:CBT

[Gel91] Andrew Gelfand. A computer based training program on typesetting for desktop publishing professionals. Master’s thesis, New York Institute of Technology, New York, NY, USA, 1991. 261 pp.

Geschke:1986:PPD

[Ges86] Charles M. Geschke. PostScript — a page description language. Informationstechnik it: Computer, Systeme, Anwendungen, 6:370–376, 1986. CODEN IC-SAES.

Gantz:2000:LM

[GF00] Howard Gantz and William Fishburne. Letters: Margination. Dr. Dobb’s Journal of Software Tools, 25(9):10, 12, September 2000. CODEN DDJOEB. ISSN 1044-789X. See [Fis00].

Griffiths:1997:LL

[GH97] David F. (Francis) Griffiths and Desmond J. Higham. Learning LaTEX. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-383-8. x + 84 pp. LCCN Z253.4.L38G75 1997.

Ghemawat:1987:QEM

[Ghe87] Mahipal S. Ghemawat. Quick and efficient method of correct typesetting of tables on computers. Computers in Biology and Medicine, 17(4):293–295, 1987. CODEN CBMDAW. ISSN 0010-4825.

Ghosh:1984:BDI

[Gho84] P. K. Ghosh. Basic design issues in a multi-lingual type font design and typesetting workstation. J. Inst. Electron.

Telecommun. Eng, 30(6):196– 202, November 1984. CODEN JIETAU. ISSN 0377-2063.

Gibbs:1995:PBN

[Gib95] Ron Gibbs. Printing benefits from new technologies. Laser focus world, 31(11):77–82, November 1995. CODEN LFWOE8. ISSN 0740-2511.

Gilchrist:1989:TTP

[Gil89] Martin Gilchrist. TEX and typesetting—A publisher’s view. Notices of the American Mathematical Society, 36(9):1199–1201, November 1989. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gill:1991:EGB

[Gil91] Evan Gill. Eric Gill: a bibliography. St. Paul’s Bibliographies, Winchester, UK, second edition, 1991. ISBN 0-906795-53-2. LCCN Z8342.2 .G5 1991. Revised by D. Steven Corey and Julia MacKenzie. Original edition published by Cassell, London, UK, 1953.

Gulbins:2000:MTK

[GK00] Jürgen Gulbins and Christine Kahrmann. Mut zur Typographie: ein Kurs für Desktoppublishing. Springer-Verlag Inc., New York, NY, USA, 2000. ISBN 3-540-67541-8. x + 430 pp. LCCN ????

Glinert:1984:LFV

[GL84] Ephraim P. Glinert and Richard E. Ladner. Large font virtual terminal interface. a software prosthesis for the visually impaired.

Communications of the ACM, 27 (6):567–572, June 1984. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gehani:1988:DFT

[GL88] Narain Gehani and Steven Lally.

Document formatting and typesetting on the UNIX system. Vol. 2: grap, mv, ms, and troff. Silicon Press, Summit, NJ, USA, 1988. ISBN 0-9615336-3-3. xiii + 304 pp. LCCN Z52.5.U54 G431 1988. US$30.95.

Goldberg:1983:INP

[GM83] Jeffrey L. Goldberg and Thomas C.

Miller. Interactive name placement for provisional maps. Technical Papers of the American Congress of Surveying and Mapping, pages 314–321, 1983. CO-DEN TPAMDF.

Grobelnik:1994:ISE

[GMZZ94] M. Grobelnik, D. Mladenic, D. Zupanic, and B. Znidar. Integrated system for encyclopaedia typesetting based on TEX. In Goossens and Rahtz [GR94], pages 386–387. ISSN 0896-3207.

Goldman:1969:CAT

[Gol69] M. S. Goldman. Computeraided typesetting with a CRT. The Computer Journal, 12 (4):357–364, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/120357.sgm.abs.html;

http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/357.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/358.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/359.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/360.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/361.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/362.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/363.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_
04/tiff/364.tif.

Goldfarb:1973:DCI

[Gol73]  Charles F. Goldfarb.  De-

sign considerations for integrated text processing systems. Technical Report G320-2094, IBM Cambridge Scientific Center, Cambridge, MA,

USA, ????  1973.  URL

http://www.sgmlsource.com/
history/G320-2094/G320-2094.
htm.

Goodstein:1982:GTC

[Goo82] David Henry Goodstein. Glossary, typesetting/computer/communications terms. National Composition Association, Arlington, VA, USA, 1982. 47 pp.

Gorbach:2001:TP

[Gor01] Rudolf Paulus Gorbach. Typographie professionell. Galileo Press, Bonn, Germany, 2001. ISBN 3-934358-73-X. ca. 350 pp. LCCN ???? DM 99.90, EUR 49.90 (ab 1.1.2002), sfr 88.00, S 730.00.

GPH:1990:GPH

[Gos90] Gospel Publishing House, Springfield, MO, USA. Gospel Publishing House Typesetting Manual, 1990. 186 pp.

Goudy:1978:GTD

[Gou78] Frederic W. Goudy. Goudy’s type designs: his story and specimens: being a virtual facsimile of The Typophiles chap books XIII and XIV, a halfcentury of type design and typography. Myriade Press, New Rochelle, NY, USA, second edition, 1978. ISBN 0-918142-05-9. 285 pp. LCCN Z232.G68 G6 1978. US$25.00 (hardcover), US$14.95 (paperback). edited by Paul A. Bennett, with an introduction by J. Ben Lieberman, and a bibliography by George J. McKay.

Goossens:1994:AMJ

[GR94] M. Goossens and S. Rahtz, editors. 15th Annual meeting:

— July 1994, Santa Barbara, CA, volume 15(3) of TUGBOAT 1994. TEX Users Group, San Francisco, CA, USA, 1994. ISSN 0896-3207.

Graham:1966:CTS

[Gra66] Joseph Turnbull Graham. Computer typesetting: a select bibliography. Hertfordshire County Council Technical Library and Information Service, Hatfield, Herts., UK, 1966. 39 + iv pp. LCCN Z253.3.A12 G7. Cover title. “A HERTIS publication”. At head of title: Watford College of Technology.

Graham:1969:CTS

[Gra69] Joseph Turnbull Graham. Computer typesetting, 1966–68: a select bibliography. Hertis, Hatfield, Herts., UK, 1969. ISBN 0-85267-009-5. 4 + 60 pp. LCCN Z253.3.A12 G7 1969. Cover title. On t.p. Watford College of Technology. “A HERTIS publication”. Updated version of Computer typesetting compiled in 1966.

Graham:1975:NLC

[Gra75] Joan Graham. New look at cold comp. IEEE Transactions on Professional Communication, PC-18(3):135–138, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Grady:1986:WSA

[Gra86] William E. Grady. Work simplification approach to error reduction in printing production. Proceedings — Fall Industrial Engineering Conference (Institute of Industrial Engineers), pages 185–190, 1986. CODEN PFICEE.

Graetzer:1996:MI

[Grä96] George Grätzer. Math into LaTEX: an introduction to LaTEX2ε and AMS-LaTEX 1.2. Birkhauser Boston Inc., Cambridge, MA, USA, 1996. ISBN 0-8176-3805-9, 3-7643-3805-9. xxvii + 451 pp. LCCN Z253.4.A65G69 199. US$49.50.

Graetzer:2000:ML

[Grä00a] George Grätzer. Math into LaTEX. Birkhauser Boston Inc. and Springer-Verlag Inc., Cambridge, MA, USA and New York, NY, USA, third edition, 2000. ISBN 0-8176-4131-9, 3-7643-4131-9. xxv + 584 pp. LCCN Z253.4.L38 G745 2000. US$49.95.

Gratzer:2000:ML

[Grä00b] George Grätzer. Math into LaTEX. Birkhauser Boston Inc. and Springer-Verlag Inc., Cambridge, MA, USA and New York, NY, USA, third edition, 2000. ISBN 0-8176-4131-9, 3-7643-4131-9. xxv + 584 pp. LCCN Z253.4.L38 G745 2000. US$49.95.

Goossens:1997:LGC

[GRM97] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The LaTEX Graphics Companion: Illustrating Documents with TEX and PostScript. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85469-4. xxi + 554 pp.

LCCN Z253.4.L38G663 1997. US$39.75.

Greenberg:1970:ECS

[GS70] J. S. Greenberg and R. E. Schubert. Electronic composition system. IEEE Transactions on Aerospace and Electronic Systems, AES-6(3):270–5, May 1970. CODEN IEARAX. ISSN 0018-9251.

Garfinkel:2021:NCM

[GS21] Simson Garfinkel and Eugene H. Spafford. News: Charles M. Geschke (1939–2021). Communications of the ACM, 64(7):22, July 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3467481.

Grebert:1992:CGP

[GSKM92] Igor Grebert, David G. Stork, Ron Keesing, and Steve Mims. Connectionist generalization for production: an example from gridfont. Neural Networks, 5 (4):699–710, July–August 1992. CODEN NNETEB. ISSN 0893-6080.

GeyerSchulz:1993:TAL

[GSMN93] A. Geyer-Schulz, J. Matulka, and G. Neumann. Typesetting APL with LaTEX–the APL front end APL2TEX. APL Quote Quad, 23(3):22–27, March 1993. CODEN APLQD9. ISSN 0163-6006.

Guldin:1962:CUA

[Gul62] Mark F. Guldin. A comparison of the use and acceptance of phototypesetting with other typesetting processes among commercial publishers and university presses. Thesis (M.S.), South Dakota State University, Brookings, SD, USA, 1962. 100 pp.

Guy:20xx:VMF

[Guyxx]  Martin W. Guy.  Various

[Mergenthaler] 202 font widths and other ditroff programs and resources. Web site,

20xx. URL http://medialab.
freaknet.org/martin/tape/
stuff/ditroff/.

Ganzinger:1985:FTL

[GW85] Harald Ganzinger and Walter Willmertinger. FOAM: a twolevel approach to text formatting on a microcomputer system. Software—Practice and Experience, 15(4):327–341, April 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Haas:1986:TCS

[Haa86] Roswitha T. Haas. Typesetting chemical structure formulas with the text formatter TEX/ LaTEX. Thesis (M.S.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1986. vii + 121 pp.

Hersch:1998:EPA

[HAB98] Roger Hersch, Jacques André, and Heather Brown, editors.

Electronic publishing, artistic imaging, and digital typography: 7th International Conference on Electronic Publishing, EP ’98 held jointly with the 4th International Conference on Raster Imaging and Digital Typography, RIDT ’98, St. Malo, France, March/April 1998: proceedings, volume 1375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64298-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z250.7.I58 1998.

Hahn:1991:LER

[Hah91] Jane Hahn. LaTEX for everyone: a reference guide and tutorial for typesetting documents using a computer. Personal TEX Inc., 12 Madrona Street, Mill Valley, CA 94941, USA, second edition, 1991. ISBN 0-9631044-0-3. xi + 346 pp. LCCN Z253.4.L38 H34 1991 Permanent Reserve.

Hahn:1993:LER

[Hah93] Jane Hahn. LaTEX for Everyone —A Reference Guide and Tutorial for Typesetting Documents Using a Computer. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, second edition, 1993. ISBN 0-13-605908-2. xi + 346 pp. LCCN Z253.4.L38 H35 1993. US$19.95.

Haley:1980:PGI

[Hal80] Allan Haley. Phototypography: a guide to in-house typesetting and design. Scribner, New York, NY, USA, 1980. ISBN 0-684-16381-0. 143 pp. LCCN TR1010.H34.

Haley:1981:PGI

[Hal81] Allan Haley. Phototypography:

a guide to in-house typesetting and design. Hale, London, UK, 1981. ISBN 0-7091-9243-6, 0-684-16381-0 (U.S.). 143 pp. LCCN TR1010 .H34.

Hammond:1968:PTU

[Ham68] W. Hammond. Phototypesetting today. user’s viewpoint of “State of art”. In International Mar Information Symposium, Washington, DC, USA, Oct 31–Nov 1 1968, pages 59–64. Mar Technology Society, Washington, DC, USA, 1968.

Hamm:1975:PTA

[Ham75] Roberto Hamm. Pour une typographie arabe: contribution technique à la démocratisation de la culture arabe. Sindbad, 1-3, rue Feutrier, Paris 75018, France, 1975. ISBN ???? 193 pp. LCCN Z251.A6 H18.

Hart:1883:APT

[Har83] M. C. Hart. The amateur printer; or, typesetting at home. A complete instructor for the amateur in all the details of the printer’s art. With explanatory engravings. Dick & Fitzgerald, New York, NY, USA, 1883. 4 + ii + 5–44 + 28 pp.

Hart:1991:HRC

[Har91] Horace Hart. Hart’s Rules; For Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 39th edition, 1991. ISBN 0-19-212983-X. 182 pp. LCCN Z254 .H37 1983. UK£8.

Haralambous:1992:THQ

[Har92] Y. Haralambous. Typesetting the Holy Qur’an with TEX. In Anonymous [Ano92], page 2.1.1.

Haralambous:1994:TK

[Har94a] Y. Haralambous. Typesetting Khmer. Electronic Publishing— Origination, Dissemination, and Design, 7(4):197–??, ???? 1994. CODEN EPODEU. ISSN 0894-3982.

Haralambous:1994:THB

[Har94b] Y. Haralambous. Typesetting the holy Bible in Hebrew, with TEX. In Goossens and Rahtz [GR94], pages 174–191. ISSN 0896-3207.

Haugland:1994:EAT

[Hau94] Edward Joseph Haugland. An economic analysis of technological evolution in the typesetting industry. Thesis (M.S.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1994. 63 pp.

Haug:2022:FTI

[Hau22] Martin E. Haug. Fast typesetting with incremental compilation. Master’s thesis, Technische Universität Berlin, Faculty IV — Electrical Engineering and Computer Science, Berlin, Germany, June 2022.

Hayashi:1987:TTH

[Hay87] Tsunetoshi Hayashi. Transporting TEX into the Hitac VOS3 environment. Software —Practice and Experience, 17 (1):17–28, January 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hersch:1995:PTG

[HBBG95] Roger D. Hersch, Claude Betrisey, Justin Bur, and Andre Gurtler. Perceptually tuned generation of grayscale fonts. IEEE Computer Graphics and Applications, 15(6):78–89, November 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Holder:1988:FSI

[HD88] S. Holder and J. Dengler. Fontand size invariant character recognition with greyvalue image features. Proceedings — International Conference on Pattern Recognition, pages 252– 254, 1988. CODEN PICREG. ISBN 0-8186-0878-1. Available from IEEE Service Cent (catalog number 88CH2614-6). Piscataway, NJ, USA.

Heller:1998:EGD

[Hel98] Steven Heller, editor. The Education of a Graphic Designer. Allworth Press, New York, NY, USA, 1998. ISBN 1-880559-99-4. xiii + 273 pp. LCCN NC590.E38 1998. US$18.95.

Heller:1999:DLC

[Hel99] Steven Heller. Design Literacy (Continued): Understanding Graphic Design. Allworth Press, New York, NY, USA, 1999. ISBN 1-58115-035-0. xiv + 281 pp. LCCN NC998.H45 1999. US$19.95.

Henderson:1973:ICT

[Hen73] Ronald L. Henderson. An introduction to computer typesetting/photo-composition. Report — Dataflow Systems, inc; no.

878. Dataflow Systems, Inc., Bethesda, MD, USA, 1973. 151 pp.

Hershey:1967:CC

[Her67] Allen V. Hershey. Calligraphy for computers. Technical Report TR-2101, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, August 1967.

Hershey:1969:FIP

[Her69] Allen V. Hershey. FORTRAN IV programming for cartography and typography. Technical Report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1969. iv + 97 pp. URL https://

                                        books.google.com/books?id=
                                        _FNCAAAAIAAJ.

Hershey:1970:PRF

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:1971:APF

[Her71] Allen V. Hershey. Auxiliary programming for the Fortran typographic system. Technical Report TR-2645, U.S. Naval Weapons Laboratory, Dahlgren, VA, USA, October 1971.

Hershey:1972:CSS

[Her72] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1(3):373– 385, November 1972. CODEN CGIPBG. ISSN 0146-664X.

Hertz:1978:GCPa

[Her78a] Andrew Hertz. Graphic Communications — Preparatory Area. Book I — Typography and Modern Typesetting. Student Manual. New Jersey Vocational-Technical Curriculum Laboratory, Building 4103, Kilmer Campus, Rutgers University, New Brunswick, New Jersey 08903, January 1978. 172 pp.

Hertz:1978:GCPb

[Her78b] Andrew Hertz. Graphic Communications — Preparatory Area. Book I — Typography and Modern Typesetting. Teacher’s Manual. New Jersey Vocational-Technical Curriculum Laboratory, Building 4103, Kilmer Campus, Rutgers University, New Brunswick, New Jersey 08903, January 1978. 34 pp.

Hershey:1979:TCC

[Her79]  Allen V. Hershey.  Terres-

trial and celestial cartography. Technical Report NSWC-DL-TR-3789, U.S. Naval Surface Weapons Center, Dahlgren, VA, USA, May 1979.

Hershey:1981:ACT

[Her81] Allen V. Hershey. Advanced computer typography. Technical Report NPS012-81-005, U. S. Naval Postgraduate School, Monterey, CA 93940, USA, December 1981.

Hershey:1995:CTT

[Her95] A. V. Hershey. Cartography and typography in True BA-SIC. Technical Report NPS-09-95-003, Naval Postgraduate School, Monterey CA, USA, September 1995.

Hasegawa:1996:VTF

[HFW^{+}96] Susumu Hasegawa, Masayuki Fujisawa, Hajime Watanabe, Yoshiko Imaki, Sachiko Iida, and Daisuke Matsunaga. Variety Touch Font: a font generating system. Shapu Giho/Sharp Technical Journal, 64:27–30, April 1996. CODEN STEJD9. ISSN 0285-0362.

Handa:1983:LP

[HH83] Michael K. Handa and Jonathan A. Hunt. ‘Typesetting’ with lowcost printers. High Technology (Boston), 3(11):24, 26, 28– ??, November 1983. CODEN HTECD3.

Hutner:1999:CCF

[HK99] Martin Hutner and Jerry Kelly.

A century for the century: fine printed books from 1900 to 1999. Grolier Club, New York, NY, USA, 1999. ISBN 0-910672-29-6 (paperback), 0-910672-28-8. xxx + 110 pp. LCCN Z1033.F5 H87 1999.

Hutner:2004:CCF

[HK04] Martin Hutner and Jerry Kelly.

A century for the century: fine printed books from 1900 to 1999. Grolier Club, Danbury, CT, USA, revised and enlarged edition, 2004. ISBN 1-56792-220-1 (hardcover). lxii + 110 pp. LCCN Z1033.F5 H87 2004. URL http:

//www.loc.gov/catdir/toc/
ecip0421/2004019023.html.

Hirschberg:1987:NAF

[HL87] D. S. Hirschberg and L. L. Larmore. New applications of failure functions. Journal of the ACM, 34(3):616–625, July 1987. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0004-5411/
28875.html.

Heller:2001:TTC

[HM01] Steven Heller and Philip B.

Meggs. Texts on Type: Critical Writings on Typography. Allworth Press, New York, NY, USA, 2001. ISBN 1-58115-082-2. xi + 275 pp. LCCN Z250.T365 2001. US$19.95.

Haas:1987:TCS

[HO87] Roswitha T. Haas and Kevin C. O’Kane. Typesetting chemical structure formulas with the text formatter TEX/LaTEX. Computers and Chemistry, 11(4):251– 271, 1987. CODEN COCHDK. ISSN 0097-8485.

Hobby:1993:GAT

[Hob93] John D. Hobby. Generating automatically tuned bitmaps from outlines. Journal of the ACM, 40(1):48–94, January 1993. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Hodes:1988:PSM

[Hod88] Carol L. Hodes. Professional software for mathematical text. IEEE Potentials, 7(4), December 1988. CODEN IEPTDF. ISSN 0278-6648 (print), 1558-1772 (electronic).

Hodges:2003:GHS

[Hod03] Elaine R. S. Hodges, editor. The Guild handbook of scientific illustration. John Wiley and Sons, Inc., New York, NY, USA, second edition, 2003. ISBN 0-471-36011-2. xv + 623 pp. LCCN Q222 .G85 2003. URL http:

                                        //www.loc.gov/catdir/bios/
                                        wiley046/2002009096.htm;
                                        http://www.loc.gov/catdir/
                                        description/wiley0310/2002009096.
                                        htm;   http://www.loc.gov/
                                        catdir/toc/wiley031/2002009096.
                                        htm.

Hoenig:1998:TUL

[Hoe98] Alan Hoenig. TEX Unbound:

LaTEX and TEX Strategies for Fonts, Graphics, & More. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-509686-X (paperback), 0-19-509685-1 (hardcover). ix + 580 pp. LCCN Z253.4.L38H64 1997. US$60.00 (hardcover), US$35.00 (paperback). URL http://

www.oup-usa.org/gcdocs/gc_
0195096851.html.

Hoffman:1983:FF

[Hof83] Paul E. Hoffman. Fancy font.

BYTE Magazine, 8(10):??, October 1983. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hohti:1988:TAT

[Hoh88] Aarno Hohti. Typesetting APL with TEX. ACM SIGAPL APL Quote Quad, 18(3):13–16, March 1988.

Holmes:1984:ECP

[Hol84] Alan Holmes. Electronic composition: a practical guide to modern developments and systems in photocomposition and computer aided typesetting. Practical printing handbooks. Emblem, Northwood, 1984. ISBN 0-946513-02-3 (paperback). 124 pp. LCCN ????

Holub:1987:NCI

[Hol87a] Allen Holub. NR: a C implementation of Nroff. Dr. Dobb’s Journal of Software Tools, 12(3):

48–??, March 1987. CODEN DDJTEQ. ISSN 0888-3076.

Holub:1987:NHE

[Hol87b] Allen Holub. Nroff: Hashing, expressions, and roman numerals. Dr. Dobb’s Journal of Software Tools, 12(2):90–??, February 1987. CODEN DDJTEQ. ISSN 0888-3076.

Holkner:2006:GMO

[Hol06] Alex Holkner. Global multiple objective line breaking. Honours thesis, School of Computer Science and Information Technology, RMIT University, Melbourne, Victoria. Australia, October 2006. 35 pp. URL

http://citeseerx.ist.psu.
edu/viewdoc/download?doi=
10.1.1.585.8487&rep=rep1&
type=pdf.

Hopkins:2012:TLM

[Hop12] Richard L. (Richard Lee) Hopkins. Tolbert Lanston and the Monotype: the origin of digital typesetting, volume 21 of Typographic curiosities. University of Tampa Press, Tampa, FL, USA, 2012. ISBN 1-59732-100-1 (hardcover), 1-59732-101-X (paperback). xvii + 214 pp. LCCN Z126.5 .H67 2012.

Hossain:2024:TSR

[Hos24] Anushah Hossain. Text standards for the “rest of world”: The making of the Unicode standard and the OpenType format. IEEE Annals of the History of Computing, 46(1):20– 33, January/March 2024. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Houston:2013:SCAa

[Hou13a] Keith Houston. Shady Characters: Ampersands, Interrobangs and Other Typographical Curiosities. Particular Books, London, UK, 2013. ISBN 1-84614-647-X (hardcover), 1-84614-648-8. xi + 340 pp. LCCN P301.5.P86 H68 2013b.

Houston:2013:SCSb

[Hou13b] Keith Houston. Shady Characters: the Secret Life of Punctuation, Symbols, and Other Typographical Marks. W. W. Norton & Co., New York, NY, USA, 2013. ISBN 0-393-06442-5. xi + 340 pp. LCCN P301.5.P86 H68 2013.

Howland:1986:TAU

[How86] John E. Howland. Typesetting APL using a Macintosh. ACM SIGAPL APL Quote Quad, 16 (4):301–305, 1986. ISBN 0-901865-35-4.

Hoyle:2013:GGP

[Hoy13] John Thomas Hoyle. Going with Goudy to Philadelphia: a story of what man has wrought. Hill & Dale Private Press and Typefoundry, Terra Alta, WV, USA, 2013. 19 pp.

Huss:1973:DPM

[Hus73] Richard E. Huss. The development of printers’ mechanical typesetting methods, 1822– 1925. Published for the Bibliographical Society of the University of Virginia by the University Press of Virginia, Charlottesville, 1973. ISBN 0-8139-0336-X. 307 pp. LCCN Z253.H87.

Huss:1976:DCH

[Hus76] Richard E. Huss. Dr. Church’s “hoax”: an assessment of Dr. William Church’s typographical inventions in which is enunciated Church’s law. Graphic Crafts, Lancaster, PA, USA, 1976. ISBN ???? xiv + 78 pp. LCCN Z253 .H872. US$15.00.

Huss:1985:MET

[Hus85a] Richard E. Huss. Models of early typesetting machines at the Smithsonian Institution. Huss, Lancaster, PA, USA, 1985. vii + 32 + 24 pp.

Huss:1985:PCM

[Hus85b] Richard E. Huss. The printer’s composition matrix: a history of its origin and development. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, 1985. ISBN 0-938768-09-3. x + 56 pp. LCCN Z253 .H88 1985. US$45.00.

Hussain:1995:CRC

[Hus95] F. Hussain. On computerisation and representation of characters. IEE Colloquium (Digest), 191:11/1–11/9, 1995. CODEN DCILDN. ISSN 0963-3308.

Hartsough:1982:DPF

[HYC82] Christopher Hartsough, Yuzo Yamamoto, and E. David Callender. Documentation Production from a Formal Database. ACM Press, New York, NY, USA, 1982. ISBN 0-89791-080-X. 64–77 pp. LCCN QA76.9.D6 I57 1982. ACM (Order number 61 1821). Baltimore, MD, USA.

Hirose:1989:BAW

[HYH89] Yoshio Hirose, Koichi Yamashita, and Shimpei Hijiya. Back-propagataion algorithm which varies the number of hidden units. In International Joint Conference on Neural Networks — volume II (Jun 18–22 1989: Washington, DC, USA), pages 625–?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89CH2765-6.

Hirose:1991:BAW

[HYH91] Yoshio Hirose, Koichi Yamashita, and Shimpei Hijiya. Back-propagation algorithm which varies the number of hidden units. Neural Networks, 4(1):61–66, 1991. CO-DEN NNETEB. ISSN 0893-6080.

Ip:1987:CCW

[IBC87] Chok-Ho Ip, Daniel M. Berry, and Kam Pui Chow. Cwpr, a Chinese/Japanese Word-Processing System for Use with the UNIX Device-Independent Troff System. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0780-7. 570–577 pp. LCCN ?QA75.5 .I5761 1987. IEEE Service Cent. Piscataway, NJ, USA.

IBM:1967:ITS

[IBM67] IBM Corporation. IBM 1130 Typesetting System (RPO). IBM Corporation, San Jose, CA, USA, 1967. iv + 30 pp.

                                        URL http://bitsavers.org/
                                        pdf/ibm/1130/GA26-3531-0_
                                        1130typesetting.pdf.

ITICP:1965:ATC

[ICC65a] Automatic typesetting = Composition automatique des textes = Automatisches setzen von texten: International Technical Information Congress (1st: Paris). ????, Paris, France, 1965.

ICCT:1965:PIC

[ICC65b] Proceedings of International Conference on Computerized Typesetting, October 25 and 26, 1967, Washington, DC, USA. Research and Engineering Council of the Graphic Arts Industry, Washington, DC, USA, 1965. LCCN Z253 .I54 1965.

IEEE:1993:IPN

[IEE93] IEEE, editor. IPCC 93 proceedings: The new face of technical communication: people, processes, products: October 5–8, 1993, Philadelphia, PA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1466-2, 0-7803-1465-4, 0-7803-1467-0. LCCN T 10.5 I54 1993.

Ion:1988:TUI

[II88] Patrick D. F. Ion and Bonita D. Ross Ion. TEX and its uses for information presentation. Bulletin of the American Society for Information Science, 14(4):21– 22, April–May 1988. CODEN BASICR. ISSN 0095-4403.

Ikedo:1995:SHG

[Ike95] Tsuneo Ikedo. Scalable highperformance graphics processor: GVIP. Visual Computer, 11 (3):121–133, 1995. CODEN VI-COE5. ISSN 0178-2789 (print), 1432-2315 (electronic).

Ishii:1982:HED

[IKM82] Atsushi Ishii, Kenji Kouno, and Yasunori Maezawa. Highly efficient data compression method for newspaper image data. Fujitsu scientific and technical journal, V 18(N 2):199–225, June 1982. CODEN FUSTA4. ISSN 0016-2523.

Ing:1991:MLT

[Ing91] Roy T. Ing. MOLLY — a language for typesetting molecular structure diagrams. Computers and Chemistry, 15(3): 185–??, ???? 1991. CODEN COCHDK. ISSN 0097-8485.

ICS:1906:PVP

[Int06] International Correspondence Schools. Penmanship, vertical penmanship: elements of English grammar, elements of English composition, punctuation and capitalization, business correspondence, typesetting and proofreading, tabulating, title-page composition, civil service letter writing, report writing. International Textbook Co., Scranton, PA, USA, 1906. various pp.

Intergraphics:1982:HTT

[Int82] Intergraphics, Alexandria, VA, USA. High-tech typesetting: the typesetting service for personal computer users, 1982. 216 pp.

Intergraphics:1986:HTT

[Int86] Intergraphics, Inc, Alexandria, VA, USA. High-tech typesetting: the typesetting service for personal computer users, second edition, 1986. 4 + 203 + 1 pp.

ICNS:2007:UIR

[Int07] Interdivisional Committee on Nomenclature and Symbols. On the use of italic and roman fonts for symbols in scientific text. Report, International Union of Pure and Applied Chemistry (IUPAC), Research Triangle Park, NC 27709-3757, USA, 2007. 2

pp. URL http://old.iupac.
org/standing/idcns/fonts_
for_symbols.html;   http:
//old.iupac.org/standing/
idcns/italic-roman_dec99.
pdf.

ISO:2009:IQU

[ISO09] ISO. ISO 80000-2: Quantities and units — Part 2: Mathematical signs and symbols to be used in the natural sciences and technology. International Organization for Standardization, Geneva, Switzerland, December 2009. 40

pp. URL https://www.iso.
org/standard/31887.html.

Jacobs:1974:HEC

[Jac74] Marvin Jacobs. How to establish a cold typesetting department and train operating personnel. North American Pub. Co., Philadelphia, PA, USA, 1974. ISBN 0-912920-35-1. 184 pp. LCCN Z253.J23 1974.

Jacobs:1979:FTP

[Jac79] Marvin Jacobs. Forms, typesetting and paste-up techniques. Formsman, Cleveland, Oh., 1979. 104 pp.

Janson:1986:TTP

[Jan86] B. J. Janson. Technical typesetting: a publisher’s view. In Constance U. Greaser and Ethel G. Langlois, editors, Seeking scholarly perspectives: Proc. of the Eighth Annual Meeting of the Society for Scholarly Publishing, pages 66–68. Society for Scholarly Publishing, Washington, DC, USA, 1986. US$15.00.

Jasper:1967:ACT

[Jas67] W. P. Jasper, editor. Advances in Computer Typesetting: Proceedings of the 1966 International Computer Typesetting Conference: Sussex University, July 1966. The Institute of Printing, London, UK, 1967. xiii + 306 pp. LCCN Z253.3 .I5

Janiczek:1974:FAT

[JK74] P. M. Janiczek and George H. Kaplan. FORTRAN automatic typesetting system. United States Naval Observatory Circular; 149. U.S. Naval Observatory, Washington, DC, USA, 1974. 71 pp.

Jones:1978:BDM

[JSG78] Dan Jones, Don Stupeck, and Gerald Gallagher. Brushless DC motor font drive. In Proceedings of the Annual Symposium on Incremental Motion Control System Devices, 7th, Chicago, IL, USA, May 24–26 1978, pages 189–202. Incremental Motion Control Syst Soc, Urbana-Champaign, Ill, 1978. CODEN IMCSD3.

Jury:2002:AFR

[Jur02] David Jury. About Face: Reviving the Rules of Typography. Rotovision, Hove, East Sussex BN3 1DD, England, 2002. ISBN 2-88046-677-6. 159 pp. LCCN ???? US$40.00.

Justus:1972:TMT

[Jus72] Paul E. Justus. There is more to typesetting than setting type. IEEE Transactions on Professional Communication, PC-15 (1):13–16, March 1972. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Kadavy:2011:DHR

[Kad11]  David Kadavy.  Design for

Hackers: Reverse Engineering Beauty. John Wiley and Sons, Inc., New York, NY, USA, 2011. ISBN 1-119-99895-6. 352 (est.) pp. LCCN ????

KDADP:1987:RIE

[Kan87] Kansas Dept. of Administration.

Division of Printing and Kansas. Division of Information Systems and Communications, Topeka, KS, USA. Request for information: electronic typesetting system, 1987. 39 + 8 pp.

Kass:1972:EMC

[Kas72] Jerry Kass. Editing mathematical copy. Journal of Technical Writing and Communication, 2 (4):307–330, October 1972. CO-DEN JTWCAA.

Kato:1990:WNC

[KAT90] Hirokatsu Kato, Eiki Azuma, and Hiroshi Tada. Workstationbased networked composition system SMI/EDIAN. Sumitomo Metals, 42(3):79–100, July 1990. CODEN SUKIA6. ISSN 0371-4101.

Kerrick:1986:ERO

[KB86] David D. Kerrick and Alan C.

Bovik. Efficient recognition of omni-font characters using models of human pattern perception. In Proceedings of the 1986 IEEE International Conference on Systems, Man, and Cybernetics, Atlanta, GA, USA, pages 1287–1291. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. IEEE Service Cent. Piscataway, NJ, USA.

Kernighan:1975:STM

[KC75] Brian W. Kernighan and Lorinda L. Cherry. System for typesetting mathematics. Communications of the ACM, 18(3): 151–157, March 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kernighan:1978:TMU

[KC78] Brian W. Kernighan and Lorinda L. Cherry. Typesetting mathematics: user’s guide. Technical report, University of California, Berkeley, Computing Services, Berkeley, CA, USA, August 15, 1978. 10 pp.

Kernighan:1980:STM

[KC80] Brian W. Kernighan and Lorinda L. Cherry. A system for typesetting mathematics. Technical report, University of California, Berkeley, Computing Services, Berkeley, CA, USA, 1980. 8 pp. Revision of 1975 paper.

Kraus:1994:SMC

[KD94] Eugene J. Kraus and Edward R. Dougherty. Segmentation-free morphological character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 2181:14– 23, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Keeler:1980:SCA

[Kee80] F. Laurence Keeler. System for Computer Automated Typesetting (SCAT) of computer authored texts. Technical Report TAEG-88, Training Analysis and Evaluation Group, System for Computer Automated Typesetting, Chief of Naval Education and Training, Orlando, FL, USA, 1980. 151 pp.

Keith:1983:MT

[Kei83] Michael Keith. Microcomputer typography. Proceedings — Symposium on Small Computers in the Arts, pages 99–102, 1983. CODEN PSSAED. ISBN 0-8186-0499-9. IEEE Service Cent. Piscataway, NJ, USA.

Keller:1985:TA

[Kel85] Arthur M. Keller. Typesetting by authors. In Lucarella [Luc85], pages 1–14. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Kelly:2010:ABT

[Kel10] Jerry Kelly. The art of the book in the twentieth century: a study of eleven influential book designers from 1900 to 2000. RIT Cary Graphic Arts Press, Rochester, NY, USA, 2010. ISBN 1-933360-46-1. xviii + 179 pp. LCCN Z246 .K45 2010.

Kelly:2011:AMA

[Kel11] Jerry Kelly. About more alphabets: the types of Hermann Zapf, volume 3 of Typophile chap book. N.S. The Typophiles, New York, NY, USA, 2011. ISBN 0-9842744-0-5. 111 pp. LCCN ???? With a foreword by Robert Bringhurst.

Kelly:2019:HZW

[Kel19] Jerry Kelly. Hermann Zapf and the world he designed. The Kelly-Winterton Press, New York, NY, USA, 2019. ISBN 0-578-42521-1 (hardcover). 364 pp. LCCN Z250.A2 Z3754 2019.

Kernighan:1981:PLT

[Ker81] B. W. Kernighan. PIC: a language for typesetting graphics. In Abrahams [Abr81], pages 92– 96. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6.

Kernighan:1982:PGL

[Ker82a] Brian W. Kernighan. PIC— a graphics language for typesetting: user manual. Computing science technical report; no. 85. Bell Laboratories, Murray Hill, NJ, USA, revised edition, March 1982. 24 pp.

                                        URL http://cm.bell-labs.
                                        com/cm/cs/cstr/116.ps.gz;
                                         http://www.cs.bell-labs.
                                        com/cm/cs/cstr/116.ps.gz;
                                         http://www.maht0x0r.net/
                                        library/computing/pic.pdf.

Kernighan:1982:PLT

[Ker82b] Brian W. Kernighan. PIC: a language for typesetting graphics. Software—Practice and Experience, 12(1):1–21, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kernighan:1982:TIT

[Ker82c] Brian W. Kernighan. A typesetter independent TROFF. Computing Science Technical Report 97, Bell Laboratories, Murray Hill, NJ 07974, USA, March 1982.

Khoubyari:1994:FIU

[KH94] Siamak Khoubyari and Jonathan J.

Hull. Font identification using visual global context. Proceedings of SPIE — The International Society for Optical Engineering, 2181:116–124, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Khoubyari:1996:FFW

[KH96] Siamak Khoubyari and Jonathan J.

Hull. Font and function word identification in document recognition. CVIU: Computer Vision and Image Understanding, 63(1):66–74, January 1996. CO-DEN CVIUF4.

Khazaeli:2000:CTL

[Kha00] Cyrus Dominik Khazaeli. Crashkurs Typo und Layout: vom Zeilenfall zum Screendesign. Rowohl, Reinbek bei Hamburg, Germany, 2000. ISBN 3-499-19815-0. 383 pp. LCCN ????

Killmon:1983:ETI

[Kil83] Mark C. Killmon. Ergonomics for the typesetting industry. National Composition Association, Arlington, VA, USA, 1983. 55 pp.

Kimball:1980:DAC

[Kim80] Glenn S. Kimball. Design of an advanced cathode-ray tube (CRT) phototypesetter. Proceedings of the Society of Photo-Optical Instrumentation Engineers, 250(??):8–13, ???? 1980. CODEN SPIECJ. ISSN 0361-0748.

Kim:1990:DSH

[Kim90]  Kyongsok Kim.  Databases

supporting Hangul (or Korean script). Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, pages 485–490, November 1990. CODEN PICYE3. ISBN 0-87942-597-0. ISSN 0884-3627. IEEE catalog number 90CH2930-6.

Kingston:1991:ELP

[Kin91] Jeffrey H. Kingston. Eq: a Lout package for typesetting mathematics. Number 410 in Technical report / University of Sydney. Basser Dept. of Computer Science. University of Sydney, Sydney, NSW, Australia, 1991. 16 pp.

Kingston:1992:ELP

[Kin92] Jeffrey H. Kingston. Eq: a Lout package for typesetting mathematics. Technical report / University of Sydney. Basser Dept. of Computer Science; 452 Technical report (University of Sydney. Basser Dept. of Computer Science); 452. University of Sydney, Sydney, NSW, Australia, December 22, 1992. 15 pp.

Kingston:1993:DIL

[Kin93] Jeffrey H. Kingston. Design and implementation of the Lout document formatting language. Software—Practice and Experience, 23(9):1001–1041, September 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kinross:2002:UTP

[Kin02]  Robin Kinross.  Unjustified

Texts: Perspectives on Typography. Hyphen Press, London, UK, 2002. ISBN 0-907259-17-0. 391 pp. LCCN ???? US$30.00.

Kirschenbaum:2005:GGH

[Kir05] Valerie Kirschenbaum. Goodbye Gutenberg: hello to a new generation of readers and writers. Designer Writers series. Global Renaissance Society, New York, NY, USA, 2005. ISBN 0-9745750-3-8. 416 pp. LCCN Z116.A3 K57 2005. URL http:/

/www.goodbyegutenberg.com/.

Kirschenbaum:2016:TCL

[Kir16] Matthew G. Kirschenbaum.

Track changes: a literary history of word processing. The Belknap Press of Harvard University Press, Cambridge, MA, USA, 2016. ISBN 0-674-41707-0. xvi + 344 + 16 pp. LCCN Z52.4 .K57

2016. URL http://www.hup.
harvard.edu/catalog.php?isbn=
9780674417076.

Kitamura:1984:DFG

[Kit84] J. Kitamura. Draft font generation. IBM Technical Disclosure Bulletin, 27(4B):2504–??, September 1984. CODEN IBM-TAA. ISSN 0018-8689.

Kawata:1993:OFR

[KKM^{+}93] Tetsuro Kawata, Kenich Kawachi, Nobuaki Miyakawa, Ichiro Kawazome, Hiromi Ysumatsu, Susumu Haga, and Masaya Takenaka. Outline font rendering processor with an embedded RISC CPU for high-speed hint processing. Proceedings of the Custom Integrated Circuits Conference, pages 12.5.1– 12.5.4, 1993. CODEN PCICER. ISBN 0-7803-0826-3. ISSN 0886-5930. IEEE catalog number 93CH3214-4.

Kawata:1994:OFR

[KKM^{+}94] Tesuro Kawata, Kenichi Kawauchi, Nobuaki Miyakawa, Ichiro Kawazome, Hiromi Yasumatsu, Susmumu Haga, and Masaya Takenaka. Outline font rendering processor with an embedded RISC CPU for high-speed hint processing. IEEE Journal of Solid-State Circuits, 29 (3):280–289, March 1994. CO-DEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic).

Kernighan:1982:UDP

[KL82] B. W. Kernighan and M. E. Lesk. UNIX document preparation. In Nievergelt et al. [NCNS82], pages 1–20. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Kleper:1976:UPT

[Kle76] Michael L. Kleper. Understanding photo typesetting. North American Pub. Co., Philadelphia, PA, USA, 1976. ISBN 0-912920-46-7. ix + 166 pp. LCCN Z253.2.K58 1976.

Kleper:1978:TTH

[Kle78] Michael L. Kleper. A touch of typesetting history. Graphic Dimensions, Rochester, NY, USA, 1978. 1 pp.

Kleper:1979:HBB

[Kle79] Michael L. Kleper. How to build a basic typesetting system. Graphic Arts Research Center, Rochester Institute of Technology: available from Graphic Dimensions, Rochester, NY, USA, 1979. ISBN 0-89938-002-6. 90 pp. LCCN TR1010.K58.

Kleper:1982:TM

[Kle82] Michael L. Kleper. Typesetting by microcomputer. Graphic Dimensions, Pittsford, NY, USA, 1982. 22 pp.

Kleper:1987:IHD

[Kle87] Michael L. Kleper. The illustrated handbook of desktop publishing and typesetting. TAB Professional and Reference Books, Blue Ridge Summit, PA, USA, 1987. ISBN 0-8306-2700-6, 0-8306-0700-5 (paperback). xiv + 770 pp. LCCN Z286.D47K55 1987.

Kleper:1990:IHD

[Kle90] Michael L. Kleper. The illustrated handbook of desktop publishing and typesetting. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, second edition, 1990. ISBN 0-8306-3350-2 (paperback). 927 pp. LCCN Z286.D47K55 1990.

Kobayashi:1987:MHL

[KM87] Makoto Kobayashi and Yoshiaki Mima. Multi-Font Handling for Large Character Set on Workstations. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0770-X. 2–8 pp. LCCN HF5548 .I44 1987. IEEE Service Cent. Piscataway, NJ, USA.

Kai:1989:HSO

[KMNO89] Naoyuki Kai, Tsutomu Minagawa, Ichiro Nagashima, and Masahide Ohhashi. High speed outline font rasterizing LSI. Proceedings of the Custom Integrated Circuits Conference, pages 24.6/1–4, May 1989. CO-DEN PCICER. ISSN 0886-5930. Available from IEEE Service Cent. Piscataway.

Kume:1990:LDP

[KNS90] S. Kume, Y. Nishio, and Y. Satoh. Laser direct prepress system. Proceedings of SPIE — The International Society for Optical Engineering, 1252:234– 239, 1990. CODEN PSISDG. ISBN 0-8194-0299-0. ISSN 0277-786X (print), 1996-756X (electronic).

Knuth:1979:TMN

[Knu79] Donald E. Knuth. TEX and : New Directions in Typesetting. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1979. ISBN 0-932376-02-9. xi + 201 + 105 pp. LCCN Z253.3 .K58 1979.

Knuth:1984:LP

[Knu84a] D. Knuth. Literate programming. The Computer Journal, 27 (2):97–111, 1984. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Knuth:1984:TB

[Knu84b] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-13447-0 (hardcover), 0-201-13448-9 (paperback). ix + 483 pp. LCCN Z253.4.T47 K58 1986. US$15.95 (paperback), US$32.95 (hardcover). Second printing, revised, October 1984. Sixth printing, revised, January 1986; also published as Computers & Typesetting, Vol. A. Twenty-fourth printing, summer 1993, contains final revisions. Twenty-first printing, June 1992.

Knuth:1986:MP

[Knu86a] Donald E. Knuth. - : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:1986:TP

[Knu86b] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986.

Knuth:1986:MB

[Knu86c]  Donald E. Knuth. The  -

book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58 1986.

Knuth:1986:CMT

[Knu86d] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:1986:CT

[Knu86e] Donald Ervin Knuth. Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13447-0 (v.A), 0-201-13437-3 (v.B), 0-201-13445-4 (v.C), 0-201-13438-1 (v.D), 0-201-13446-2 (v.E). LCCN Z253.4.T47K58 1986.

Knuth:1986:CTE

[Knu86f] Donald Ervin Knuth. Computers and typesetting: errata and changes as of 25 September 1986. TEX Users Group, San Francisco, CA, USA, 1986. 9 pp.

Knuth:1986:M

[Knu86g] Donald Ervin Knuth. The METAFONTbook, volume C of Computers and typesetting.

Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hard), 0-201-13444-6 (soft). xi + 361 pp. LCCN Z250.8.M46K58 1986.

Knuth:1987:MP

[Knu87] Donald Ervin Knuth. META-FONT: the program. Computers and typesetting; D Knuth, Donald Ervin, 1938- Computers and typesetting; D. Addison-Wesley, Reading, MA, USA, reprinted with corrections. edition, 1987. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46K578 1986.

Knuth:1988:TP

[Knu88] Donald Ervin Knuth. TEX: the program, volume B of Computers and typesetting. Addison-Wesley, Reading, MA, USA, reprinted with corrections may 1988. edition, 1988. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47K578 1986.

Knuth:1989:CTE

[Knu89] Donald E. Knuth. Computers and typesetting: errata and changes as of 11 October 1989. Computers and typesetting. TEX Users Group, San Francisco, CA, USA, 1989. various pp.

Knuth:1991:TP

[Knu91a] Donald Ervin Knuth. TEX: the program, volume B of Computers and typesetting. Addison-Wesley, Reading, MA, USA, 4th printing, rev. edition, 1991. ISBN 0-201-13437-3. xvii + 600 pp. LCCN Z253.4.T47K578 1986.

Knuth:1991:T

[Knu91b] Donald Ervin Knuth. The TEXbook, volume A of Computers and typesetting. Addison-Wesley, Reading, MA, USA, 11th printing, rev. edition, 1991. ix + 483 pp.

Knuth:1991:TTL

[Knu91c] Donald Ervin Knuth. Textures typesetting and layout for the Apple Macintosh, 1991. 6 computer disks + 2 manuals.

Knuth:1993:TP

[Knu93a] Donald Ervin Knuth. TEX: the program, volume B of Computers and typesetting. Addison-Wesley, Reading, MA, USA, 1993 corrected edition, 1993. ISBN 0-201-13437-3. xvii + 600 pp. LCCN Z253.4.T47K578 1986.

Knuth:1993:T

[Knu93b] Donald Ervin Knuth. The TEXbook, volume A of Computers and typesetting. Addison-Wesley, Reading, MA, USA, 23rd printing, with corrections. edition, 1993. ISBN 0-201-13448-9 (paperback). ix + 483 pp. LCCN Z253.4.T47K58 1984.

Knuth:1999:DT

[Knu99] Donald E. Knuth. Digital Typography, volume 78 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 1999. ISBN 1-57586-010-4 (paperback), 1-57586-011-2 (hardcover). xvi + 685 pp. LCCN Z249.3.K59 1998. US$90.00 (cloth), US$39.95 (paperback).

Kokaji:1987:MPN

[Kok87] Norio Kokaji. Magnetic printer for newspaper CTS. IEEE Transactions on Magnetics, MAG-23(5):2889–2891, September 1987. CODEN IEMGAQ. ISSN 0018-9464.

Kopec:1993:LFM

[Kop93] Gary E. Kopec. Least-squares font metric estimation from images. IEEE Transactions of Image Processing, 2(4):510–519, October 1993. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Korbuly:1975:NAC

[Kor75]  Dorothy K. Korbuly.  New

approach to coding displayed mathematics for photocomposition. IEEE Transactions on Professional Communication, PC-18(3):283–287, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Knuth:1981:BPL

[KP81] D. E. Knuth and M. E. Plass.

Breaking paragraphs into lines. Software—Practice and Experience, 11(11):1119–1184, November 1981. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khan:1987:RPC

[KPB87] Simon Khan, Theo Pavlidis, and Henry S. Baird. On the recognition of printed characters of any font and size. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-9 (2):274–288, March 1987. CO-DEN ITPIDJ. ISSN 0162-8828.

Kendall:1985:IOC

[KR85] M. Kendall and A. Reed. Introduction of an on-line computer typesetting service. University Computing, 7(3):134–138, Winter 1985. CODEN UNCOET. ISSN 0265-4385.

Krantz:1996:PMW

[Kra96] Steven G. Krantz. A Primer of Mathematical Writing: Being a disquisition on having your ideas recorded, typeset, published, read & appreciated. American Mathematical Society, Providence, RI, USA, 1996. ISBN 0-8218-0635-1. 216 pp. LCCN QA42.K73 1996. US$19.00. URL http:/

/www.ams.org/bookstore. In-

cludes discussions of computers and TEX.

Krantz:2001:HTM

[Kra01] Steven G. Krantz. Handbook of Typography for the Mathematical Sciences. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-149-6. xv + 173 pp. LCCN Z253.4.T47 K65 2001.

Kronenberg:1967:CTC

[Kro67] Shirley Kronenberg. Cold type composition; equipment and technique, a handbook of typesetting and layout. Shirley Kronenberg, Stamford, CT, USA, 1967. 138 pp.

Krtolica:1994:BCA

[Krt94] Radovan V. Krtolica. Box connectivity approach to multifont character recognition. Proceedings of SPIE — The International Society for Optical Engineering, 2181:315–321, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic).

Kellerman:1987:TPT

[KS87] David Kellerman and Barry Smith. TEXTURES: professional typesetting for the Macintosh: user’s guide. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-17221-6. 123 pp. LCCN Z253.4.T47U85 1987.

Kunkel:1979:CTP

[Kun79] George Z. Kunkel. Copyflow:

typesetting procedures for book composition. National Composition Association, Arlington, VA, USA, 1979. 82 pp.

Kurlansky:2017:PPT

[Kur17] Mark Kurlansky. Paper: paging through history. W. W. Norton & Co., New York, NY, USA, 2017. ISBN 0-393-23961-6 (hardcover), 0-393-35370-2 (paperback). xx + 389 pp. LCCN TS1090 .K87 2017.

Kernighan:1989:PMP

[KV89] Brian W. Kernighan and Christopher J. Van Wyk. Page makeup by postprocessing text formatter output. Computing Systems, 2 (2):103–132, Spring 1989. CO-DEN CMSYE2. ISSN 0895-6340.

Kinross:2002:KMP

[KVM02] Robin Kinross, Jaap Van Triest, and Karel Martens, editors. Karel Martens: printed matter/ drukwerk. Hyphen Press, London, UK, second edition, 2002. ISBN 0-907259-20-0. 168 pp. LCCN ???? US$40. Translations by Françoise Berserik and Harry Lake, and text by Hugues Boekraad.

Labuz:1984:IDB

[LA84] Ronald Labuz and Paul Altimonte. The interface data book for word processing/typesetting, 1984- 1985. The Bowker graphics library. Bowker’s composition series. Bowker, New York, NY, USA, 1984. ISBN 0-8352-1908-9 (paperback). xi + 195 pp. LCCN Z52.4.L33 1984.

Labuz:1985:IDB

[LA85] The Interface data book for word processing/typesetting, page various, 1985. R.R. Bowker Co., New York, NY, USA.

Labuz:1984:HTW

[Lab84] Ronald Labuz. How to Typeset from a Word Processor: a Practical Manual for Type Buyers and Typesetters. R. R. Bowker Company, New York, NY, USA, 1984. ISBN 0-8352-1899-6. xii + 218 pp. LCCN Z253.3.L33 1984. US$24.95.

Labuz:1988:TTT

[Lab88] Ronald Labuz. Typography and typesetting: Type design and manipulation using today’s technology. Van Nostrand Reinhold Co., New York, NY, USA, 1988. ISBN 0-442-25966-2. 272 pp. LCCN Z250.L13 1988.

Labovitz:2017:ETO

[Lab17] John Labovitz. The electric typesetter: The origins of computing in typography. Printing History (New Series), (21):??, Winter 2017. ISSN 0192-9275. URL https://

printinghistory.org/publications/
printing-history/ns/#21.

Lagally:1992:AST

[Lag92a] K. Lagally. ArabTeX: a system for typesetting Arabic. In Anonymous [Ano92], page 9.4.1.

Lagally:1992:ATA

[Lag92b] Klaus Lagally. ArabTEX — typesetting Arabic with vowels and ligatures. In Zlatuška [Zla92], pages 153–172. ISBN 80-210-0480-0.

Lagus:1995:APG

[Lag95] Krista Lagus. Automated pagination of the generalized newspaper using simulated annealing. Master’s thesis, Helsinki University of Technology, Helsinki, Finland, 1995.

Lamport:1986:LDP

[Lam86]  Leslie Lamport.  LaTEX: a

Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-15790-X. xiv + 242 pp. LCCN Z253.4.L38 L35 1986.

Lammarsch:1993:NTS

[Lam93] J. Lammarsch. A new typesetting system: is it really necessary? In Anonymous [Ano93a], pages 167–170. ISSN 0896-3207.

Lamport:1994:LDP

[Lam94]  Leslie Lamport.  LaTEX: a

Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-52983-1. xvi + 272 pp. LCCN Z253.4.L38L35 1994.

Bell:1940:STM

[Lan40] Lanston Monotype Machine Co., Philadelphia, PA, USA. Specimen of types made by John Bell, now available for composition on the Monotype typesetting machine, 1940. 15 pp.

LMMC:1949:DFM

[Lan49] Lanston Monotype Machine Company, Philadelphia, PA, USA. Display faces on the Monotype; one-line specimens of the Monotype faces for which matrices are made to cast type to be used in hand composition, with a listing of sizes available for both machine and hand typesetting, 1949. 30 pp.

Lansburgh:1964:SST

[Lan64] W. N. Lansburgh. Sättningsregler.

(Swedish) [Typesetting rules]. Almqvist & Wiksells, Uppsala, Sweden, second edition, 1964. viii + 475 pp.

Landau:1971:WCC

[Lan71] Robert M. Landau, editor.

Workshop on Computer Composition, ASIS, Potomac Valley Chapter, Proceedings, Washington, DC, December 8–9, 1970, Hydrocarbon Processing. American Society for Information Science, Medford, NJ, 1971. ISBN 0-87715-002-8. LCCN Z253.3.A2 1970.

Lang:1988:JMT

[Lan88] Laura Lang. Just my type. ID Systems, 8(6):21–22, 25, July– August 1988. CODEN IDSYE5. ISSN 0892-676X.

Lau:1990:SUI

[Lau90] David Chik-Wai Lau. A semiinteractive user interface for the TEX typesetting system. Thesis (M.S. in Computer Science), University of Texas at Austin, Austin, TX, USA, 1990. viii + 578 pp.

Liebson:1974:CTT

[LB74] Laurence S. Liebson and Albert L. Bushkoff. Computerized text-editing and typesetting make headlines. Electronics, 47 (11):111–118, May 1974. CO-DEN ELECAD. ISSN 0883-4989.

Lee:1995:SQP

[LBJ95] Matthew Lee, Ben-Zion Barta, and Peter Juliff, editors. Software quality and productivity: theory, practice, education and training: Conference — 1995. Chapman and Hall, Ltd., London, UK, 1995. ISBN 0-412-62960-7. LCCN QA76.76.D47 S665 1995.

LBSC:1985:LTY

[LBS85] LePrint: typesetting for your dot-matrix or laser printer, 1985. 1 program file (IBM PC) on 3 computer disks manual.

Logar:1994:PCC

[LCO94] Antonette M. Logar, Edward M. Corwin, and William J. B. Oldham. Performance comparisons of classification techniques for multi-font character recognition. International Journal of Human Computer Studies, 40(3):403– 423, March 1994. ISSN 1071-5819.

LeMaire:1988:SCA

[Le 88] Lisa Le Maire. A study of costs and alternatives for typesetting at Visible Computer Supply Corporation. Technical report, Aurora University, Aurora, IL, USA, 1988. 37 pp.

Levine:1984:TDI

[Lev84] Charles Joseph Levine. Typesetting digital images on a Compugraphics-8400. Thesis (B.S.), Department of Electrical Engineering and Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, 1984. 69 pp. Supervised by Ronald L. MacNeil.

Levien:1987:BPT

[Lev87] Raphael L. Levien. BYSO PRINT tutorial: a guide to code based typesetting on a PC. Levien Instrument Co., McDowell, VA, USA, 1987. vii + 113 pp.

Levelt:1995:IPI

[Lev95] A. H. M. Levelt, editor. IS-SAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montreal, Canada, ISSAC - PROCEEDINGS- 1995. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Liao:1991:FGB

[LH91] Chia-Wei Liao and Jun S.

Huang. Font generation by beta-spline curve. Computers & Graphics (Pergamon), 15 (4):527–534, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lieberman:1977:TT

[Lie77] J. Ben Lieberman. Type and typefaces. Myriade Press, New Rochelle, NY, USA, second edition, 1977. ISBN 0-918142-01-6 (hardcover), 0-918142-02-4 (paperback). 142 pp. LCCN Z250 .L6 1977. US$30.00 (hardcover), US$19.95 (paperback).

LPC:1930:SPA

[Lin30] Linotype Parts Corporation. Star parts and accessories for typesetting machines, composing room supplies. The Company, New York, NY, USA, 1930. 76 pp.

LSC:1939:MTM

[Lin39] Linotype Supply Company, New York, NY, USA. To maintain typesetting machines efficiently and economically, 1939. 8 pp.

Lingelbach:1952:FAI

[Lin52] William E. Lingelbach. Franklin’s “American instructor”. Early Americanism in the art of writing. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 96(4):367– 387, August 1952. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

                                        URL https://www.jstor.org/
                                        stable/3143795.

Lindegren:1982:ALP

[Lin82] Erik Lindegren. ABC of lettering and printing typefaces: a complete guide to the letters and typefaces used for typesetting and printing. Greenwich House: Distributed by Crown Publishers, New York, NY, USA, 1982. ISBN 0-517-38334-9. 348 pp. LCCN Z250.L627 1982.

Lincoln:1990:SCT

[Lin90] Harry B. Lincoln. A study of computer techniques for music research. Report Project number 9-B-101, Department of Music, State University of New York at Binghamton, Binghamton, NY 13901, June 30, 1990.

URL https://books.google.
com/books?id=B9AXAQAAIAAJ.

Lin:1995:RAC

[Lin95] Zhen Yong Lin. Relational attribute code representation for printed multi-font Chinese character recognition. Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, 2:1924–1927, 1995. CODEN PICYE3. ISSN 0884-3627. IEEE catalog number 95CB35767.

LSC:19xx:LPA

[Linxx] Linotype Supply Company, New York, NY, USA. Linsco parts and accessories for typesetting machines, 19xx. 71 pp.

Lippold:1978:ARG

[Lip78] William Lippold. The ABP reference guide on copy processing and typesetting. American Business Press, New York, NY, USA, 1978. 38 pp.

Lipkin:1999:LLV

[Lip99] Bernice Sacks Lipkin. LaTEX for Linux: a vade mecum. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 0-387-98708-8 (softcover). xxxi + 568 pp.

LCCN Z 253.4 L38 L56 1999. US$49.95.

Lane:1990:FDM

[LJ90] Ramon Lane and Richard John. Forty dot/millimeter resistive ribbon thermal print head. Proceedings of SPIE — The International Society for Optical Engineering, 1252:168–177, 1990. CODEN PSISDG. ISBN 0-8194-0299-0. ISSN 0277-786X (print), 1996-756X (electronic).

Lesk:1977:CTT

[LK77] M. E. Lesk and B. W. Kernighan. Computer typesetting of technical journals on UNIX. AFIPS Conference Proceedings, 46(??):879–888, ???? 1977. CODEN AFPGBT.

Lau:1993:HMA

[LK93] J. Lau and P. Karow. hz-Program: Micro-typography for advanced typesetting. In Morreale [Mor93], pages 58–60. ISSN 0097-966X (print), 2154-6738 (electronic), 2154-6746.

Lim:1995:OCF

[LK95] Soon-Bum Lim and Myung-Soo Kim. Oriental character font design by a structured composition of stroke elements. Computer Aided Design, 27(3):193–207, March 1995. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

Lee:1995:CWT

[LLW95] J. H. M. Lee, J. C. K. Leung, and C. C. K. Wong. Construction of a WYSIWYG LaTEX typesetting system using object-oriented design. In Lee et al. [LBJ95], pages 135–138. ISBN 0-412-62960-7. LCCN QA76.76.D47 S665 1995.

Lacombe:2021:CYA

[LMT^{+}21] Guilhem Lacombe, Kseniia Masalygina, Anass Tahiri, Carole Adam, and Cédric Lauradoux. Can you accept LaTEX files from strangers? Ten years later. arXiv.org, ??(??):1–10, February 1, 2021. URL https:/

/arxiv.org/abs/2102.00856.

Lee:1990:FRN

[LO90] Ming-Chih Lee and William J. B. Oldham. Font recognition by a neural network. International Journal of Man-Machine Studies, 33(1):41–61, July 1990. CODEN IJMMBC. ISSN 0020-7373.

Lopez:1994:MVM

[Lop94] Robert J. Lopez, editor. Maple V: mathematics and its application: Maple summer workshop and symposium — August 1994, Troy, NY. Birkhauser Boston Inc., Cambridge, MA, USA, 1994. ISBN 3-7643-3791-5. LCCN QA76.95.M36 1994.

Lorenzen:2019:VS

[Lor19] Lee Lorenzen. The Ventura story. IEEE Annals of the History of Computing, 41(3):75–88, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Lowood:1994:LTT

[LR94] Henry F. Lowood and Robin E. Rider. Literary technology and typographic culture: The instrument of print in early modern science. Perspectives on Science, 2(1):1–37, Spring 1994. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

Luvisetto:1992:LER

[LU92] M. L. Luvisetto and E. Ugolini. LaTEX: easy recipes. Technical Report INFN/TC-92/05, Ist. Naz. Fis. Nucl., Bologna, Italy, March 1992. 17 pp.

Lucarella:1985:PFE

[Luc85] Dario Lucarella, editor. Proceedings of the First European Conference on TEX for Scientific Documentation, 16–17 May 1985, Como, Italy. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

LCC:1989:ILF

[Lud89] Ludlow Composition Company, Los Angeles, CA, USA. Intertype and Linotype faces, 1989. 33 pp.

Lunde:1999:CIP

[Lun99] Ken Lunde. CJKV Information Processing: Chinese, Japanese, Korean & Vietnamese Computing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-224-7. 1174 pp. LCCN ???? US$64.95. URL http:

//www.oreilly.com/catalog/
cjkvinfo/.

Lundmark:2002:QQS

[Lun02] Torbjörn Lundmark. Quirky Qwerty: The Story of the Keyboard Your Fingertips. New South Wales University Press Ltd., Sydney, Australia, 2002. ISBN 0-86840-436-5. 172 pp. LCCN ???? US$19.95.

Li:1986:NAR

[LZ86] Bin Li and Shuxiang Zhao.

New approach to recognition of both handwritten and multi-font printed Chinese characters. Proceedings — International Conference on Pattern Recognition, pages 641–643, 1986. CODEN PICREG. ISBN 0-8186-0742-4. IEEE Service Cent. Piscataway, NJ, USA.

Mack:1975:LCC

[Mac75] Paul F. Mack. Lower composition costs through optical scanning and photocomposition. IEEE Transactions on Professional Communication, PC-18 (3):279–282, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

MacKay:1990:ITS

[Mac90] Pierre A. MacKay. The internationalization of TEX with special reference to Arabic. Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, pages 481–484, November 1990. CO-DEN PICYE3. ISBN 0-87942-597-0. ISSN 0884-3627. IEEE catalog number 90CH2930-6.

Macrakis:1996:GLT

[Mac96] Michael S. Macrakis, editor. Greek letters: from tablets to pixels: Proceedings of an international symposium held at the Institut français d’Athènes, Athens, June 7–10, 1995. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, 1996. ISBN 1-884718-27-2. LCCN Z251.G7 G74 1996. US$49.95.

Mandry:1987:SVP

[Man87] Thomas Mandry. Schriftsatz — Vorbereitung und Produktion von Buechern. Informationstechnik it: Computer, Systeme, Anwendungen, 29(1):24– 26, 1987. CODEN ICSAES.

Martin:1972:PDC

[Mar72] David Martin. Progress in documentation: Computer typesetting and information. Journal of Documentation, 28(3):247–261, September 1972. CODEN JDO-CAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Marquart:1983:MC

[Mar83]  R. Gary Marquart.  Mak-

ing of Chemlaw. Proceedings — National Online Meeting, pages 359–366, 1983. CODEN PNOMDR. ISBN 0-938734-05-9. ISSN 0739-1471.

Marovac:1987:HFE

[Mar87] Nenad Marovac. Handling fonts in electronic publishing systems. Computers and Graphics, 11 (3):289–295, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marshall:1991:OFOa

[Mar91a]  Gerald F. Marshall.  Opti-

cal feedback overcomes scanline wobble in laser typesetting. Laser focus world, 27(4):167–??, April 1991. CODEN LFWOE8. ISSN 1043-8092.

Marshall:1991:OFOb

[Mar91b] Gerald F. Marshall. Optical feedback overcomes scanline wobble in laser typesetting. Laser Focus (Littleton, Massachusetts), 27(5), May 1991. CODEN LAFOEO. ISSN 0740-2511.

Mast:1984:FHM

[Mas84]  Claudia Mast.  Farewell to

hot metal — new technology changes newspaper editing. Siemens Review, 51(1):15–17, January–February 1984. CO-DEN SZTEA6. ISSN 0302-2528.

Masterson:2005:BDP

[Mas05] Pete Masterson. Book design and production: a guide for authors and publishers. Aeonix Publishing Group, El Sobrante, CA, USA, 2005. ISBN 0-9669819-0-1 (paperback). 320 pp. LCCN Z116.A3 M37

                                        2005.  URL http://www.
                                        loc.gov/catdir/toc/fy0711/
                                        2005901925.html.

MAVCC:1982:TP

[MAV82] Typesetting procedures, 1982. 25 slides mono. + 1 script.

Mayer:1969:TCT

[May69] Peter C. Mayer. Technical change in the typesetting of daily newspapers. Thesis (Ph.D. in economics), Dept. of Economics, University of California, Berkeley, Berkeley, CA, USA, 1969. vi + 168 pp.

Muller-Brockmann:1996:GSG

[MB96] Josef Müller-Brockmann. Grid systems in graphic design: a visual communication manual for graphic designers, typographers, and three dimensional designers = Raster-Systeme für die visuelle Gestaltung. Niggli, Sulgen, Switzerland, fourth edition, 1996. ISBN 3-7212-0145-0. 176 pp. LCCN ???? US$78.00. English version by Dennis Q. Stephenson.

Maddox:1977:FCM

[MBG77] Michael E. Maddox, James T. Burnette, and James C. Gutmann. Font comparisons for 5×7 dot matrix characters. Human Factors, 19(1):89–93, February 1977. CODEN HUFAA6. ISSN 0018-7208.

Morris:1975:ACT

[MC75a] Robert Morris and Lorinda L. Cherry. Appendix A. Common technical English words, and Appendix B. Possible typo’s and spelling errors. IEEE Transactions on Professional Communication, PC-18(1):57–64, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See [MC75b].

Morris:1975:CDT

[MC75b] Robert Morris and Lorinda L.

Cherry. Computer detection of typographical errors. IEEE Transactions on Professional Communication, PC-18 (1):54–56, March 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See supplemental word list [MC75a].

McCord:1984:FS

[McC84] E. McCord. Font substitution. IBM Technical Disclosure Bulletin, 27(1B):547–548, June 1984. CODEN IBMTAA. ISSN 0018-8689.

McClelland:1988:LTD

[McC88] Bruce McClelland. From logon to typeset document in under ten minutes. Online, 12(6):66– 71, November 1988. CODEN ONLIDN. ISSN 0146-5422.

McCarthy:2020:DTS

[McC20] Steven McCarthy. Digital typography at Stanford. She Ji: The Journal of Design, Economics, and Innovation, 6(4): 546–560, Winter 2020. ISSN 2405-8726.

McGrew:1986:AMT

[McG86] Mac McGrew. American metal typefaces of the Twentieth Century. Myriade Press, New Rochelle, NY, USA, 1986. ISBN 0-918142-20-2. xvi + 395 pp. LCCN Z250 .M34 1986.

McGrew:1993:AMT

[McG93] Mac McGrew. American metal typefaces of the Twentieth Century. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, second edition, 1993. ISBN 0-938768-34-4 (hardcover), 0-938768-39-5 (paperback). xx + 376 pp. LCCN Z250 .M137 1993. US$75.00 (hardcover), US$49.95 (paperback).

McSherry:1984:CTG

[McS84] James E. McSherry. Computer typesetting: a guide for authors, editors, and publishers. Opendoor Press, Arlington, VA, USA, 1984. ISBN 0-912162-05-8 (paperback). 106 pp. LCCN Z253.3 .M38 1984. US$9.50.

Meggs:1998:HGD

[Meg98] Philip Meggs. A History of Graphic Design. John Wiley and Sons, Inc., New York, NY, USA, third edition, 1998. ISBN 0-471-29198-6. xiii + 510 pp. LCCN Z246.M43 1998. US$75.00.

Mergenthaler:1898:BOM

[Mer98] Ottmar Mergenthaler. The biography of Ottmar Mergenthaler, inventor of the Linotype. Mergenthaler Linotype Company, Baltimore, MD. USA, 1898. 71 pp. LCCN Z253 .M48. See also extended edition [MS89].

Metzner:1975:MUO

[Met75] A. W. Kenneth Metzner. Multiple use and other benefits of computerized publishing. IEEE Transactions on Professional Communication, PC-18(3):274– 278, September 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Mittelbach:2023:LCPb

[MFB^{+}23] Frank Mittelbach, Ulrike Fischer, Javier Bezos, Johannes Braams, and Joseph Wright. The LaTeX Companion: Part II. Tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, third edition, 2023. ISBN 0-201-36300-3 (Part II) (hardcover), 0-13-816657-9 (Part II) (PDF), 0-13-816652-8 (Part II) (ePub). xxxviii + 970 pp. LCCN Z253.4.L38 M58 2023.

Mittelbach:2023:LCPa

[MFW23] Frank Mittelbach, Ulrike Fischer, and Joseph Wright. The LaTeX Companion: Part I. Tools and techniques for computer typesetting. Addison-Wesley, Reading, MA, USA, third edition, 2023. ISBN 0-13-465894-9 (Part I) (hardcover), 978-013-465-9-5-9-6 (Part I) (PDF), 978-013-465-9-6-0-2 (Part I) (e-pub), 0-13-816648-0-X (Part I + II). xxviii + 947 pp. LCCN Z253.4.L38 M58 2023.

Messina:1970:EPA

[MH70] Carla G. Messina and Joseph Hilsenrath. Edit-insertion programs for automatic typesetting of computer printout. NBS technical note 500 NBS technical note; 500. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1970. iii + 47 pp.

Martin:1998:NRD

[MH98] Shannon E. Martin and Kathleen A. Hansen. Newspapers of record in a digital age: from hot type to hot link. Praeger series in political communication. Praeger, Westport, CT, USA, 1998. ISBN 0-275-95960-0. ISSN 1062-5623. xiv + 159 pp. LCCN ????

Miller:1955:FTS

[Mil55] C. William Miller. Franklin’s type: Its study past and present. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 99(6):418– 432, December 1955. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

                                        URL https://www.jstor.org/
                                        stable/3143924.

Milde:2012:IMS

[Mil12] Günter Milde. isomath: Mathematical style for science and technology. Web site, September 4, 2012. URL http://get-

                                        software.net/macros/latex/
                                        contrib/isomath/isomath.html.

Murphy:1989:MTO

[MK89] David Murphy and David Kember. Mathematical typesetting from an operator perspective. Educational Technology, 29(4): 40–44, April 1989.

HMLS:1982:HT

[MLS82] History of typesetting, 1982.

1 videocassette (U-matic) (33 min.).

Marks:1973:TCC

[MM73] Robert H. Marks and A. W. Kenneth Metzner. Typewriter composition cuts journal costs, speeds publication. IEEE Transactions on Professional Communication, PC-16(3):73– 79, September 1973. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

McGilton:1990:TTU

[MM90] Henry McGilton and Mary Mc-Nabb. Typesetting tables on the UNIX system. Trilithon Press, Los Altos, CA, USA, 1990. ISBN 0-9626289-0-5. xxii + 282 pp. LCCN Z253.4.U53M33 1990.

Moitra:1979:DAH

[MMN79] Abha Moitra, S. P. Mudur, and A. W. Narwekar. Design and analysis of a hyphenation procedure. Software—Practice and Experience, 9(4):325–337, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Millar:1977:AFP

[MN77] J. B. Millar and M. C. Newey. Arbitrary font printing system. Papers and Presentations — Proceedings of the Digital Equipment Computer Users Society, 3 (5):1641–1647, ???? 1977. CO-DEN DEPPAD.

Mudur:1979:DST

[MNM79] S. P. Mudur, A. W. Narwekar, and Abha Moitra. Design of software for text composition. Software—Practice and Experience, 9(4):313–323, April 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Molnar:1993:TGT

[Mol93] Suzanne M. Molnar. A tour of graphical typesetting on the Macintosh. The UMAP Journal, 14(2):97–112, 1993.

Morison:1973:TT

[Mor73]  Stanley Morison.  A tally

of types. Cambridge University Press, Cambridge, UK, 1973. ISBN 0-521-20043-1 (hardcover), 0-521-09786-X (paperback). 137 pp. LCCN Z232.C17 M6 1973. US$55.00. Edited by Brooke Crutchley.

Morgan:1987:FDP

[Mor87] Bradford A. Morgan. Future of desktop publishing in technical communications. In IPCC 87 International Professional Communication Conference: Engineering Communication, a Byte into the Future, Winnipeg, Manitoba, Canada, pages 231–232.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. IEEE Service Cent. Piscataway, NJ, USA.

Morris:1989:RDT

[Mor89] R. A. Morris. Rendering digital type. A historical and economic view of technology. The Computer Journal, 32(6):524– 532, December 1989. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Morreale:1993:SAM

[Mor93] J. Morreale, editor. Seminar:

Annual meeting — May 1993, Seattle, WA, volume 24 of SID International Symposium Digest of Technical Papers 1993. SID, ????, 1993. ISSN 0097-966X (print), 2154-6738 (electronic), 2154-6746.

Ma:1994:AGA

[MP94] Xiaohu Ma and Zhigeng Pan.

Automatic generation algorithm of high-quality outline font using Bézier curve. Zidonghua Xuebao/Acta Automatica Sinica, 20 (1):121–125, January 1994. CO-DEN THHPAY. ISSN 0254-4156.

Macdonald:1981:TDM

[MPP81] John B. Macdonald, Mary Kay Podlecki, and Milt J. Pappas. Technical documentation by MAGIC. IEEE Computer Graphics and Applications, 1(2): 27–30, 32, 34, 36, April 1981. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Ma:1995:AGC

[MPZ95] Xiaohu Ma, Zhigeng Pan, and Fuyan Zhang. Automatic generation of Chinese outline font based on stroke extraction. Journal of computer science and technology, 10(1):42–52, January 1995. CODEN JCTEEM. ISSN 1000-9000.

Mittelbach:1992:FHQ

[MR92a] Frank Mittelbach and Chris Rowley. The future of high quality typesetting: structure and design. In Zlatuška [Zla92], pages 255–256. ISBN 80-210-0480-0. Invited talk.

Mittelbach:1992:PQH

[MR92b] Frank Mittelbach and Chris Rowley. The pursuit of quality: How can automated typesetting achieve the highest standards of craft typography? In Vanoirbeek and Coray [VC92], pages 261–273. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Marshall:1993:PST

[MRK93] Gerald F. Marshall, Gerald A. Rynkowski, and Mehrdad Ketabchi. Polygonal scanners for TV and HDTV laser projectors: spatial and temporal tolerances versus resolution. Proceedings of SPIE — The International Society for Optical Engineering, 1988:223–241, 1993.

CODEN PSISDG. ISBN 0-8194-1237-6. ISSN 0277-786X (print), 1996-756X (electronic).

McThompson:1983:ECF

[MRS83] O. W. McThompson, R. F.

Romon, and L. A. Sendelbach. Ergonomic character font for CRT displays. IBM Technical Disclosure Bulletin, 26(4):2120– ??, September 1983. CODEN IBMTAA. ISSN 0018-8689.

Morris:1989:CFS

[MRT89] R. J. T. Morris, L. D. Rubin, and H. Tirri. Comparison of feedforward and self-organizing approaches to the font orientation problem. In International Joint Conference on Neural Networks — volume II (Jun 18–22 1989: Washington, DC, USA), pages 291–297. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89CH2765-6.

Mursi:1984:CFS

[MS84] Mona M. Mursi and Mohamed Saba. Character font simulation system. Modeling and Simulation on Microcomputers, pages 131–135, 1984. CODEN MSMIE8.

Mergenthaler:1989:BOM

[MS89] Ottmar Mergenthaler and Carl Schlesinger. The biography of Ottmar Mergenthaler, inventor of the Linotype. Oak Knoll series on the history of the book. Oak Knoll Books, New Castle, DE, USA, 1989. ISBN 0-938768-12-3, 0-938768-13-1 (special edition). xviii + 125 pp. LCCN Z253 .M48 1989. URL http:

//www.nytimes.com/2014/11/
17/business/media/carl-schlesinger-
88-dies-helped-usher-out-
hot-type-.html.  New edi-

tion with added historical notes based on recent findings, researched and edited by Carl Schlesinger. Introduction by Elizabeth Harris.

Mercer:1994:SP

[MTZ94] Alan Mercer, C. Bernhard Tilanus, and Hans-Juergen Zimmermann. Style and principles. European Journal of Operational Research, 72(1):1–3, January 1994. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Mulvany:1994:IB

[Mul94] Nancy C. Mulvany. Indexing books. Chicago guides to writing, editing, and publishing. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-55014-1. xiii + 320 pp. LCCN Z695.9 .M8 1994. URL

http://www.loc.gov/catdir/
description/uchi052/93022934.
html;   http://www.loc.gov/
catdir/toc/uchi051/93022934.
html.

Muller:1996:JMB

[Mül96] Lars Müller. Josef Müller-Brockmann — Designer: a Pioneer of Swiss Graphic Design. Chronicle Books, ????, 1996.

ISBN 1-56898-063-9. 262 pp. LCCN ???? US$95.00.

Mulvany:2005:IB

[Mul05] Nancy C. Mulvany. Indexing books. University of Chicago Press, Chicago, IL, USA, second edition, 2005. ISBN 0-226-55276-4. xiv + 315 pp. LCCN Z695.9 .M8 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0621/2005004214-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0621/
2005004214-d.html;   http:
//www.loc.gov/catdir/toc/
ecip057/2005004214.html.

Munro:1988:BTI

[Mun88] Kathy Munro. Bringing typesetting in-house: the options for publishers. Blueprint, London, UK, 1988. ISBN 0-948905-16-6 (paperback). xi + 132 pp. LCCN ????

Murray:2019:FTF

[Mur19] David J. Murray. Frame Technology and FrameMaker. IEEE Annals of the History of Computing, 41(3):62–74, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Mitchell:2005:BTD

[MW05] Michael Mitchell and Susan Wightman. Book typography: a designer’s manual. Libanus Press, Marlborough, Wiltshire, UK, 2005. ISBN 0-948021-66-7. xi + 434 pp. LCCN Z246 .M57 2005.

                                        URL http://catalogue.bnf.
                                        fr/ark:/12148/cb40927963h.

Mitchell:2017:TSH

[MW17] Michael Mitchell and Susan Wightman. Typographic Style Handbook. MacLehose Press, London, UK, 2017. ISBN 0-85705-753-7. 287 pp. LCCN Z246 .M58 2017.

Myers:1983:TTG

[Mye83] Patti Myers. Telecommunicating for typesetting: a guide for the user of an information system. National Composition Association, Arlington, VA, USA, 1983. 47 pp.

Myers:1984:TMM

[Mye84a] Patti Myers. Typesetting from magnetic media: a guide for the user of an information system. National Composition Association, Arlington, VA, USA, 1984. 47 pp.

Myers:1984:WPI

[Mye84b] Patti Myers. Word processing’s impact on today’s typesetting. An NAPL management publication. The Association, Research and Educational Foundation, Teaneck, NJ, USA, 1984. v + 17 pp.

Myhill:1990:PTI

[Myh90] M. Myhill. Publishing TEXnical INFORMATION: typesetting by database. Program, 24(3): 281–289, July 1990. CODEN PRGMBD. ISSN 0033-0337.

Nass:1994:LCE

[Nas94] Richard Nass. Laser-printer controller eases memory demands. Electronic Design, 42(22):139– 140, October 1994. CODEN ELODAW. ISSN 0013-4872.

NCA:1982:WWT

[Nat82] National Composition Association, Arlington, VA, USA. Who’s who in typesetting, 1982. 64 pp.

NCSL:1986:DPT

[Nat86] National Conference of State Legislatures. Desk-top publishing and typesetting, 1986. 2 sound cassettes (ca. 115 min.).

NCA:1987:GTC

[Nat87a] National Composition Association, 1730 N. Lynn St., Arlington, VA 22209, USA. Glossary of typography, computerized typesetting and electronic publishing terms, 1987. 112 pp.

NCAUS:1987:GTC

[Nat87b] National Composition Association, Arlington, VA, USA. Glossary of typography, computerized typesetting and electronic publishing terms, 1987. 112 pp.

NCAUS:1981:NRT

[NCA81] NCA reports on typesetting, 1981. National Composition Association, Arlington, VA, USA.

Nievergelt:1982:DPS

[NCNS82] J. Nievergelt, G. Coray, J.-D Nicoud, and A. C. Shaw, editors. Document preparation systems. Elsevier North-Holland, Inc., New York, NY, USA, 1982. ISBN 0-444-86493-8. xiv + 274 pp. LCCN Z244 .D63 1982. US$46.50.

NC:1989:NR

[NCP89] NCPA reports, 1989. National Composition and Prepress Association, Arlington, VA, USA.

Nedbal:1989:PTM

[Ned89] Richard Nedbal. Photoplotting on a typesetting machine. Printed Circuit Design, 6(9):33– 36, September 1989. CODEN PCIDEU. ISSN 0884-9862.

Negru:1988:CT

[Neg88] John Negru. Computer typesetting. Van Nostrand Reinhold Co., New York, NY, USA, 1988. ISBN 0-442-26696-0. xiv + 185 pp. LCCN Z253.3.N4 1988. US$39.95.

Nelson:1988:HHS

[Nel88] Charles Nelson. Help with help: a selected bibliography of desktop publishing resources. Technical Communication, 35 (4):313–317, November 1988. CODEN TLCMBT. ISSN 0049-3155.

Nies:1980:BYO

[Nie80] James B. Nies. Buying your own typesetting equipment. Communication: Journalism Education Today (C:JET), 14(1):14– 15, Fall 1980.

NIST:1998:TSS

[NIS98] NIST. Typefaces for symbols in scientific manuscripts. Report, National Institute for Standards and Technology, Gaithersburg, MD, USA, January 4, 1998. 2 pp. URL https:/

/physics.nist.gov/cuu/pdf/
typefaces.pdf.

NIST:2004:SUR

[NIS04] NIST. SI unit rules and style conventions: Check list for reviewing manuscripts. Report, National Institute for Standards and Technology, Gaithersburg, MD, USA, Septem-

ber 2004.  2 pp.  URL

https://physics.nist.gov/
cuu/Units/checklist.html.

Nizamani:1990:DMD

[Niz90] A. S. Nizamani. On dot matrix displays for Arabic text. IEEE Transactions on Consumer Electronics, 36(2):76–83, May 1990. CODEN ITCEDA. ISSN 0098-3063.

Nagashima:1989:DOF

[NKMO89] Ichiro Nagashima, Naoyuki Kai, Tsutomu Minagawa, and Masahide Ohhashi. Design of an outline font rasterizing LSI. In Symposium on VLSI Circuits Digest of Technical Papers (May 25–27 1989: Kyoto, Japan), pages 117–118. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. IEEE catalog number 89TH0262-6.

Noot:1983:STF

[Noo83] Han Noot. Structured text formatting. Software—Practice and Experience, 13(1):79– 94, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Namane:1990:CSC

[NSA90] A. Namane and M. A. Sid-Ahmed. Character scaling by contour method. IEEE Transactions on Pattern Analysis and Machine Intelligence, 12(6):600– 606, June 1990. CODEN IT-PIDJ. ISSN 0162-8828.

Ocker:1971:PHE

[Ock71] W. A. Ocker. Program to hyphenate English words. IEEE Trans Eng Writing Speech, EWS-14(2):53–9, June 1971.

Ocker:1975:PHE

[Ock75] Wolfgang A. Ocker. Program to hyphenate English words. IEEE Transactions on Professional Communication, PC-18 (2):78–84, June 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Oppenheim:1985:OIS

[Opp85] Charles Oppenheim. Online information services: Present plans and future prospects: ‘the Pergamon Infoline view’. International Journal of Micrographics & Video Technology, 4(2): 85–89, 1985. CODEN IJMTDZ. ISSN 0743-9636.

Padmanabhan:1987:MCD

[Pad87]  K. Padmanabhan.  Multi-

language CRT display circuit and software. IETE Technical Review (Institution of Electronics and Telecommunication Engineers), 4(7):285–291, July 1987. CODEN ITREEI. ISSN 0255-9609.

Perrousseaux:2005:HET

[PAL16] Yves Perrousseaux, Jacques André, and Christian Laucou. Histoire de l’écriture typographique. (French) [History of typographic writing]. Atelier Perrousseaux, Méolans-Revel, France, 2005-2016. ISBN 2-911220-13-7 (vol. 1), 2-911220-24-2 (vol. 2, pt. 1), 2-911220-84-6 (vol. 3), 2-911220-34-X (vol. 2/2), 2-36765-005-5 (vol. 4/1), 2-36765-006-3 (vol. 4/2). LCCN Z124 .P47 2005. URL http://

tinyurl.com/ja-xxieme. Four

volumes in six.

Pappas:1989:PRT

[Pap89] T. L. Pappas. Product reviews — TEXnology on the IBM PC. Computer, 22(8):111–120, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1991:TTD

[Pap91] T. L. (Frank) Pappas. TEX typesetting with a difference. Computer, 24(6):92–93, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parisi:1975:CIA

[Par75]  Paul A. Parisi.  Composi-

tion innovations at the American Society of Civil Engineers. IEEE Transactions on Professional Communication, PC-18 (3):244–273, September 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Partl:1990:HMT

[Par90] Hubert Partl. How to make TEX and LaTEX international. Computer Physics Communications, 61(1–2):190–200, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

                                        www.sciencedirect.com/science/
                                        article/pii/001046559090116I.

Parida:1993:CTG

[Par93a] Laxmi Parida. Computational technique for general shape deformations for use in font design. Computers & Graphics (Pergamon), 17(4):349–356, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parida:1993:VIC

[Par93b] Laxmi Parida. Vinyas: An interactive calligraphic type design system. IFIP Transactions B: Computer Applications in Technology, B-9:287–299, 1993. CO-DEN ITBTEH. ISBN 0-444-81564-3. ISSN 0926-5481.

Pal:1994:OBI

[PC94] U. Pal and B. B. Chaudhuri.

OCR in Bangla: An Indo-Bangladeshi language. Proceedings — International Conference on Pattern Recognition, 2:269– 273, 1994. CODEN PICREG. ISSN 1051-4651. IEEE catalog number 94CH3440-5.

Peck:2018:RNE

[Pec18] Robert McCracken Peck. The remarkable nature of Edward Lear. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 162(2):158–190, ???? 2018. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.amphilsoc.org/
sites/default/files/2018-
11/attachments/Peck.pdf.

Anonymous:1958:PEE

[PEE58] Printing equipment engineer, 1958. ISSN 0096-9400. Willsea Pub. Co., Cleveland, OH, USA.

Perry:1988:PAC

[Per88] Tekla S. Perry. ‘PostScript’ prints anything: a case history. IEEE Spectrum, 25(5):42–46, May 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Phelps:1961:OTT

[Phe61] Charles E. Phelps. The origin of trade typesetting: a brief sketch. Andersen Typesetting Co., Minneapolis, MN, USA, 1961. 10 pp.

Phelps:1989:NWT

[Phe89] Paul E. Phelps. Now we also have to teach typesetting. Community College Journalist, 17 (1):11–13, Winter 1989.

Phillips:1968:CPT

[Phi68] Arthur H. Phillips. Computer peripherals and typesetting: a study of the man-machine interface incorporating a survey of computer peripherals and typographic composing equipment. His Majesty’s Stationary Office, London, UK, 1968. xxii + 665 pp. LCCN Z253.3 .P49.

Phillips:1979:WLC

[Phi79] R. J. Phillips. Why is lower case better? Applied Ergonomics, 10(4):211–214, December 1979. CODEN AERGBW. ISSN 0003-6870.

PHM:1980:ITH

[PHM80] Introduction to typography history of typesetting technology, 1980. 2 filmstrips guide + 8 black line masters.

Pierson:1971:CPE

[Pie71] J. L. Pierson. Computer program for electronic typesetting. IEEE Trans Eng Writing Speech, EWS-14(2):46–52, June 1971.

Pinkham:1981:PCP

[Pin81] Charles B. Pinkham. Producing cost-effective publications with the aid of wp. Proceedings — International Technical Communication Conference, pages A. 64– A. 68, 1981. CODEN PITCDL. ISBN 0-914548-34-4.

Pinchback:1994:DSA

[Pin94] R. M. Pinchback. Denotational semantics applied to the typesetting of Maple expressions. In Lopez [Lop94], pages 22– 27. ISBN 3-7643-3791-5. LCCN QA76.95.M36 1994.

Plass:1982:CBL

[PK82] M. F. Plass and D. E. Knuth.

Choosing better line breaks. In Nievergelt et al. [NCNS82], pages 221–242. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Potter:2002:WDT

[PK02] Norman Potter and Robin Kinross. What is a Designer: Things, Places, Messages. Hyphen Press, London, UK, fourth edition, 2002. ISBN 0-907259-16-2. 181 pp. LCCN ???? US$20.00.

Plass:1981:OPT

[Pla81]  Michael F. Plass.  Optimal

pagination techniques for automatic typesetting systems. Thesis (Ph.D.), Stanford University, Stanford, CA, USA, 1981. vi + 72 pp.

Plooij:19xx:DBG

[Ploxx] Edward Bernard Plooij. The Difference between the Gutenberg way of typesetting and my method for composing Arabic. ????, ????, 19xx. 10 pp.

Podell:1984:MME

[Pod84] Edwin J. Podell. Mathematics must be effective in technical communication. IEEE Transactions on Professional Communication, PC-27(2):97–100, June 1984. CODEN IEPCBU. ISSN 0099-9474.

Pohlen:2015:LBU

[Poh15]  Joep Pohlen.  Letterfontäne

(über Buchstaben) ; [das ultimative Handbuch zur Typografie]. Taschen, Köln, Germany, 2015. ISBN 3-8365-5450-X. 640 pp. LCCN ????

Post:1978:HEC

[Pos78] Jack M. Post. How to earn, control and maintain your profits in typesetting. Gama Communications, Salem, NH, USA, 1978. 136 pp.

Powers:1962:ETT

[Pow62] Stanley Piazza Powers. The effect of three typesetting styles on the speed of reading newspaper content. Thesis (M.A in Journalism and Communication), University of Florida, Gainesville, FL, USA, 1962. vii + 89 pp.

Powers:1986:PCP

[Pow86] Jack Powers. Personal computerbased pagination systems. Proceedings — IEEE Computer Society International Conference, pages 78–81, 1986. CODEN PCICDQ. ISBN 0-8186-0692-4. IEEE Service Cent. Piscataway, NJ, USA.

Powell:1990:TDP

[Pow90] Patrick Powell. Typesetting with DITROFF PostScript fonts and characters using Typeset — Tpsnew. Technical report TR 90-3, University of Minnesota, Institute of Technology, Computer Science Dept., Minneapolis, MN, USA, 1990. 9 pp.

Poynor:2001:OGL

[Poy01a] Rick Poynor. Obey the Giant: Life in the Image World. Birkhauser Boston Inc., Cambridge, MA, USA, 2001. ISBN 3-7643-6565-X. 224 pp. LCCN NC998.4.P69 2001. US$28.00.

Poynor:2001:T

[Poy01b] Rick Poynor. Typographica.

Princeton Architectural Press, ????, 2001. ISBN 1-56898-298-4. 144 pp. LCCN ???? US$50.00.

Anonymous:1930:PP

[PP’30] Printing production, 1930. Penton Publishing Company, Cleveland, OH, USA.

Perfect:1983:RCI

[PR83] Christopher Perfect and Gordon Rookledge. Rookledge’s classic international typefinder: the essential handbook of typeface recognition and selection. Sarema Press, London, UK, 1983. ISBN 0-85331-469-1. 268 pp.

Perfect:2004:RCI

[PR04] Christopher Perfect and Gordon Rookledge. Rookledge’s classic international typefinder: the essential handbook of typeface recognition and selection. Laurence King Publishing Ltd., London, UK, revised edition, 2004. ISBN 1-85669-406-2 (paperback). 279 pp. LCCN Z250 .P42 2004. Revised by Phil Baines, with a foreword by Adrian Frutiger.

Prabha:1987:AFT

[PRB87] Chandra Prabha, Duane Rice, and John Bunge. Access to the full text of nonfiction books: Design considerations from a study of public library users. Proceedings of the ASIS annual meeting, 24:196–200, 1987. CODEN PAISDQ. ISBN 0-938734-19-9. ISSN 0044-7870.

UAPS:1980:GDT

[Pri80] Printing Services, University of Arkansas, Fayetteville, AK, USA. A guide to design, typesetting, printing, new issue edition, 1980. various pp.

Pringle:1981:JFH

[Pri81]  Alison M. Pringle.  Justi-

fication with fewer hyphens. The Computer Journal, 24 (4):320–323, November 1981. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/320.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/321.tif;   http:/
/www3.oup.co.uk/computer_

journal/hdb/Volume_24/Issue_
04/tiff/322.tif;   http:/
/www3.oup.co.uk/computer_
journal/hdb/Volume_24/Issue_
04/tiff/323.tif.

Panoutsopoulos:1984:MEB

[PS84] A. Panoutsopoulos and C. Y.

Suen. Method of encoding binary patterns and font generation. In Proceedings of the 1984 IEEE International Conference on Systems, Man and Cybernetics, Halifax, NS, Can, pages 131–137. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. IEEE Service Cent. Piscataway, NJ, USA.

Anonymous:1965:PT

[PT’65] Photo typesetting, 1965. Society Publications, London, UK.

PI:1971:ICT

[Pub71] Publicate Incorporated: Dataflow Systems, Inc., Bethesda, MD, USA. An introduction to computer typesetting/photocomposition, 1971. 10 + 151 pp.

Anonymous:1954:PW

[PW’54] Printing world, 1954. ISSN 0032-8715. Stonhill and Gillis, London, UK.

Raciti:1970:DCC

[Rac70] S. A. Raciti. Digital control of a CRT phototypesetting system. Comput Des, 9(11):107–12, November 1970.

Radcliffe:1977:IAP

[Rad77] P. A. B. Radcliffe. Introduction of automation to printing and publishing. Electronics and power, 23(3):209–212, March 1977. CODEN ELP-WAQ. ISSN 0013-5127.

Rahtz:1987:PW

[Rah87] Sebastian Rahtz. The processing of words. In Sebastian Rahtz, editor, Information technology in the humanities: tools, techniques and applications, Ellis Horwood series in computers and their applications, pages 69–79 (of 184). Halsted Press, New York, USA, 1987. ISBN 0-470-20852-X. LCCN AZ105 .I54 1987. US$39.95.

Randall:1986:EET

[Ran86] Starr Duane Randall. The effect of editing and typesetting technology on the typographical error rate of selected newspapers. Thesis (Ph.D.), Department of Communication, University of Utah, 1986. xi + 99 pp.

Rehe:1978:UTJ

[Reh78] Rolf F. Rehe. Is unjustified typesetting justified? Technical Communication, 25(1):7–10, First Quarter 1978. CODEN TLCMBT. ISSN 0049-3155.

Reines:1966:VFC

[Rei66] J. Reines. Variable font character recognition machine. In Pattern Recognition, Proceedings of IEEE Workshop, Las Croabas and Dorado, Oct 24–26, 1968.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1966.

Ryabova:1977:OEL

[RF77] E. K. Ryabova and M. A. Faermark. Ob effektivnosti lyuminestsentnykh lamp raznogo spektra pri osveshchenii rabot s gartom v poligrafii. [efficiency of fluorescent lamps with various spectral characteristics in the lighting of hard lead printing works]. Svetotekhnika, ??(7): 12–14, July 1977. CODEN SVE-TAG.

Rice:1978:BD

[Ric78] Stanley Rice. Book Design: Systematic Aspects. R. R. Bowker Company, New York, NY, USA, 1978. ISBN 0-8352-1044-8 (vol. 1), 0-8352-1045-6 (vol. 2). xii + 274 pp. LCCN Z116.A3 R52.

Rice:1981:CTH

[Ric81] Stanley Rice. CRT typesetting handbook. Van Nostrand Reinhold Co., New York, NY, USA, 1981. ISBN 0-442-23889-4. vi + 409 pp. LCCN Z253.3.R53.

Ristroph:1988:MFG

[Ris88] John H. Ristroph. Multi-system font generation. Computers & industrial engineering, 15:467– 474, 1988. CODEN CINDDL. ISSN 0360-8352 (print), 1879-0550 (electronic).

Romano:1974:ATB

[RM74] Frank J. Romano and Eileen Mc-Manus. Automated typesetting: the basic course. GAMA Communications, Salem, NH, USA, 1974. 120 pp.

Romano:2019:HDP

[RM19] Frank J. Romano and Miranda Mitrano. History of Desktop Publishing. Oak Knoll Press, New Castle, DE, 2019. ISBN 1-58456-380-X (hardcover), 1-58456-381-8 (paperback). viii + 399 pp. LCCN Z253.53

                                        .R65 2019. URL https://www.
                                        oakknoll.com/pages/books/
                                        133473/frank-romano-with-
                                        miranda-mitrano/history-of-
                                        desktop-publishing.

Robinson:1972:SCT

[Rob72] John Robinson. Strike-on cold typesetting and the campus newspaper. Thesis (M.A.), East Texas State University, Commerce, TX, USA, 1972. vii + 75 pp.

Roehrich:1985:FAS

[Roe85] Johannes Roehrich. Fehlerlose Automatische Silbentrennung in Fachsprachlichen Texten. Angewandte Informatik, Applied Informatics, 27(5):198–206, May 1985. CODEN AWIFA7.

Rohne:1974:PP

[Roh74] Carl F. Rohne. Phototypesetting for proposals. Technical Communication, 21(3 3rd Q):12–14, ???? 1974. CODEN TLCMBT. ISSN 0049-3155.

Romano:1975:HES

[Rom75] Frank J. Romano. How to evaluate and specify a computer typesetting and data management system. National Composition Association, Arlington, VA, USA, 1975. 42 pp.

Romano:1976:HDS

[Rom76] Frank J. Romano. How to develop and sell new typesetting markets, 1976. 1 cassette. 2track. mono.

Romano:1986:MWTa

[Rom86a] Frank J. Romano. Machine writing and typesetting: the story of Sholes and Mergenthaler and the invention of the typewriter and the linotype. GAMA, Salem, NH, USA, 1986. ISBN 0-938853-00-7. 122 pp. LCCN Z253 .R581 1986.

Romano:1986:MWTb

[Rom86b] Frank J. Romano. Machine writing and typesetting: the story of Sholes and Mergenthaler and the invention of the typewriter and the linotype. GAMA, Salem, NH, USA, 1986. 146 pp.

Romano:2014:HLC

[Rom14] Frank J. Romano. History of the Linotype Company. RIT Press, Rochester, New York, 2014. ISBN 1-939125-03-0 (hardcover), 1-933360-60-7 (paperback), 1-933360-75-5 (e-book). xv + 463 pp. LCCN Z253 .R736 2014.

Rose:1992:HTP

[Ros92a] Kristoffer H. Rose. How to typeset pretty diagram arrows with TEX — design decisions used in Xy-pic. In Zlatuška [Zla92], pages 183–210. ISBN 80-210-0480-0.

Rose:1992:TDX

[Ros92b] Kristoffer H. Rose. Typesetting diagrams with XY-pc: user’s manual. In Zlatuška [Zla92], pages 273–292. ISBN 80-210-0480-0.

Rosen:2015:AHE

[Ros15] Michael Rosen. Alphabetical: how every letter tells a story. Counterpoint, Berkeley, CA, USA, 2015. ISBN 1-61902-483-7 (hardcover). xv + 431 pp. LCCN P211 .R65 2015.

Rogener:1996:TMM

[RPP96] Stefan Rogener, Albert-Jan Pool, and Ursula Packhauser. Typen machen Marken Machtig [English: Branding with type: How type sells]. Adobe Press, Mountain View, CA, USA, 1996. ISBN 1-56830-248-7. 120 pp. LCCN HF5825 .R5613 1995. US$18.00, CDN$25.00, UK£16.50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15683/bud/1568302487.html.

Edited by E. M. Ginger and translated by Stephanie Tripier.

Rosenblum:1983:CTC

[RRLG83] Louis Rosenblum, B. D. Rosenblum, G. P. Low, and W. W. Garth IV. Computerized Typesetting of Chinese: Text Processing on the Sinotype III. Instrument Society of America, Research Triangle Park, NC, USA, 1983. ISBN 0-87664-741-7. 511– 527 pp. LCCN TA165.M52 1983.

Rich:1965:MHA

[RS65] R. P. Rich and A. G. Stone.

Method for hyphenating at the end of a printed line. Communications of the ACM, 8(7): 444–445, July 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramsay:1981:CVC

[RS81] Mike Ramsay and Theodore A.

Shaffer. Custom video control delivers advanced font graphics. Electronics, 54(25):151– 154, December 1981. CODEN ELECAD. ISSN 0013-5070.

Rumble:1984:ADD

[RS84] John R. Rumble, Jr. and N. L.

Seidenman. Analysis and display of data in science and technology. Database Manage in Sci and Technol, pages 75–91, 1984. ISBN 0-444-86865-8.

Rubinstein:1988:DTI

[Rub88] Richard Rubinstein. Digital Typography—An Introduction to Type and Composition for Computer System Design. Addison-Wesley, Reading, MA, USA, November 1988. ISBN 0-201-17633-5. xi + 340 pp. LCCN Z253.3 .R8 1988. US$32.25. Reprinted with corrections.

Sachs:1980:ETS

[Sac80] Jonathan Sachs. Economical typesetting from small computer text files. SIGSMALL Newsletter (ACM Special Interest Group on Small Computing Systems and Applications), 6(2): 184–188, September 1980. CO-DEN SNSLD9.

Sacco:1986:TBP

[Sac86] Ralph Sacco. Typesetting and basic pasteup, 1986. 1 videocassette (17 min.).

Samson:1963:TTJ

[Sam63] Peter Samson. TJ-2: Type justifying program. Memo PDP-9, PDP-1 Computer, Massachusetts Institute of Technology, Cambridge, MA, USA, May 9, 1963. 4 pp. URL http://www.

                                        dpbsmith.com/tj2.html.

Samet:1982:HLD

[Sam82] Hanan Samet. Heuristics for the line division problem in computer justified text. Communications of the ACM, 25(8):564– 571, 1982. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sandberg:1995:COE

[San95] Jonathan S. Sandberg. Counting OCR errors in typeset text. Proceedings of SPIE — The International Society for Optical Engineering, 2422:184–195, 1995. CODEN PSISDG. ISBN 0-8194-1769-6. ISSN 0277-786X (print), 1996-756X (electronic).

Sanidas:1999:BRB

[San99] Amy E. Sanidas. Book review: Newspapers of record in a digital age: From hot type to hot link, by Shannon E. Martin and Kathleen A. Hansen. Journal of the American Society for Information Science, 50(9):840–841, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saxe:1992:AIH

[Sax92] Stephen O. Saxe. American iron hand presses. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, 1992. ISBN 0-938768-35-2 (hardcover), 0-938768-36-0 (paperback). xiv + 108 pp. LCCN Z249 .S23 1992. US$35.00 (hardcover), US$12.95 (paperback).

Silver:1993:PRW

[SB93] N. Clayton Silver and Curt C.

Braun. Perceived readability of warning labels with varied font sizes and styles. Safety Science, 16(5-6):615–625, August 1993. CODEN SSCIEO. ISSN 0925-7535.

Smura:1989:FII

[SBSG89] Edwin J. Smura, Barbara Beeton, Karla Savage, and Alan Griffee. Font information interchange standard ISO/IEC 9541. Computer Communications, 12 (2):93–96, April 1989. CO-DEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

Schell:1983:SPA

[Sch83] K. J. Schell. Security printers application of lasers. Proceedings of SPIE — The International Society for Optical Engineering, 396:131–140, 1983. CO-DEN PSISDG. ISBN 0-89252-431-6. ISSN 0277-786X (print), 1996-756X (electronic).

Schaer:1986:IMF

[Sch86a] Hansrudolf Schaer. Die Integration Mathematischer Formeln in den Dokumenteneditor Lara. Informationstechnik it: Computer, Systeme, Anwendungen, 28(6): 352–360, 1986. CODEN IC-SAES.

Schulze:1986:ET

[Sch86b] Bernd Schulze. Einfuehrung in TEX. Informationstechnik it: Computer, Systeme, Anwendungen, 28(6):332–341, 1986. CO-DEN ICSAES.

Schreiber:1995:SCA

[Sch95] Fred Schreiber. Simon de Colines: an annotated catalogue of 230 examples of his press, 1520–1546. Friends of the Brigham Young University Library, Provo, UT, USA, 1995. ISBN ???? lxxxiv + 242 pp. LCCN Z232.C67 H37 1995. US$150. With an introduction by Jeanne Veyrin-Forrer.

Sylla:1988:HFD

[SDB88] Cheickna Sylla, Colin G. Drury, and A. J. G. Babu. Human factors design investigation of a computerized layout system of text-graphic technical materials. Human Factors, 30(3):347–358, June 1988. CODEN HUFAA6. ISSN 0018-7208.

Serio:1984:CTP

[Ser84] John N. Serio. Controlling typesetting with a personal computer. In Conference Record — PCC ’84, The Practical Aspects of Engineering Communication, Atlantic City, NJ, USA, pages 137–139. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. IEEE Service Cent. Piscataway, NJ, USA.

Serio:1988:MDP

[Ser88] John N. Serio. Maximizing desktop publishing software: High resolution at low cost. Technical Communication, 35(4):310–312, November 1988. CODEN TL-CMBT. ISSN 0049-3155.

Sewell:1989:WPL

[Sew89] Wayne Sewell. Weaving a program: literate programming in WEB. Van Nostrand Reinhold Co., New York, NY, USA, 1989. ISBN 0-442-31946-0. xx + 556 pp. LCCN QA76.73.W24 S491 1989. US$34.95.

Seybold:1966:BRB

[Sey66]  John W. Seybold.  Book

review: 10/66-2R Computer Typesetting: Experiments and Prospects. 1965. Michael T. Barnett. MIT Press, Cambridge, Mass. 245 pp. American Documentation, 17(4):219–220, October 1966. CODEN AMDOA7. ISSN 0096-946X.

Seybold:1982:DPS

[Sey82] J. W. Seybold. Document preparation systems and commercial typesetting. In Nievergelt et al. [NCNS82], pages 243–264. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Seybold:1984:WDTb

[Sey84a] J. W. Seybold. The world of digital typesetting: Supplement. Seybold Publications, San Francisco, CA, USA, 1984. 35 pp.

Seybold:1984:WDTa

[Sey84b] John W. Seybold. The world of digital typesetting. Seybold Publications, San Francisco, CA, USA, 1984. ISBN 0-918514-08-8. x + 428 pp. LCCN Z253.3.S435 1984.

Skillin:1974:WT

[SG74] Marjorie E. Skillin and Robert Malcolm Gay. Words into Type. Prentice-Hall, Inc., Englewood Cliffs, NJ, USA, third edition, 1974. ISBN 0-13-964262-5. xx + 583 pp. LCCN PN160.S52 1974. US$39.95.

Spiekerman:1993:SSS

[SG93] Erik Spiekerman and E. M. Ginger. Stop Stealing Sheep & Find Out How Type Works. Adobe Press, Mountain View, CA, USA, second edition, 1993. ISBN 0-672-48543-5. 174 pp. LCCN Z250.S738 1993. US$19.95, CDN$24.95.

Sofroniou:1994:TTU

[SH94] N. Sofroniou and G. D. Hutcheson. Technical typesetting using LaTEX. Behavior research methods, instruments, and computers, 26(3):369–??, ???? 1994. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Scull:2019:KAT

[SH19] John Scull and Hansen Hsu. The killer app that saved the Macintosh. IEEE Annals of the History of Computing, 41(3):42–52, July 2019. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Shapiro:1977:ECT

[Sha77] Charles Shapiro. Encyclopedia of contemporary typesetting. Graphic Arts Technical Foundation, Pittsburgh, PA, USA, 1977. ISBN 0-88362-014-6. vii + 192 pp. LCCN ????

Shafran:1985:NPC

[Sha85] Joan K. Shafran. New page:

the changing face of composition. Proceedings — Symposium on Small Computers in the Arts, pages 91–96, 1985. CODEN PSSAED. ISBN 0-8186-0689-4. IEEE Service Cent. Piscataway, NJ, USA.

Sherman:1950:GMT

[She50] Frank M. Sherman, editor. The genesis of machine typesetting: a series of talks on the origin and early development of various machines now in general use for setting type by mechanical means. M and L Typesetting and Electrotyping Co., Chicago, IL, USA, 1950. 52 pp.

Siebenmann:1986:TWU

[Sie86] L. C. Siebenmann. Toward wider use of TEX typesetting. A pre-TEX manifesto. Notices of the American Mathematical Society, 33(4):597–607, August 1986. ISSN 0002-9920 (print), 1088-9477 (electronic).

ACM:1982:USP

[SIG82] User services, proceedings of the Tenth ACM SIGUCCS conference (Chicago, IL, Nov. 10–12, 1982). ACM Press, New York, NY, USA, 1982.

Simon:1987:SCO

[Sim87] John J. Simon. From sand to circuits: and other inquiries. Harvard Univ. Of. for Info. and Tech., Cambridge, MA, USA, 1987. ISBN 0-674-32575-3. xi + 271 pp. LCCN TK7828.S48 1986. US$24.95. Includes chapter on computer typesetting.

Sinclair:1998:TW

[Sin98] Joseph T. Sinclair. Typography on the Web. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-645545-7. xxvi + 498 pp. LCCN TK5105.888. US$35.95, CDN$55.95.

Skrivanek:1959:SAT

[Skr59] Richard F. Skrivanek. Some attitudes toward typesetting machines in the nineteenth century. Thesis (M.A.), University of Minnesota, Minneapolis, MN, USA, 1959. 17 pp.

Stevens:1967:ATT

[SL67] Mary Elizabeth Stevens and John Llewellyn Little. Automatic typographic-quality typesetting techniques: a stateof-the-art review, volume 99 of National Bureau of Standards monograph. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1967. vi + 98 pp. URL https:

//digital.library.unt.edu/
ark:/67531/metadc13276/.

Slingerland:1981:STT

[Sli81] David W. Slingerland. Summary of trade and tariff information printing and typesetting machinery and printing plates, TSUS items 668.21, 668.23 (pt.), 668.25, 668.38, and 668.50 (pt.). USITC publication; 841 Control no. 6-4-7, U.S. International Trade Commission, Washington, DC, USA, 1981. vi + 31 pp.

Slingerland:1984:STT

[Sli84] David W. Slingerland. Summary of trade and tariff information: printing and typesetting machinery and printing plates: TSUS items 668.21, 668.23 (pt.), 668.25, 668.32, 668.34, 668.36, 668.38 and 668.50 (pt.). USITC publication; 841 Control no. 6-4-7 (supp.), U.S. International Trade Commission, Washington, DC, USA, 1984. 20 pp.

Smart:1979:T

[Sma79] John E. Smart. Typesetting. Science Museum, London, UK, 1979. 4 pp.

Smit:1984:HAH

[Smi84] Paul R. Smit. Hyphenation algorithm for HPWord. Hewlett– Packard Journal: technical information from the laboratories of Hewlett–Packard Company, 35(9):26–30, September 1984. CODEN HPJOAX. ISSN 0018-1153.

Smith:1989:CLF

[Smi89] N. E. Smith. Coping with a large Forth system. In Forsley [For89], pages 113–115. ISBN 0-914593-09-9. LCCN TJ212.2.A15 1989.

Smith:1990:PCS

[Smi90] N. E. Smith. Parsing character strings in Forth. In 1990 Rochester Forth Conference: Embedded Systems, pages 133–134. Institute for Applied FORTH Research, Rochester, NY, USA, 1990. ISBN 0-914593-10-2. LCCN QA76.73.F24 R59 1990.

Smith:2017:BLB

[Smi17] Eddie Smith. From boiling lead and black art: An essay on the history of mathematical typography. Web site., October 13, 2017. URL http:/

                                        /practicallyefficient.com/
                                        2017/10/13/from-boiling-lead-
                                        and-black-art.html.

Sennhauser:1994:DIC

[SO94] Rene Sennhauser and Krystyna W.

Ohnesorge. Document image compression using document analysis and block-class-specific data compression methods. Proceedings of SPIE — The International Society for Optical Engineering, 2186:146–155, 1994. CODEN PSISDG. ISBN 0-8194-1481-6. ISSN 0277-786X (print), 1996-756X (electronic).

Suzuki:1977:DTT

[SOE77] Saatoshi Suzuki, Satoshi Otake, and Shigemitsu Ebihara. Design of train timetable automatic editing system by smallscale computers. Hitachi Review, 26(3):121–126, ???? 1977. CO-DEN HITAAQ.

Soiffer:1995:MTM

[Soi95] Neil Soiffer. Mathematical typesetting in Mathematica. In Levelt [Lev95], pages 140–149. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Somberg:1990:CAR

[Som90] Benjamin L. Somberg. Character aspect ratio and design tradeoffs. Proceedings of the Human Factors Society, pages 1461–1464, 1990. CODEN PHFSDQ. ISSN 0163-5182.

Southall:2005:PTT

[Sou05]  Richard Southall.  Printer’s

type in the twentieth century: manufacturing and design methods. British Library, London, UK, 2005. ISBN 0-7123-4812-3 (British Library), 1-58456-155-6 (Oak Knoll Press). xv + 238 pp. LCCN Z250.A2 S68 2005.

Smura:1987:TNB

[SP87] Edwin J. Smura and Archie D. Provan. Toward a new beginning: the development of a standard for font and character encoding to control electronic document interchange. IEEE Transactions on Professional Communication, PC-30(4):259–264, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Slater:1972:CTP

[SPD72] J. F. Slater, A. H. Phillips, and C. J. Duncan, editors. Computer-aided typesetting: proceedings of an international conference held in London. Transcripta Books, London, UK, 1972. ISBN 0-903012-06-5. LCCN ????

Spivak:1980:JTG

[Spi80] M. Spivak. The joy of TEX: a gourmet guide to typesetting technical text by computer: with numerous explicit illustrations (version 1). American Mathematical Society, Providence, RI, USA, 1980. 134 pp. US$9.00.

Spivak:1982:JTG

[Spi82] Michael Spivak. The joy of TEX: a gourmet guide to typesetting technical text by computer: with numerous explicit illustrations. American Mathematical Society, Providence, RI, USA, version 0. edition, 1982. iv + 42 + 83 + 47 pp.

Spivak:1983:JTG

[Spi83] Michael Spivak. The joy of TEX: a gourmet guide to typesetting technical text by computer. American Mathematical Society, Providence, RI, USA, repr. with additions. edition, 1983. various pp.

Spivak:1986:JTG

[Spi86] Michael Spivak. The joy of TEX: a gourmet guide to typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-2999-8. xviii + 290 pp. LCCN Z253.4.T47S673 1986.

Spivak:1990:JTG

[Spi90] Michael D. Spivak. The Joy of TEX: a Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, second revised edition, 1990. ISBN 0-8218-2997-1. xxii + 309 pp. LCCN Z253.4.T47 S6731 1990.

US$40.00. URL https://ctan.
org/pkg/joy-of-tex.

Srinivasan:1993:UDP

[Sri93] Balasubramaniam (Balasubramaniam) Srinivasan. UNIX Document Processing and Typesetting. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-0605-4. xiv + 445 pp. LCCN Z52.5.U54S65 1993.

Seybold:1975:SRG

[SS75] John W. Seybold and Patricia Breuer Seybold. The Seybold report glossary and guide for automated typesetting. Seybold Publications, San Francisco, CA, USA, 1975. 44 + 17 pp.

Schmidt:1986:GTT

[SS86] Alfred Schmidt and Horst Stenzel. Graphik in TEX-Texten und Texte in Gks-Graphiken. Informationstechnik it: Computer, Systeme, Anwendungen, 28(6): 348–351, 1986. CODEN IC-SAES.

Sugita:1983:MKG

[SSST83] Takuya Sugita, Hisao Sakamoto, Kimiko Shima, and Masahide Tsukamoto. Multi-font Kanji generator. Transactions of the Institute of Electronics and Communication Engineers of Japan, Section E, E66(6):377– 382, June 1983. CODEN TIEEDU. ISSN 0387-236X.

Sakamoto:1985:CGH

[ST85] Munekazu Sakamoto and Mikio Takagi. Computerized Generation of High-Quality Fonts of the Japanese Alphabets. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0621-5. 786–791 pp. LCCN QA 75.5 C5865 1985. IEEE Service Cent. Piscataway, NJ, USA.

SCM:1971:PTT

[Sta71a] Statistics Canada. Manufacturing and Primary Industries Division, Ottawa, ON, Canada. Platemaking, typesetting and trade bindery industry. Industrie du clichage, de la composition et de la reliure commerciale, preliminary edition, 1971. various pp.

SCDIMP:1971:PTT

[Sta71b] Statistique Canada. Division des industries and manufacturières et primaires, Ottawa, ON, Canada. Platemaking, typesetting and trade bindery industry Industrie du clichage, de la composition et de la reliure commerciale, préliminaire edition, 1971. various pp.

Stallybrass:2006:BFP

[Sta06] Peter Stallybrass. Benjamin Franklin: Printed corrections and erasable writing. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 150(4):553–567, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.
jstor.org/stable/4599024.

Stewart:1919:TPI

[Ste19] Alexander A. Stewart. Typesetting: a primer of information about working at the case, justifying, spacing, correcting, making-up, and other operations employed in setting type by hand.

Typographic technical series for apprentices; pt. II, no. 16 Typographic technical series for apprentices; pt. 2, no. 16. Committee on Education, United Typothetae of America, Chicago, IL, USA, 1919. 94 + ix pp.

Stevens:1965:AIS

[Ste65] Mary Elizabeth Stevens. Automatic Indexing: A State-of-the-Art Report, volume 91 of National Bureau of Standards monograph. U.S. National Bureau of Standards, Gaithersburg, MD, USA, March 30, 1965. vi + 220 pp. URL https:

//digital.library.unt.edu/
ark:/67531/metadc70462/.

Stevens:1970:AIS

[Ste70] Mary Elizabeth Stevens. Automatic Indexing: A State-of-the-Art Report, volume 91 of National Bureau of Standards monograph. U.S. National Bureau of Standards, Gaithersburg, MD, USA, second edition, February 30, 1970. vi + 290 pp. URL https:

//digital.library.unt.edu/
ark:/67531/metadc171070/.

Steven:1987:PRC

[Ste87] Peter M. Steven. Problems of and requirements for character recognition systems. In RESNA ’87: Meeting the Challenge, Proceedings of the 10th Annual Conference on Rehabilitation Technology, San Jose, CA, USA, pages 666–668. RESNA-Assoc for the Advancement of Rehabilitation Technology, Washington, DC, USA, 1987.

Steinberg:1996:FHY

[Ste96] S. H. (Sigfrid Henry) Steinberg.

Five hundred years of printing. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, 1996. ISBN 1-884718-20-5 (Oak Knoll paperback), 1-884718-19-1 (Oak Knoll hardback), 0-7123-0414-2 (British Library hardback), 0-7123-0438-X (British Library paperback). x + 262 pp. LCCN Z124 .S8 1996. US$45.00 (hardcover), US$29.95 (paperback). Revised by John Trevitt.

Strange:1993:AWD

[Str93] Catherine L. Strange. Around the world in 80 days. typesetting Arabic, Hebrew, Cyrillic, and Thai. In IEEE [IEE93], pages 43–49 (or 43–48??). ISBN 0-7803-1466-2. LCCN T 10.5 I54 1993. IEEE catalog number 93CH3367-0.

Schuermann:1973:VLQ

[SV73] Juergen Schuermann and Leonore Vogt. Vergleich Linearer und quadratischer Klassifikatoren am Beispiel der Schriftart OCR-A. [comparison of linear and quadratic classifiers using the font OCR-A as an example]. NTZ Nachrichtentechnische Zeitschrift, NTZ Communications Journal, 26(6):269–274, June 1973. CODEN NNNCAZ.

Semenzato:1992:AAL

[SW92] Luigi Semenzato and Edward Wang. Aleph: a language for typesetting. In Vanoirbeek and Coray [VC92], pages 65–75. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

Swanson:1971:MTC

[Swa71]  Ellen Swanson.  Mathemat-

ics into Type: copyediting and proofreading of mathematics for editorial assistants and authors. American Mathematical Society, Providence, RI, USA, 1971. ISBN 0-8218-0053-1. x + 98 pp. LCCN T11 .S77.

Swanson:1979:MTC

[Swa79]  Ellen Swanson.  Mathemat-

ics into Type: copyediting and proofreading of mathematics for editorial assistants and authors. American Mathematical Society, Providence, RI, USA, revised edition, 1979. ISBN 0-8218-0053-1. ix + 90 pp. LCCN ????

Swanson:1986:MTC

[Swa86]  Ellen Swanson.  Mathemat-

ics into Type: copyediting and proofreading of mathematics for editorial assistants and authors. American Mathematical Society, Providence, RI, USA, revised edition, 1986. ISBN 0-8218-0053-1. ix + 90 pp. LCCN ???? Reprinted with corrections.

Swanson:1998:MTC

[Swa98]  Ellen Swanson.  Mathemat-

ics into Type: copyediting and proofreading of mathematics for editorial assistants and authors. American Mathematical Society, Providence, RI, USA, revised and updated. edition, 1998. ISBN 0-8218-0644-0. ix + 90 pp. LCCN ???? Reprinted with corrections.

Sadek:1991:TPC

[SZ91] George Sadek and Maxim Zhukov. Typography, polyglot: a comparative study in multilingual typesetting. The Center, New York, NY, USA, 1991. 35 pp.

Szoka:1982:GCR

[Szo82] Kathryn Szoka. A guide to choosing the right chart type. IEEE Transactions on Professional Communication, PC-25 (2):98–101, June 1982. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

UE:1970:T

[T’770] Typesetting, 1970. 35 min. sd. b and w. 16 mm.

Tuescher:1982:LDS

[TA82] Daniel Tuescher and Max Arnold. Latin/Arabic dual-font SP 300 ATU teleprinter. Hasler Review, 15(3):79–81, Autumn 1982. CODEN HSLRA5. ISSN 0374-3306.

Taylor:1916:CBC

[Tay16] Henry Huntly Taylor. Comparative basic costs of typesetting by hand, linotype, and monotype, of, 8, 10, and 12 point non-technical text matter, set both leaded and solid. Harvard University. Graduate School of Business Administration. Student reports on business problems, no. 1. Harvard University Press, Cambridge, MA, USA, 1916. 54 pp.

Tucker:1977:IAM

[TBF^{+}77] W. H. Tucker, R. H. T. Bates, Susan D. Frykberg, R. J. Howarth, W. K. Kennedy, M. R. Lamb, and R. G. Vaughan. Interactive aid for musicians. International Journal of Man-Machine Studies, 9(6):635–651, November 1977. CODEN IJMMBC. ISSN 0020-7373.

Tseng:1992:EKS

[TC92] L. Y. Tseng and C. T. Chuang. Efficient knowledge-based stroke extraction method for multifont Chinese characters. Pattern Recognition, 25(12):1445– 1458, December 1992. CO-DEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Tou:1982:ATF

[TCLM82] Julius T. Tou, Jack M. Cheng, D. Z. Lu, and John McCarthy. Automatic type font identification for falsified documents. Proceedings — International Conference on Pattern Recognition, pages 1234–1240, 1982. CODEN PICREG. IEEE Service Cent. Piscataway, NJ, USA.

Tenney:1976:STS

[Ten76] Leland Arthur Tenney. The selection of a typesetting system for journalism instruction at Oklahoma State University and production of the Daily O’Collegian. Thesis (M.S.), Oklahoma State University, Stillwater, OK, USA, 1976. v + 31 pp.

Terry:1968:PAT

[Ter68] David Wesley Terry. A problem in automatic typesetting. Thesis (M.S.), University of Kentucky, Lexington, KY, USA, 1968. vi + 83 pp.

TAPCCS:1984:TIP

[Tex84] Texas Adult Probation Commission. Communications Section, Austin, TX, USA. Typesetting/ illustration/paste-up/audio-visual/ guide, 1984. ii + 28 pp.

Thompson:1973:SEI

[TH73] Robert C. Thompson and Joseph Hilsenrath. SETAB: an edit/insert program for automatic typesetting of spectroscopic and other computerized tables. NBS technical note; 740. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1973. iii + 25 pp.

Thieme:1983:SBM

[Thi83] Romeo Thieme. Satz und Bedeutung mathematischer Formeln. (German) [Typesetting and meaning of mathematical formulas]. Werner-Verlag GmbH, Düsseldorf, Germany, 1983. ISBN 3-8041-3549-8. viii + 88 + 4 pp. LCCN QA41 .T45 1934a, QA41 .T44 1983. Reprint of the 1934 original. Edited by Karl Billmann, Helmut Bodden and Horst Nacke.

Thimbleby:1991:LTT

[Thi91] H. Thimbleby. Low-tech TEX? (typesetting program). In Proceedings of the Fourth Annual Conference on Computers and the Writing Process, pages 124– 130. Univ. Sussex, Brighton, UK, 1991.

Tervonen:1980:MCG

[THL80] Mikko Tervonen, Hannu Hakalahti, and Pentti Lappalainen. Microprogrammable character generator for CRT phototypesetting system. EUROMICRO Symposium on Microprocessing and Microprogramming, pages 179– 190, 1980. CODEN ESMMDX. ISBN 0-444-86098-3.

Thompson:1977:ABD

[Tho77] Susan Otis Thompson. American book design and William Morris. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1977. ISBN 0-8352-0984-9. xvii + 258 pp. LCCN Z116.A3 T6 1977. US$29.95.

Thompson:1996:ABD

[Tho96] Susan Otis Thompson. American book design and William Morris. Oak Knoll Books, 414 Delaware St., New Castle, DE 19720, USA, second expanded edition, 1996. ISBN 1-884718-25-6 (hardcover), 1-884718-26-4 (Oak Knoll paperback), 1-884718-25-6 (Oak Knoll hardcover), 0-7123-4515-9 (British Library paperback). xxix + 258 pp. LCCN Z116.A3 T47 1996. US$49.95 (hardcover), US$34.95 (paperback).

Thomas:2021:TLT

[Tho21] Edd Thomas. Turning letters into tones: A century ago, the optophone allowed blind people to hear the printed word. IEEE Spectrum, 58(7):34–39, 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

TBL:1952:TSB

[Til52] Tillotsons (Bolton) Ltd, ????

Type specimen book, incorporating the house style of typesetting and standard printers’ and authors’ proof corrections, new edition, 1952. 134 pp.

TBL:1955:THS

[Til55] Tillotsons (Bolton) Limited, ???? Tillotsons house style of typesetting: with standard printers’ and authors’ proof corrections, 1955. 95 pp.

Tachikawa:1983:MKS

[TKE83] Michiyoshi Tachikawa, Tatsuo Kasahara, and Koichi Ejiri. Multi-font Kanji synthesis by chain coded strokes. In Japan Display ’83, Proceedings of the 3rd International Display Research Conference, Kobe, Japan, pages 172–175. Soc for Information Display, Los Angeles, Calif, USA, 1983.

Townsend:2015:TPM

[Tow15] Adam Townsend. Is there a perfect maths font? Web site, August 13, 2015. URL

                                        http://chalkdustmagazine.
                                        com/blog/is-there-a-perfect-
                                        maths-font/.

Thrift:2002:DG

[TR02] Julia Thrift and Lucien E. Roberts. The Designer and the Grid. Rotovision, Hove, East Sussex BN3 1DD, England, 2002. ISBN 2-88046-678-4. 159 pp. LCCN ???? US$45.00.

Trickett:1969:DST

[Tri69] T. Trickett. Design of standard type font for optical character recognition. Honeywell Computer J, 3(1):3–11, Winter 1969.

Tufte:1983:VDQ

[Tuf83] Edward R. Tufte. The Visual Display of Quantitative Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1983. ISBN 0-9613921-0-X, 0-9613921-4-2. 197 pp. LCCN K27.S8 T84. US$34.00.

Tufte:1990:EI

[Tuf90] Edward R. Tufte. Envisioning Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1990. 126 pp. LCCN QA90 .T82 1990. US$48.00.

Tufte:1997:VEI

[Tuf97] Edward R. Tufte. Visual Explanations: Images and Quantities, Evidence and Narrative. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1997. ISBN 0-9613921-2-6. 156 pp. LCCN P93.5.T846 1997. US$45.00.

Tulley:1985:DDP

[Tul85] C. Tulley. Document delivery to a phototypesetting system. IERE Conference Proceedings, 63:137–143, 1985. CODEN IERCB6. ISBN 0-903748-60-6. ISSN 0538-0006.

Tuthill:1983:TUS

[Tut83] B. Tuthill. Typesetting on the UNIX system. BYTE Magazine, 8(19):253–265, October 1983. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

TDC:2002:TT

[Typ02] Type Directors Club. Typography 22 (Typography). Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 2002. ISBN 0-8230-5557-4; 0-06-018586-4. 298 pp. LCCN ???? US$59.95.

Uphaus:1995:ITA

[UB95] James A. Uphaus, Jr. and Richard Barry Bryant. Innovative tool for assessing the impact of cluster defects on dotmatrix display cockpit symbology. IEEE Proceedings of the National Aerospace and Electronics Conference, 1:558–565, 1995. CODEN NASEA9. IEEE catalog number 95CH35797.

Uphaus:1990:ICR

[UBR90] James A. Uphaus, Jr., Kristen Barthelemy, and John M. Reising. Improve character readability in spite of pixel failures — a better font. IEEE Proceedings of the National Aerospace and Electronics Conference, 1:278–283, 1990. CODEN NASEA9. IEEE catalog number 90CH2881-1.

Utting:1983:TSC

[UCBB83] I. A. Utting, D. J. Caul, H. Brown, and S. E. Binns. A typesetting server for the Cambridge ring. In Dallas and Spratt [DS83], pages 243–248. ISBN 0-444-86852-6. LCCN TK5105.5 .I3427 1983.

Uchikata:1984:HCF

[UKKS84] Hideki Uchikata, Seiji Koyama, Takayuki Kageyama, and Masahiro Sasaki. High-quality character font generation system. National Technical Report (Matsushita Electric Industry Company), 30 (3):124–132, June 1984. CO-DEN NTROAV. ISSN 0028-0291.

Ulukut:1986:CAP

[Ulu86] Sevin Ulukut. Computer Aided Publishing — die Revolution in der Technischen Dokumentation. Informationstechnik it: Computer, Systeme, Anwendungen, 6:386–390, 1986. CODEN ICSAES.

Umeda:1985:CMP

[UM85] Michio Umeda and Shinichi Meguro. Construction of a multifont printed Chinese character reader. Denki Tsushin Kenkyusho Kenkyu Jitsuyoka Hokoku, 34(1):47–58, 1985. CO-DEN DTKKAA. ISSN 0415-3200.

Umeda:1982:RMP

[Ume82] Michio Umeda. Recognition of multi-font printed Chinese characters. Proceedings — International Conference on Pattern Recognition, pages 793–796, 1982. CODEN PICREG. IEEE Service Cent. Piscataway, NJ, USA.

UMOSSRI:1974:GOB

[Uni74] University of Maine at Orono.

Social Science Research Institute and New England Regional Commission, Boston, MA, USA. Greater opportunities in Bangor and Brewer for typesetting, 1974. 21 pp.

UAPS:1978:GDT

[Uni78] University of Arkansas Printing Services, Fayetteville, AK, USA. A guide to design, typesetting, printing, 1978. 61 pp.

USNYBAA:1985:PWP

[Uni85a] University of the State of New York, The State Education Dept. Division of Education Business Management Services, Bureau of Administrative Analysis, Albany, NY, USA. Preparing word processing documents for typesetting, 1985. 8 pp.

UWMACC:1985:TUM

[Uni85b] University of Wisconsin, MACC Academic Computing Center, Madison, Wis. Typeset: user manual for the 1100, revision 1, 1985. various pp.

USENIX:1987:PSU

[USE87] USENIX Association, editor. Proceedings of the Summer 1987 USENIX Conference: June 8– 12, 1987, Phoenix, Arizona, USA. USENIX, Berkeley, CA, USA, 1987.

Vacha:1992:TTP

[Vác92] Ondrřej Vácha. TEX for typesetting in a publishing house. In Zlatuška [Zla92], pages 59–65. ISBN 80-210-0480-0.

VanWyk:1980:LTG

[Van80] Christopher John Van Wyk. A language for typesetting graphics. Cover title. thesis (ph. d.), Stanford University, Stanford, CA, USA, 1980. v + 59 pp.

VanWyde:1981:GTL

[Van81a] C. T. Van Wyde. A graphics typesetting language. In Abrahams [Abr81], pages 97– 107. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6.

VanWyk:1981:GTL

[Van81b] Christopher J. Van Wyk. A graphics typesetting language. ACM SIGPLAN Notices, 16 (6):99–107, June 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vanoirbeek:1992:EPE

[VC92] C. Vanoirbeek and G. Coray, editors. EP92—Proceedings of Electronic Publishing, ’92, International Conference on Electronic Publishing, Document Manipulation, and Typography, Swiss Federal Institute of Technology, Lausanne, Switzerland, April 7-10, 1992. Cambridge University Press, New York, NY, USA, 1992. ISBN 0-521-43277-4. LCCN Z 286 E43 I58 1992. UK£35. Prepared with LaTEX, and Microsoft Word converted to LaTEX.

vanderLaan:1992:TCT

[vdL92] Kees van der Laan. Typesetting crosswords via TEX. In Zlatuška [Zla92], pages 217–224. ISBN 80-210-0480-0.

Vismara:1993:TEA

[VDM^{+}93] Elena Vismara, Angela Donna, Francesco Minisci, Annamaria Naggi, and Nadia Pastori. Typesetting error–no article title given on the table of contents. Journal of Organic Chemistry, 58(4):959–??, February 1993. CODEN JOCEAH. ISSN 0022-3263.

vanLeunen:1992:HS

[vL92] Mary-Claire van Leunen. A handbook for scholars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised edition, 1992. ISBN 0-394-40904-3 (hardcover), 0-394-73395-9 (paperback). xi + 348 pp. LCCN PN146 .V36 1992.

vonSydow:1967:CTA

[von67] L. von Sydow. Computer typesetting of ALGOL. Communications of the ACM, 10(3):172– 174, March 1967.

Vincent:1994:DR

[VP94] Luc M. Vincent and Theo Pavlidis, editors. Document Recognition, volume 2181. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1994. CODEN PSISDG. ISBN 0-8194-1476-X. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK7895.O6 D63 1994.

vanVliet:1988:DMT

[vV88] J. C. van Vliet, editor. Document manipulation and typography: proceedings of the International Conference on Electronic Publishing, Document Manipulation, and Typography, Nice (France), April 20–22, 1988, The Cambridge series on electronic publishing. Cambridge University Press, New York, NY, USA, 1988. ISBN 0-521-36294-6. LCCN Z286.E43I57 1988.

Wade:1984:DFS

[Wad84] B. W. Wade. Digital font scaling method. IBM Technical Disclosure Bulletin, 26(9):4863–??, February 1984. CODEN IBM-TAA. ISSN 0018-8689.

Wagner:1985:DDM

[Wag85] Annette Wagner. Design Decisions Made in the Development of Type for Apple’s Lisa Computer. Natl Computer Graphics Assoc, Fairfax, VA, USA, 1985. ISBN 0-941514-09-9. 491–501 pp. LCCN ????

Walker:1981:TT

[Wal81] T. R. Walker. Typographical terminology. Proceedings — International Technical Communication Conference, pages A. 109–A. 111, 1981. CODEN PITCDL. ISBN 0-914548-34-4.

Waldock:1982:TAE

[Wal82] W. C. Waldock. Typesetting in the academic environment. In SIGUCCS’82 [SIG82], pages 25– 28.

Walker:1983:TT

[Wal83] T. R. Walker. Typographical terminology. In 1983 IEEE Professional Communication Society Conference Record: The Many Facets of Computer Communications, Atlanta, Ga, USA, pages 188–190. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. IEEE Service Cent. Piscataway, NJ, USA.

Wallis:1984:ETQ

[Wal84] Lawrence W. Wallis. Electronic typesetting: a quarter century of technological upheaval. Paradigm, Gateshead, 1984.

ISBN 0-9506104-3-7. xi + 135 pp. LCCN ????

Waldschmidt:1987:ECG

[Wal87]  H. Waldschmidt.  An edi-

tor for constructing graphics with TEX. Computer Graphics Forum, 6(4):359–364, December 1987. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wallis:1988:CCT

[Wal88] Lawrence W. Wallis. A concise chronology of typesetting developments, 1886–1986. Severnside Printers Ltd., Upton upon Severn, Eng., 1988. ISBN 0-85331-538-8. vii + 88 pp. LCCN Z253.W24 1988.

Wallis:1993:TSE

[Wal93] Lawrence W. Wallis. Typomania: selected essays on typesetting and related subjects. Severnside Printers, Upton-upon-Severn. UK, 1993. ISBN 1-874084-01-7. xiii + 161 pp. LCCN Z116.W295 1993.

WL:1979:THR

[Wan79] Wang Laboratories, Inc., Lowell, MA, USA. Typesetting hyphenation rules, 1979. various pp.

WTI:1969:PT

[War69] Warwick Typographers, Inc, St. Louis, MO, USA. Photographic typesetting, 1969. 100 pp.

WT:1985:PT

[War85] Warwick Typographers, St. Louis, MO, USA. Photographic typesetting, 1985. 202 + 2–34 pp.

Warnock:2012:SIC

[War12] John E. Warnock. Simple ideas that changed printing and publishing. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 156(4): 363–368, December 2012. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.
jstor.org/stable/23558230.

Warnock:2018:OP

[War18] John E. Warnock. The origins of PostScript. IEEE Annals of the History of Computing, 40(3):68– 76, July/September 2018. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

URL  https://ieeexplore.
ieee.org/document/8509542/.

Watt:1984:SDR

[Wat84] Roger W. Watt. Share-output device requirements. In Proceedings — SEAS Spring Meeting 1984: Processing Statistical Information, Knokke, Belgium, pages 591–603. SHARE European Association, Nijmegen, The Netherlands, 1984.

White:1995:PFS

[WB95] Ronald G. White and John Biard. Portable font specification. Dr. Dobb’s Journal of Software Tools for Professional Programmer, 20(3), March 1995. CODEN DDJTEQ. ISSN 1044-789X.

Witten:1982:PTD

[WBS82] I. H. Witten, M. Bonham, and E. Strong. On the power of traps and diversions in a document preparation language. Software —Practice and Experience, 12 (12):1119–1131, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wyatt:2018:TMR

[WD18] Christopher Scott Wyatt and Dànielle Nicole DeVoss, editors. Type Matters: the Rhetoricity of Letter Forms. Visual rhetoric. Parlor Press, Anderson, SC, USA, 2018. ISBN 1-60235-973-3 (paperback), 1-60235-977-6 (hardcover), 1-60235-978-4 (ebook), 1-60235-979-2 (e-pub), 1-60235-980-6 (iBook), 1-60235-981-4 (Kindle). xxi + 388 pp. LCCN Z250 .T935 2018.

Willberg:1999:EHT

[WF99] Hans Peter Willberg and Friedrich Forssman. Erste Hilfe in Typografie: Ratgeber für den Umgang mit Schrift. Schmidt, Mainz, Germany, 1999. ISBN 3-87439-474-3. 104 pp. LCCN ???? Library catalogs show both 1999 and 2000 for the year.

Warnock:2019:FGA

[WG19] John E. Warnock and Charles Geschke. Founding and growing Adobe Systems, Inc. IEEE Annals of the History of Computing, 41(3):24–34, July 2019. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Wolcott:1976:CCT

[WH76] Norman M. Wolcott and Joseph Hilsenrath. A contribution to computer typesetting techniques: tables of coordinates for Hershey’s repertory of occidental type fonts and graphic symbols. Technical Report 424, U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1976. 169 pp.

Wezler:1986:PXI

[WH86a] Albrecht Wezler and Ernst Hammerschmidt, editors. Proceedings of the XXXII International Congress for Asian and North African Studies, Hamburg, 25th– 30th August 1986, number IX in Zeitschrift der Deutschen Morgenländischen Gesellschaft, Supplement. Steiner, Stuttgart, Germany, 1986. ISBN 3-515-04808-1. LCCN DS1.5 .I565 1986.

Williams:1986:FSL

[WH86b] R. Don Williams and R. D.

Huchingson. Font style for large-area discrete-element displays. Digest of Technical Papers — SID International Symposium (Society for Information Display), 17:146–149, 1986. CO-DEN DTPSDS.

Wong:1995:DCT

[WH95] Paul Yiu Chung Wong and Siu Chi Hsu. Designing Chinese typeface using components. Proceedings — IEEE Computer Society’s International Computer Software & Applications Conference, pages 416–421, 1995.

CODEN PSICD2. ISSN 0730-6512. IEEE catalog number 95CB35838.

Walden:2017:DPP

[WH17] David Walden and Hiromichi Hashizume. Desktop publishing pioneer meeting at Computer History Museum. IEEE Annals of the History of Computing, 39 (3):65–67, July/September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/an/
                                        2017/03/man2017030065-abs.
                                        html.

White:1987:HST

[Whi87] Alex White. How to SPEC Type. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1987. ISBN 0-8230-2473-3. 128 pp. LCCN Z253.W46 1987. US$14.95.

Wick:1965:RTS

[Wic65] Karel Wick. Rules for Typesetting Mathematics. Publishing House of the Czechoslovak Academy of Sciences, Prague, Czechoslovakia, 1965. 99 pp. LCCN ???? Translated by Vaćlav Boublik and M. Hejlova.

Wichary:2023:SH

[Wic23] Marcin Wichary. Shift Happens. ????, ????, 2023. 1200+ (two volumes) pp.

                                        URL https://arstechnica.
                                        com/culture/2023/10/shift-
                                        happens-is-a-beautifully-
                                        designed-history-of-how-keyboards-
                                        got-this-way/.

WRC:1931:STA

[Wil31] William Reid and Company, Chicago, IL, USA. Special tools and appliances for those in care of typesetting and printing machines, 1931. 104 pp.

Wilson:1937:TAT

[Wil37] G. F. Wilson. The Trend of advertisement typography: as reflected by the typesetting department of Gee and Watson Ltd. Gee and Watson, London, UK, 1937. 49 + i pp.

WBP:1954:MT

[Wil54] William Byrd Press. Mathematics in Type. The William Byrd Press, Richmond, VA, USA, 1954. xi + 58 pp. LCCN Z250.6.M3 W5.

Willis:1989:EVG

[Wil89] Doug Willis. Easy Ventura:

guide to producing typesetting in one hour! Willis Publications, Hot Springs, Ark. (P.O. Box 1126, Hot Springs 71902), 1989. 38 + [21] pp.

Wind:1990:EFG

[Win90] Herbert Stephens Wind. An examination of fourth generation typesetting computers in Georgia newspaper. Thesis (M.A.), University of Georgia, Athens, GA, USA, 1990. v + 93 pp. Directed by E. M. Beck.

Woessner:1988:TMS

[WM88] Linda H. Woessner and Mary K.

McCaskill. A TEX macro set for typesetting NASA reports. Technical Report 101522, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1988. ?? pp.

Wohlfeil:1998:PCB

[Woh98] Stefan Wohlfeil. On the Pagination of Complex, Book-Like Documents. Shaker Verlag, Aachen, Germany and Maastricht, The Netherlands, 1998. ISBN 3-8265-3304-6. 224 pp. LCCN ???? DM 98.00.

URL http://www.shaker.de/
Online-Gesamtkatalog/Details.
idc?ID=24201&CC=59311&IDSRC=
1&ISBN=3-8265-3304-6&Reihe=
15;   http://wwwpi6.fernuni-
hagen.de/Forschung/Pagination/

.

Wolberg:1986:SOC

[Wol86]  George Wolberg.  Syntactic

Omni-Font Character Recognition System. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0721-1. 168–173 pp. LCCN TA1632 .I36 1986. IEEE Service Cent. Piscataway, NJ, USA.

Wonneberger:1986:VTT

[Won86a] R. Wonneberger. “Verheissung und Versprechen”, a third generation approach to theological typesetting. In Jacques Desarmenien, editor, TEX for scientific documentation: Proceedings of the second European conference, volume 236 of Lecture notes in computer science, pages 180–198. Springer-Verlag Inc., New York, NY, USA, 1986. ISBN 0-387-16807-9. LCCN Z253.4.T47 T491 1986.

Wonneberger:1986:TPT

[Won86b] Reinhard Wonneberger. TEX for philological typesetting. In Wezler and Hammerschmidt [WH86a], pages 595–596. ISBN 3-515-04808-1. LCCN DS1.5 .I565 1986.

Wonneberger:1990:SDP

[Won90] Reinhard Wonneberger. Structured document processing: the LaTEX approach. Computer Physics Communications, 61 (1–2):177–189, November 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/001046559090115H.

Wood:1987:TWG

[Woo87] Patrick Wood. Typesetting:

when gray isn’t. The PostScript Language Journal, 1(2):15–19, June 1987. ISSN 0891-5873.

Wright:1981:TCT

[Wri81] Donald L. Wright. Total concept in technical writing. Proceedings — International Technical Communication Conference, pages A. 112–A. 115, 1981. CO-DEN PITCDL. ISBN 0-914548-34-4.

Wright:1983:UAW

[Wri83] Walter Wright. Using an artists’ workstation. Proceedings — Symposium on Small Computers in the Arts, pages 128–142, 1983. CODEN PSSAED. ISBN 0-8186-0499-9. IEEE Service Cent. Piscataway, NJ, USA.

Weston:1926:RBI

[WT26] H. Claude (Hubert Claude) Weston and A. K. Taylor. The relation between illumination and efficiency in fine work (typesetting by hand): joint report of the Industrial Fatigue Research Board and the Illumination Research Committee. His Majesty’s Stationary Office, London, UK, 1926. v + 12 pp.

Weston:1928:EDS

[WT28] H. C. Weston and A. K. Taylor. The effect of different systems of lighting on output and accuracy in fine work (typesetting by hand); joint report of the Industrial Fatigue Research Board and the Illumination Research Committee. His Majesty’s Stationary Office, London, UK, 1928. iv + 12 pp.

Weston:1933:RBI

[WT33] Hubert Claude Weston and A. K. Taylor. The relation between illumination and efficiency in fine work (typesetting by hand). His Majesty’s Stationary Office, London, UK, 1933. 18 pp.

Warnock:1982:DIG

[WW82] John E. Warnock and D. K. Wyatt. A device independent graphics imaging model for use with raster devices. Computer Graphics, 16(3):313–319, July 1982. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Wilkinson:1993:USR

[WW93] R. A. Wilkinson and Charles L.

Wilson. Using self-organizing recognition as a mechanism for rejecting segmentation errors. Proceedings of SPIE — The International Society for Optical Engineering, 1825:378–388, 1993. CODEN PSISDG. ISBN 0-8194-1026-8. ISSN 0277-786X (print), 1996-756X (electronic).

Williams:2003:TAT

[WW03] Allison Muench Williams and J. Phillips Williams, editors. Typography 23: Annual of the Type Directors Club (Typography). Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 2003. ISBN 0-8230-5558-2. 300 (est.) pp. LCCN ???? US$59.95.

Yamakawa:1990:DPV

[YKT90] T. Yamakawa, Y. Kawabata, and H. Tamura. Desktop publishing viewed from the office automation industries. Journal of Information Processing, 31 (11):1508–1517, 1990. CODEN JOSHA4. ISSN 0447-8053.

Zaki:1987:FDA

[ZA87] M. Zaki and Al. H. Albarhamtoshy. Formal design of an Arabic text formatter for microcomputers. Computer Languages, 12 (2):123–143, 1987. CODEN CO-LADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Zapf:1954:MT

[Zap54] Hermann Zapf. Manuale typographicum. Schauer, Frankfurt am Main, West Germany, 1954. 48 + 1 pp. LCCN ????

Zapf:1968:MTT

[Zap68] Hermann Zapf. Manuale typographicum: 100 typographische Gestaltungen mit Aussagen über die Schrift, über Typographie und Druckkunst, aus Vergangenheit und Gegenwart, in 18 verschiedenen Sprachen. (German) [Typography manual: 100 typographical designs with statements about writing, about typography and printing, from past and present, in 18 different languages]. Z-Presse, Frankfurt am Main, West Germany, 1968. 119 pp. LCCN ????

Zapf:1970:MTT

[Zap70] Hermann Zapf. Manuale typographicum: 100 typographic pages with quotations from the past and present on types and printing in 16 different languages. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-24011-4 (hardcover), 0-262-74004-4 (paperback). 123 pp. LCCN Z250 .Z325.

Zapiti:1980:MGP

[Zap80] Kallie Zapiti. Model guidelines for the preparation of camera-ready typescripts by authors/typists, M. O’Connor, Ed.

IEEE Transactions on Professional Communication, PC-23 (4):204, December 1980. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Zapf:2007:ASC

[Zap07a] Hermann Zapf. Alphabet Stories: a Chronicle of Technical Developments. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-6-2. 149 pp. LCCN ????

Zapf:2007:ACT

[Zap07b]  Hermann Zapf.  Alphabet-

geschichten: Eine Chronik technischer Entwicklungen. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-5-4. 156 pp. LCCN ???? 29.00 EUR.

Zapf:1972:MTT

[ZB72] Hermann Zapf and Gertraude Benöhr. Manuale typographicum: 100 typographische Gestaltungen mit Aussagen über die Schrift, über Typographie und Druckkunst, aus Vergangenheit und Gegenwart, in 18 verschiedenen Sprachen. (German) [Typography manual: 100 typographical designs with statements about writing, about typography and printing, from past and present, in 18 different languages]. Büchergilde Gutenberg, Frankfurt am Main, West Germany, 1972. ISBN 3-7632-1628-6. 119 pp. LCCN ????

Zapf:1991:CTC

[ZD91] Hermann Zapf and John Dreyfus. Classical typography in the computer age: papers presented at a Clark Library seminar, 27 February 1988. William Andrews Clark Memorial Library, University of California, Los Angeles, CA, USA, 1991. ISBN ???? xv + 36 pp. LCCN Z250.7.Z36 1991. US$15.00. With an introduction by John Bidwell.

Zeyen:1974:PST

[Zey74] Fritz-Otto Zeyen. Program systems and typical applications of the digiset electronic photocomposing system. Siemens Review, 41(10):440–443, October 1974. CODEN SZTEA6.

Zuckerman:1984:CRT

[ZJ84] Alan E. Zuckerman and Keith W. Johnson. Converting a reference textbook into a computer searchable database: the experience with the USP DI. Proceedings — Annual Symposium on Computer Applications in Medical Care, pages 341–344, 1984. CODEN PCMCDC. ISBN 0-8186-0565-0. ISSN 0195-4210. IEEE Service Cent. Piscataway, NJ, USA.

Ziegler:1988:SZA

[ZK88] Holger G. Ziegler and Hermann Kaindl. Synthese zweier Ansaetze der Textverarbeitung. Angewandte Informatik, Applied Informatics, 30(10):449– 454, October 1988. CODEN AWIFA7. ISSN 0013-5704.

Zable:1987:FDH

[ZL87] J. L. Zable and H. C. Lee. Font design for high-speed impact line printers. IBM Journal of Research and Development, 31(6): 679–684, November 1987. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Zlatuska:1992:EPE

[Zla92]  Jiří Zlatuška, editor.  Eu-

roTEX ’92: Proceedings of the 7th European TEX Conference, Prague, Czechoslovakia, September 14–18, 1992, Proceedings of the European TEX Conference. Masarykova Universita, Brno, Czechoslovakia, September 1992. ISBN 80-210-0480-0.

Zurek:1985:TAP

[Zur85] Jerome Zurek. Take advantage of PCs: Use PCs for typesetting. Journalism Educator, 40(1):16– 19, Spring 1985.